A Bibliography of Publications in Dr. Dobb’s Journal: Software Tools for the Professional Programmer

(2000–2009)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

05 July 2016 Version 3.24

Title word cross-reference

#1 [Swa02q, Swa03k].

+ [Ste01h]. 3 [Cro00b, LJ01, Sny03]. = [Ste01h]. ^{3} [Eri02a]. N log N [Gon00]. R_{XML} [Eri01b].

.NET [Bar03, Bui02, Gri02, Gun02, Hel02, Rob03, SPRF02, Tim03, All02, DHR^{+}01, Rei01b, SPM02, dIJ02].

1 [She03]. 1.0 [Coc01c, Swa02k]. 108-bit [Bar00b]. 10th [Coc02f]. 1394 [Ale00].

2 [BC01]. 2.0 [Coc01e, Rao02, Wal03b]. 2000 [Cal00a, Coc03e, Mil00, SEC01, Ste00a, Ste00e]. 2001 [Coc02i]. 2002 [Coc03e]. 2003 [Eri03a]. 21st [LMM^{+}01, MKR^{+}01].

3.0 [Hei01, Van01]. 30th [Coc02f]. 3Dnow [Fom00b].

4 [Wal02a, Woe02a]. 4.0 [Met01].

6.1 [Van02a]. 64 [Jar01]. 64-Bit [Gut00, HBT^{+}00, KDG^{+}00].

802.11g [Coc02f].

= [SPRF02].

Academic [Bar00d]. Accelerating [Rot01b]. Accept [Coc03a]. Access [Bar00k, Gun01a, Smi01b]. Accidents [Swa03c]. Accolades [KDG^{+}00]. Ace [Byk00]. Achievement [Coc01d]. ACLU [Bar01e]. ACM [Bar00d, Bar00i, Bar00f, Coc02c, Isr01]. Acronyms [Bar01b]. Acrophilia [Swa00n]. Across [Coc03e, Mon00b]. Act [Bar00g]. ActiveX [Lar01a]. Adam [Swa03b]. Adaptive [Coc02f, Gon00, Jon02b]. Add [Bar01d]. Adding [Car02a, Car02b, Eri02c]. Addition [Dau01]. Address [CDsJ^{+}00]. Adele [Ano02b]. Adi [Coc03b]. Adleman [Coc03b]. ADO [Che01]. Adopts [Bar01h]. Advanced [Bar00l, Coc02k, DR01]. Adventures [Sto01, Sto02a, Sto02b, Sto02c, Sto02d, Sto02e, Sto02f, Sto02g, Sto02h, Sto02i, Sto02j, Sto02k, Sto02l, Sto03b, Sto03a, Sto03d, Sto03g, Sto03c, Sto03e, Sto03f]. Adversarial [Sha02d]. Adware [Ste00g]. AE [Van02b]. AES [Coc02k]. Again [Eri01f]. Agent [Bru02]. Agents [Jon02a, O’M01, Kie01b]. Aggregation [ZBKO01]. Aglets [Jon02a]. Agreement [Bar01d]. AI [Bar00k, Coc01a, Coc02l]. Ails [Eri01b]. Airport [Eri00f]. Al [TLA^{+}02]. Alert [Coc03g]. Algorithm [Ana00, Bar00j, Bar00l, Bar00b, Bar01c, Bar01d, Bar01g, Bar01e, Bay01, Ben00, Ben01, Car01, DR01, Dem01, DNM00, Get01a, Get01b, Gon00, Gut00, Gut02, HM01, Keo00, Kri01, Lev00, Mah02, Man00, Mar00, Mis00, Mit02a, Nel02b, Nil00, Pal01b, Pig02, Qur01, Rol00a, Rol00b, Rol02, Sco02, Sha02a, Smi02, Sta01a, Wie00, Win00]. Algorithms [HBT^{+}00, LMM^{+}01, TCM^{+}00]. Alias [Mit00]. Aliasing [HWT^{+}01]. Alien [Coc02b, Wil01c]. All-Pairs [Mah03]. Alley [Ana00, Bay01, Ben00, Ben01, Car01, DR01, Dem01, DNM00, Get01a, Get01b, Gon00, Gut00, Gut02, HM01, Keo00, Kri01, Mah02, Mar00, Mis00, Mit02a, Nel02b, Pal01b, Pig02, Qur01, Rol00a, Rol00b, Rol02, Sco02, Sha02a, Smi02, Sta01a, Wie00, Win00]. Alliance [Coc02j]. Allocation [Nak01]. Almost [Ste01d]. Alphabet [Coc02j]. Alphanumeric [Win00]. Alternating [Mar00]. Alternatives [Swa01a]. Always [Eri02l]. Ambulance [Sha01a, Swa01p]. Analog [DeC02, HBT^{+}00]. Analysis [Bar01d, BJ01b, FV00, Mit00, Ner00, Nis03e, She03]. Analytic [LMM^{+}01]. Analytical [Ber00]. Analyzing [LMM^{+}01, TCM^{+}00]. Anarchy [Bar00f, Swa02h]. Anatomy [Fre02]. Anders [Bar01b]. Android [SL01b]. Annealing [Pal01b]. anniversary [Coc02f]. Announced [Coc01a, Coc02b, Coc02i]. Announces [Bar00c]. Anonymity [Bar01b]. Anti [Sto03d]. Anti-patterns [Sto03d]. Antipasto [Swa00k]. Apache [Blo00, Jac02, Ram01]. API [Mus01, RG00, Rou02, Sto03g, Wal02a]. AppForge [Cro01b]. Applet [Bar00j]. Application [Coc02h, GKM01, HF00, Man02, SM00, TEM^{+}01, Yud02, Rem01b]. Applications [Abu01, Cox01b, Fom00a, FH02, Fox00a, HWZ00, Kis01, MKR^{+}01, PA02, Tor01]. Arc [Sha02h]. Architecture [Cro00a, Rot02b]. Architectures [Pie01]. Archives [PTB^{+}02]. Aren’t [BHP^{+}01, Coc02c, GDHI02, MKR^{+}01]. Arithmetic [Sco02]. Armies [Bar00h]. Arrays [Dem01, Ben01]. Arrogant [WNL^{+}00]. Art [Coc01a, Cro00a, PTB^{+}02]. Artistic [Swa03l]. ASP.NET [Hov02, MPG^{+}02, PA02, Rei02]. Aspect [Gro02]. Aspect-Oriented [Gro02]. AspectJ [Gro02]. Assembly [Sno01]. Assignment [Cha03]. Astronomy [Bar01d]. Atlantic [Bar00n, Coc03e]. Atom [Swa00b]. Attacks [PGH^{+}00, Swa03c]. Attention [Nis01a]. Attraction [Nis02c]. Attributes [Hov02]. Audio [Arn01, WWL^{+}02]. Augural [Wit00].

Authentication [Fre03, Str01, Coc01d, Dal01a]. Automated [Gso01a, Hav00, O’M01, Pug02, Smi03a]. Automatic [Bar00g, Fom00a, Mou00]. Automatically [BGGT01]. Automation [SL01b]. Automend [Coc02d]. AVL [Rol00a]. Award [Bar00d, Bar00i, Bar01f, Coc02b, Coc02c, Coc02i, Coc03b, Eri00a, Eri01a, Eri03a]. Awarded [Coc02k]. Awards [Ano02b]. Aware [Chr00]. Away [Bar00g, Coc02g, Coc03b]. Axis [Dou02].

B [PWS^{+}02]. Babbage [Bar00b]. Baby [Ste01l]. Back [Coc02d, EJS^{+}01, Eri01e, Ste02b, Swa01b]. Bad [BHP^{+}01, GDHI02, MKR^{+}01]. Balancing [Fab01, Mon00a]. Ballots [Coc03a]. Band [VMNH03]. Bandwagon [SSC^{+}00]. Banners [SZ01]. Bar [Swa00x]. Bargain [Rol03]. Bargain-Basement [Rol03]. Base [Coc01c, Ste02b]. Based [Bar00k, EBC01, FMPL^{+}03, Gös03, Gso01b, Kie01a, Lar01b, Li03a, Mit00, Pal01c, Pie01, Wil03i, deC01a]. Basement [Rol03]. Basic [Cro01b, Lar01a, PWS^{+}02, Swa01k]. Basics [Coc02d, EJS^{+}01]. Bass [Ste02b]. Batteries [Coc01c, Coc02d]. Batting [Bar00j]. Battlefronts [Swa03m]. BB [Coc02h]. Be [Bar00b, Eri00f, Ste02e, Coc01e]. Beams [Sha01b]. Beans [Rao02]. Beasts [Sha00a]. Beats [Bar01d, Sha01c]. Beautiful [Swa02e]. Bed [KDG^{+}00]. Beeb [Swa03h]. been [Whi01]. before [GKM01]. Beginning [GP01]. Behave [Swa00i]. BEHEMOTH [Bar00l]. Being [ASW^{+}01, Eri03f, FV01, Swa00o]. Bell [Bar00n]. Belong [Nis02b]. Benchmark [Bar01e, GKM01]. Berkeley [VR02]. Berners [Coc02k]. Berners-Lee [Coc02k]. Best [Kap00, Ste01j]. Bet [Bar00e]. Better [Coc01c, Coc02d, Eri00f, Swa00d]. BetterBASIC [ASW^{+}01]. Between [Woe03]. Bidness [Eri03c]. Bifurcations [Sha02d]. Big [BW00, Swa01t, Swa02w, Swa03g]. Big-Endian [Swa01t]. Bill [Bar00m]. Billionth [Coc02g]. Binary [Bar01f, Ben00]. Binding [Bro03, FMPL^{+}03, GLFO^{+}03]. Biometric [Coc03d, Dal01a]. Biometrics [Yud02]. Birth [Coc02c]. Birthday [Coc03g]. Bit [Gut00, Bar00b, HBT^{+}00, KDG^{+}00]. Bits [Mit02a, Rob01b]. Blass [Bar00m]. Blip [Swa00o]. Bloch [Bar00g]. Block [Coc02i, FMPL^{+}03, TCM^{+}00]. Block-Structured [FMPL^{+}03]. Blood [Nis03g, Sha00b]. Bluetooth [SG01, WK00b, WK00a]. Board [Bar01d]. Bob [Bet02]. Boneyard [Coc02j]. Book [CDsJ^{+}00]. Books [Ste00h, Wil03e]. Bookshelf [Cro00a, Cro01a, GS00, GP01, Gri01, Gri02, LS01, Pap02, Rei00b, Rei01a, Rot00, Rot01b, Rot01a, Rot02a, Sta00b, SG01, Tay00, Tay02, Wil00b, Wil00c, Wil00d, Wil00e, Wil00f, Wil01c, Wil01d, Wil01e, Wil02a, Wil02b, Wil02c, Wil02d, Wil02e, Wil03c, Wil03e, Wil03d, Wil03g, Wil03f, Wil03h, WG00, Woe01, Woe02b, Woe02c, Woe02a, Woe03]. Boomin [Nis03b]. Boost [LMM^{+}01, LMM^{+}01]. Boot [Coc02l]. Booty [Coc02d]. Borders [Coc03d]. Borland [Lis02]. Bottlenecks [Pap01]. Box [Col00, Wil03b]. Boy [Swa01m]. Brackets [RSW^{+}01]. Brain [Bar00h, Bar00g, Bar01h]. Branch [Sim00b]. Breadth [Swa00c]. Breadth-First [Swa00c]. Breakfast [Swa01m]. Breakthrough [Bar01g, Coc02f]. Breed [Bar01f]. Bresenham [Rie02]. Bring [Far01]. Brings [Swa02a]. Brittle [Nis01h]. Brooks [Bar00i]. Brother [Swa03g]. BSD [ASW^{+}01]. BSP [Cun00]. Bug [Coc03g, Swa02w]. Bugs [McR02]. Build [Mar03, PH00]. Builder [Coc03d]. Building [Bar00n, Chr00, CM01, Coh00, Mar02b, Pie01, Rem01b, She01, Ste00a, deC01a]. Builds [Bar01f, Hav00]. Bull [Ste02b].

Bundles [Jac01]. Burnout [Swa01m]. Burying [Arn01]. Business [WDGO03, Bar01d, Coc02j, GLFO^{+}03, Nis02k, Eri02i]. Buy [CDsJ^{+}00, Eri00g]. Buzzes [Coc02k]. Bye [Coc01e].

C [Coc03g, RSW^{+}01, BFG01, But02, Cav02, Cha03, Coh03, Dal01b, EBC01, Fra01, HWT^{+}01, Hin02, Hum03, Mac00, MC00, Mad01, MLDP02, MKR^{+}01, MW00, Mic00, Pat01, Pon03, RCE^{+}00, Rot01b, Rot01a, Rot02a, SP00, Sha03, SL01a, Ste00a, Ste00b, Ste00d, Ste00k, Ste00e, Ste00f, Ste00j, Ste00i, Ste00h, Ste00g, Ste00l, Ste00c, Ste01d, Ste01j, Ste01b, Ste01c, Ste01e, Ste01a, Ste01f, Ste01g, Ste01h, Ste01i, Ste01l, Ste01k, Ste02b, Ste02c, Ste02d, Ste02f, Ste02g, Ste02i, Ste02h, Ste02j, Ste02k, Ste02a, Ste02e, Ste03a, Tay01, Tru01, TC00, VMNH03, Wil00e, Woe00]. C# [Bar01b, BHP^{+}01, Cro01c, Ead01, Gri02, Hum03, SG01, Ste02c, Ste02k, Wil00g]. C99 [Che02]. Cache [BTTF02, Jol01]. Caching [Ben00]. Calculated [Coc01c]. Calculates [Coc01e]. Calculator [Yam00]. Calculators [Bir00]. Call [Bar00d, PWS^{+}02]. Callbacks [Hum03]. Calling [Pon03]. Caltech [Bar00f, Coc03a]. Camera [Bar00g, Dou02]. Cameras [Rob01a]. Campaign [Coc01b]. Can [Bar00e, Bar00g, Bar01g, Cal00c, Dav01, Gso00, Jen00a, Jol01, Kie01a, Kie02a, Mos00, Sea02, Smi01b, Ste01e, Wra01]. Canadians [Swa00q]. Canceled [Coc02e]. Canonicalization [Kel03]. Can’t [Eri00g]. Capabilities [Cal00c, Van01]. Capacity [Coc01c]. Cards [Bar00m]. Care [Eri02i, Nis02k]. Career [Isr01]. Carlo [Gin00]. Carly [Swa00p]. Cat [Eri00b]. Catches [Bar01d]. Causality [Sha00c]. CE [Tim03, Van01, Pug02, SSC^{+}00, Spu00, TCM^{+}00, Tuc00, Car02a, Car02b]. Cell [RCE^{+}00, Coc02d]. Cells [MPG^{+}02, MHM^{+}02]. Cellular [Coc02b]. Center [Isr01]. Centimillionaire [Sha00d].

Century [LMM^{+}01, MKR^{+}01]. Certification [GH00]. CESH [Tuc00]. Cg [Rat02]. Chains [Coc02b]. Challenge [Coc02g]. Challenges [Bar01e]. Change [CDsJ^{+}00, Swa00d]. Changes [Swa02a]. Changing [Ste01b, Tay02]. Channel [Bar01g]. Charge [Eri02g]. Chasers [Swa01p]. Cheap [Bar00h]. Checked [Gol01]. Checker [Boe03]. Checking [SL01a, BK08]. Checkpointing [CS02]. checksum [Fre02]. Chemical [Nis02c]. Child [LMM^{+}01, Sha01d]. Childhood [Swa00e]. Chilly [Swa00u]. Chimps [Sha00e]. China [Coc02c]. Ching [Eri03e]. Chip [Bar01h]. Chips [Coc01c]. Cities [Coc00g]. City [CDsJ^{+}00]. Claims [Bar00k]. Clarification [GWT^{+}01, SPRF02]. Clark [Kim01b]. Class [BC01, BHP^{+}01, Roe00]. Classics [Wil00c]. Claude [Bar01f]. Client [KDG^{+}00]. Client/Server [KDG^{+}00]. Clients [JRFH01]. Closed [Eri01i]. Closer [Bar01g]. Closest [Mah03]. CLR [SPRF02]. Cluetrain [Swa00k]. Cluster [Kop00]. CLX [Lis02]. CMake [Mar03]. CMOS [CM01]. Cobol [Bar01f, EJS^{+}01]. Code [Ben00, BHP^{+}01, Mos00, Sea02, Wan00, BK08]. Coding [SEC01]. Cofounder [Bar01c]. Colbert [KDK00]. Colby [Coc01a]. Coli [Coc03a]. Collaborative [Fox00a]. Collection [Bur00]. College [Bar00j, Bar01d]. Collegiate [Coc02i]. Color [Bar00k, Sha02e]. com [Bar00e, Coh01, Gso00, Gun02, McM00]. Combating [Coc02b]. Come [Bar00f, Ste01l]. Coming [Coc03e]. Commerce [Sta00c, Coc03g]. common [Cro01c]. Communications [Pap01]. Community [Swa02f]. Companies [Bar00e]. Company [ASW^{+}01, Swa03g]. Comparing [Dor02]. Compiler [Car02a, Car02b, Cav02, Ste01d]. Compilers [MLDP02, Ste00h]. Compiling [TLA^{+}02, Lai01a]. Complaining [Ste00l]. Complete [Bar00b, Vli00]. Complex [Gor00]. Complexity [Coc01l, LJ02]. Component [Ber01, Hei01, Joh00, Lis02, SLL00, TEM^{+}01]. Components [Gso00, Rei02]. Compression [Bar00j, Sco02, WWL^{+}02]. Compromising [Bar00d]. Computation [Bar01f]. Computational [Get01a, Swa02l]. Computer [Bar00g, Bar00m, Bar01b, Bar01d, Coc01e, Coc02e, Coc02f, Coc03b, FV00, Fre02, HBT^{+}00, KDG^{+}00, MPG^{+}02, Swa02d, Zan01, Coc02d, Bar01c, Coc02h]. Computer-Science [Coc03b]. Computers [Coc01a, Coc02b, Coc03b, Gro01, Kop00, Roj00]. Computing [ASW^{+}01, Bar00c, Bar01d, Bar01g, Ber01, Ber00, Che02, Coc02b, Coc02c, Coc02h, CL00, Fre02, LMM^{+}01, Ree00, Swa00l, Tom02, Wil03d]. Concept [SL01a]. Concludes [Coc01b, Coc02g, Coc03e, Coc03f]. Configuration [Ale00, Sin00, Wil01b]. Conformance [MLDP02]. Connected [Swa03j]. Connecting [Coc01b]. Connection [Jen00b, Tre02b]. Connectivity [Wil03b]. Considered [Ams02]. Considering [Rob00]. Consistency [Pal01c, Rob03]. Consistency-Based [Pal01c]. console [Rem01b]. Consortium [Bar01d]. Conspiracy [Swa01g, Swa01t]. Construction [Lai01b]. Containers [BFG01, Hin02, RSW^{+}01]. Content [Mil00, Web00, Coc03e]. Contest [Bar00i, Bar00j, Bar00f, Coc01a, Coc01b, Coc02c, Coc03f]. Contests [Coc03g]. Context [TLA^{+}02]. Continuing [Coc02e]. Contra [Ste02b]. Control [Coc02j, Cox01d, Dou02, KDK00, Lar03b, NN00, SSC^{+}00, Sim00b, Smy00, Sny03, deC00a]. Controlled [Coc01a]. Controls [Lar01a]. Controversies [Bar01f]. Controversy [Coc01b]. Conversation [Kim00a]. Convex [Get01a]. Coping [Kie01b]. Copy [WDGO03, Swa02p]. CORBA [BES^{+}01, DHR^{+}01, Hou00, IEZ01, Mod01, TEM^{+}01].

CORDIC [Pas00a]. Corner

[Sha00j, Sha00k, Sha00i, Sha00b, Sha00d, Sha00g, Sha00l, Sha00f, Sha00e, Sha00a, Sha00h, Sha00c, Sha01l, Sha01i, Sha01a, Sha01e, Sha01c, Sha01f, Sha01b, Sha01g, Sha01k, Sha01d, Sha01h, Sha01j, Sha02h, Sha02f, Sha02g, Sha02k, Sha02e, Sha02j, Sha02i, Sha02b, Sha02c, Sha02d]. Coroutines [Fra01]. Correcting [Swa00r]. Correction [Ano01, Ano02a, BHP^{+}01, TLA^{+}02, KDG^{+}00, RSW^{+}01, TEM^{+}01]. Correctly [Coh02]. Correlation [Qur03]. Cosmology [Coc01e]. Costs [Coc01l]. Could [Bar00g]. Count [Nis01f]. Counting [Coh01]. courtesy [Pat01]. CPRM [GWT^{+}01]. CPUs [Cha00]. Cracks [Bar00b]. Crawler [Kar00]. Creating [Jan01b, MHG02, Pas00b]. Creation [Coc01e]. Creator [Coc01d]. Creators [Bar01g]. Cross [Fra01, Ma01, Qur03, Zei01]. Cross-Correlation [Qur03]. Cross-Platform [Fra01, Ma01, Zei01]. Crowned [Bar00j]. Cryptographers [Coc03b]. Cryptographic [Bar00b]. Cryptography [Wie00]. Crystal [Coc02b]. Cubed [Rei01a]. CueCat [HWT^{+}01]. Cup [Nis02e]. Curl [Coc01a, Müf01]. Curley [RSW^{+}01]. Curve [Bar00b]. Custom [Roe00]. Customizing [Ste00e]. CVS [Sim00b]. Cycles [Coc02l].

D

[Cro00b, Sny03, Bri02a, MPG^{+}02, PWS^{+}02]. Daemons [DiM00]. Dahl [Coc02b]. Dammit [RSW^{+}01]. Damnation [Sto03b]. Dan [Ano02b]. Dark [Swa01x, Swa02o]. DARPA [Coc01d]. Data [Bar01g, Bar01e, Bro03, Coc03a, CNS^{+}00, Cro01a, FMPL^{+}03, Fox00a, GLFO^{+}03, McC02, Rei00b, Sam03, SSC^{+}00, Smi01b, TGV^{+}01, Arn01]. Database [Bar01h, BKCD00, Coc01a, GMW00, Kie00a, Sea02, WG00, Yua02]. Databases [Cox00, Oba03, Okm02, SB02, Whi01].

Datacenter [Jol03]. day [McC02]. Days [Cro01a, Ste00d]. DB [RZ00, VR02]. DBMS [GDHI02, Whi01]. DDD [Zel01]. DDJ [PTB^{+}02]. Dead [Far01]. Deadlock [GWT^{+}01, Mou00]. Death [Swa00w]. Debate [CDsJ^{+}00]. Debates [BKCD00]. Debian [EJS^{+}01, Epp00a, Sta01b]. Debugger [Chr00]. Debugging [Cav02, GB01, HWZ00, Zel01]. Debut [Bar00m]. Decaf [Bar01e]. Deciphering [Eri02e]. decision [GKM01]. Decisions [Coc02f]. Declarative [Swa00f]. Decoder [Ale00]. Decryption [Bar00i]. DeCSS [Coc02f, Coc03f]. Deep [Coc03e]. Default [Dau01]. Defect [Smi03a]. Defective [BTTF02, Dav01]. Definition [DeC02]. Déjà [Eri01f]. Delivery [Ste01c]. Delphi [TEM^{+}01, Hei01]. Demography [Coc03b]. Department [BHP^{+}01, Coc03g, Eri01g]. Dependability [Coc02b]. Dependencies [All02]. Descriptors [Cal00a]. Desert [Sha02f]. Design [Bar00i, Bar00j, Dav00, Jan01a, KSC^{+}00, Lev00, Ner00, Nis01g, PB00, Pat01]. Designations [VMNH03]. Designing [Bar00f, LJ01, LPP01]. Desktops [Lar03a]. Desupport [DHR^{+}01]. Detecting [GWT^{+}01]. Detection [Ish01, Mou00]. Determining [BDM^{+}03, Mil02b, PTB^{+}02]. Developed [Coc02h]. Developer [Bar01d, Ste02f]. Developers [Kap00]. Developing [Coc03d, LS01, Tor01]. Development [All02, BRVW00, DH02, DeC01b, Eri00j, Gso01b, Gun01a, HF00, Mar02a, SM00, Ste01i]. Device [How01]. Devices [Gib01, GB01, Por03, Tre02b]. DHTML [BHP^{+}01, Ma01]. Diagnosing [Eth01]. Diagnosis [Pal01c]. Dialogs [Ste00e]. Dictionary [Car01]. Dies [Bar00m, Bar00b, Bar01c, Bar01f, Coc01a, Coc01d]. Difficulty [Woe02c]. Diffie [ASW^{+}01]. Digita [VF00]. Digital [Bar00j, Bar00b, Bar01g, DeC02, HBT^{+}00, Kal01, Led00, RCE^{+}00, Rob01a, Swa02g].

DigitaScript [Rob01a]. Dijkstra [Coc02g]. Dirges [Swa02o]. Disaster [Bar00e]. Discarding [MPG^{+}02, MHM^{+}02]. Disgusting [Swa00q]. Disk [Mil02a, MHM^{+}02]. Display [Spu00]. Displays [Bar00k]. Disposal [Swa02v]. Distributed [Ber01, Esc02, Fom00a, Gun01a, Kis01, Smi03b, Tom02, Tor01]. Distributing [Bar01d]. Distribution [Gem00, McC02, SA00]. DMCA [Coc03f]. DNA [Bar01f, Coc02c, Sha02h]. Do [Coh02, Cox01a, HL00, Jac01, Jen00b, Jen02b, PH00, Rao02, Rei00a, Swa02p, Win01, Yua02]. Dobb [Ano02b, Eri00a, Eri01a, Eri03a]. Documents [Nac02]. Does [Bar00d, Hag02]. Dog [Coc01c]. Domain [Bar00l, PK02]. Domain-Specific [PK02]. Don’t [Bar00k, Eri00h, Eri02l, Sto03g]. Doone [Swa00w]. Doors [Eri01i, Eri02e]. Dot [Bar00e]. Dot-com [Bar00e]. Dots [Coc01b]. Double [Coc02f]. Double-Gate [Coc02f]. Doubling [Wil02c]. Doug [Kim00a]. Down [Coc02k, Coc02j, McR02, Wil02c]. Dr [Eri03a]. Dr. [Ano02b, Eri00a, Eri01a, Sha00j, Sha00k, Sha00i, Sha00b, Sha00d, Sha00g, Sha00l, Sha00f, Sha00e, Sha00a, Sha00h, Sha00c, Sha01l, Sha01i, Sha01a, Sha01e, Sha01c, Sha01f, Sha01b, Sha01g, Sha01k, Sha01d, Sha01h, Sha01j, Sha02h, Sha02f, Sha02g, Sha02k, Sha02e, Sha02j, Sha02i, Sha02b, Sha02c, Sha02d]. Drafted [Coc03e]. Dream [Eri01m]. Dries [Bar00g]. Driven [Hou01]. Driver [Rao02]. Drivers [Spu00, Ste01h]. Drop [Nis02b]. Duelists [Sha02g]. Dumb [BHP^{+}01, Eri01g, WNL^{+}00]. Duo [Rot01a]. Dynamic [Bar03, Joh00, Mod01, Nak01, Nis01a, Pum03, She01].

E-Address [CDsJ^{+}00]. E-Books [Ste00h]. E-Commerce [Sta00c]. E-Mail [Swa00m, Kie01b]. E-mails [Swa02q]. E-Waste [Coc02d]. E. [Coc03a]. Early [ASW^{+}01, Eri02f]. Eating [Swa02r]. eBook [Bai01, WWL^{+}02]. EC [Yam00]. Ecco [Sha00j, Sha00k, Sha00i, Sha00b, Sha00d, Sha00g, Sha00l, Sha00f, Sha00e, Sha00a, Sha00h, Sha00c, Sha01l, Sha01i, Sha01a, Sha01e, Sha01c, Sha01f, Sha01b, Sha01g, Sha01k, Sha01d, Sha01h, Sha01j, Sha02h, Sha02f, Sha02g, Sha02k, Sha02e, Sha02j, Sha02i, Sha02b, Sha02c, Sha02d]. Eclipse [Coc02k]. eCos [Mas02, Tho00]. Editor [Ste01d, TCM^{+}00]. Editorial [Bar00a, Eri00f, Eri00b, Eri00c, Eri00d, Eri00e, Eri00g, Eri00h, Eri00i, Eri00j, Eri00k, Eri00l, Eri00m, Eri01h, Eri01e, Eri01g, Eri01f, Eri01d, Eri01i, Eri01j, Eri01k, Eri01c, Eri01l, Eri01m, Eri01b, Eri02c, Eri02e, Eri02f, Eri02g, Eri02h, Eri02i, Eri02a, Eri02d, Eri02l, Eri02b, Eri02k, Eri02j, Eri03h, Eri03d, Eri03c, Eri03f, Eri03e, Eri03b, Eri03g]. Editors [Ste01h]. Edsger [Coc02g]. Efficiency [TC00]. Efficient [Kri01, Man00]. EJB [GKM01, LPP01, Rao02, TEM^{+}01]. Elections [Nis01b]. Electronic [Bar00m, Bar01e, SA00, Eri01l]. Elements [Ano02a, Ple02a]. Eliminate [Bar01d]. Ellipses [Ste00j]. Elliptic [Bar00b]. Ellis [Coc01d]. Ellison [PEYS00]. Embeddable [Jon01, Jon03]. Embedded [BES^{+}01, Cun00, DH02, FH02, Lin01, Met01, Nis01a, Nis01b, Nis01c, Nis01d, Nis01e, Nis01h, Nis01g, Nis01f, Nis01i, Nis01j, Nis02i, Nis02b, Nis02c, Nis02d, Nis02f, Nis02h, Nis02j, Nis02k, Nis02l, Nis02a, Nis02e, Nis02g, Nis03a, Nis03b, Nis03d, Nis03c, Nis03e, Nis03f, Nis03g, Okm02, RSW^{+}01, Ric00, Tho00, VC01, Wil03a, Woe01, Woe02b, Wri03, Coc03f, Coc03d, Fis02, Lar01a, Nis01j]. Embrace [Eri00c]. Embracing [Swa02h]. Emitting [Bar00k]. Encoding [Bar00m]. Encrypt [BTTF02, Dav01]. Encryption [Bar00l, Coc02k, DR01]. End [Swa00e]. Endian [Swa01t]. Endoscope [Bar00g]. Ends [DHE03]. Endures [BDM^{+}03, Nis02g]. Enduring [Swa01k].

Engelbart [Kim00a]. Engine

[Yue01, Gon01, Web00]. Engineering [Coc03a]. Engines [Kie00a]. Enhancing [Sha02a]. Enigma [Coc02h]. Enough [Bar00f, PGH^{+}00]. Entangled [Bar00l]. Entity [LPP01]. Enums [TCM^{+}00]. Environment [Müf01, Phi01, Dau01]. Environments [Dav00, Jol03, Kie02b, Tay02, Tru02]. EOF [Sco02]. Epicenter [Ste03b]. Epoch [Coc01b]. Equals [MHM^{+}02, Coh02]. Equivalence [Swa02l]. Equivalent [Coc01e]. Era [Eri01l]. ERP [Bar00e]. Error [Swa00r]. Error-Correcting [Swa00r]. ESC [Nis02d]. Escape [Sha00f]. Essays [Swa03j]. Establish [Jen00b]. Ethernet [Sta00a]. Etiquette [Coc03g]. EULA [WWL^{+}02]. Euro [Yam00]. Europeans [Bar00g]. EUVL [Bar01h]. Evangelizing [Coc01d]. Event [Bru02, Kis01]. Event-Logging [Kis01]. Events [Hou00]. Every [Swa01c]. Everyone [Coc01c]. Everything [Woe03]. Examination [Pat01]. Examining [CK00, Cro00b, Hel02, Hig01, Hig02, IEZ01, McM00, NL00, Tho02, Tim03, Van01, Van02a, Van02b]. Examples [Bar01c]. Excellence [Ano02b, Eri00a, Eri01a, Eri03a]. Exception [BJ01a, SP00]. Exceptions [Car02a, Car02b, Gol01]. Executable [Coc03f]. Executables [BHP^{+}01]. Expands [Coc02j]. Expert [Bar01f]. Exploiting [BGGT01, Gut00, HBT^{+}00, KDG^{+}00]. Exploring [Kie01c]. Explosion [Bar01g]. Expo [Swa03a]. Exposed [Ste03b]. Exposing [Wan00]. Expressions [Mad01]. Extend [Cal00c, Eri00c, Wra01]. Extended [Hov02]. Extensible [Coc03a, VF00]. Extensions [Dal01b, Rot02b, Ang01]. External [TP02]. Extinguish [Eri00c]. Extraction [Pat01]. Eye [Sha02k].

Facility [Kis01]. Fact [Ell00]. Facts [Wil03d]. Fail [Bar00k, Coc03f, She01].

Fail-Over [She01]. Failures [Bar01d, Ste01f]. Fair [Bar01c]. Fairchild [Bar01c]. Fallacies [Wil03d]. Falling [Swa02s]. Fame [Bar00l]. Famous [Eri00f]. Fantasies [Eri01k, Swa01f, Swa01g]. Farm [Bar00e]. Fast [Bar00f, Dem01, PGH^{+}00, WNL^{+}00]. Faster [Bar00l, Kie02a]. Fastest [Bar00k, Nil00]. Faux [Swa02i]. FDF [Hig01]. Fear [Swa02i]. Feature [Sto03g]. Feet [Bar00k]. Fellows [Bar01c, Coc02h]. Femtoseconds [Bar00f]. Fewer [Bar00f]. Fi [Bar00g]. Fiberoptic [Coc01c]. Fibre [Bar01g]. Fiction [Ell00, Eri01l]. Fiddle [Ste02b]. Field [Bar00f]. Fights [Coc03g]. File [Gem00, Ven00a]. Filesystem [Esc02]. Filling [Wil02d]. Filtering [Led00, RCE^{+}00]. Filters [Cal00b, Hal00, Qur03]. Fin [Sto02l]. Finally [Coc02k]. Find [PH00, Ste01e, Web00]. Find-By-Content [Web00]. Finding [Get01a, Kap00]. Finds [Bar01f]. Finite [Ano02a, Ple02a]. First [Bar00b, Ste02k, Swa00c, Coc02f]. Five [Swa01c, Swa02q, Wil03e]. Five-Sevenths [Swa01c]. Fix [GMTA02, TEM^{+}01]. Fixed [Qur01]. Fixed-Point [Qur01]. Fjords [Coc01e]. Flames [Swa00n, Swa00o, Swa00p, Swa00w, Swa00q, Swa00r, Swa00s, Swa00t, Swa00u, Swa00x, Swa00v, Swa00y, Swa01m, Swa01n, Swa01o, Swa01t, Swa01p, Swa01u, Swa01q, Swa01r, Swa01s, Swa01v, Swa01x, Swa01w, Swa02m, Swa02o, Swa02p, Swa02r, Swa02s, Swa02v, Swa02n, Swa02w, Swa02t, Swa02u, Swa02q, Swa02x, Swa03i, Swa03l, Swa03g, Swa03h, Swa03j, Swa03k, Swa03m]. Flapper [Swa01v]. Flaw [Bar01f, Swa01n]. Flexible [Sav00]. Flyby [KSC^{+}00]. Flyer [Wil00b]. Folding [Swa00d]. Foo [Swa00x]. Forecasting [WWL^{+}02]. Forensic [FV00]. Form [Bar00g]. Formal [Swa00u]. Formalized [Coc01a]. Forms [RZ00]. Fortran [Coc03e, Nag01, Nag02]. Fortune [Swa02q]. Forums [Bar00m]. Forward [Bar00k]. Found [Bar00b]. Founder [Bar00b]. Fourth [Coc02l]. Foxy [Sha01e]. Frame [Sta00b]. Framework [Coh00, DHR^{+}01, Lis02, Rei01b, Zei01, dIJ02, HF00, Tre02b]. Fred [Bar00i]. Free [Bar00f, Coc01d, BES^{+}01]. FreeBSD [Coc01d]. Freedom [Bar01e]. Freenet [Bar00f]. Frenzy [GWT^{+}01]. Frequent [Wil00b]. Fritz [Coc03e]. FrontPage [WWL^{+}02]. Fthreads [Nag01]. fuel [Coc02d]. fully [CM01]. Function [Tru01]. Functions [Get01b, Mit02b]. Funds [Bar01g, Coc01d]. Future [ASW^{+}01, Bar00d, Bar01g, Eri00l, Eri02g, GDHI02, Kim00b, Sch00, Swa01b, Whi01]. Fuzzy [Dor02].

G [Coc03b, Bar00d, LJ01]. Game [Cro00b, DHR^{+}01, Eri02k]. Games [Coc02c, Coc02l, Eri00b]. Gaps [Wil02d]. Garbage [Bur00]. Gate [Coc02f]. Gates [Bar00m]. Gateway [deC00a]. GCC [BHP^{+}01, Ste01g]. Geek [Swa01t]. Gene [Wil00f]. Generalized [Hum03]. Generate [Bar01h, Sea02]. Generating [Get01b, Jen02a]. Generator [Bri02b, Zyz03]. Generic [Ana00, Wad00a, Tor01, SLL00, Tre02b]. Genetic [Bar01g]. Geometry [Bar00j, Get01a]. German [Bar01g]. Get [Coc01d, Coc02i, Eri02l, Nis02l, Ste01d, Wil02e]. Gets [Bar00d]. Getting [Coc03a, Eri00d, Eri02g, RCE^{+}00, Swa01o]. Gifts [Swa01l]. Gigabit [Sta00a]. Girls [Swa01t]. Giving [LMM^{+}01]. GJ [Wad00a]. Globalized [PA02]. Glow [Coc02c]. GNOME [GWT^{+}01]. GNU [Coc01d, GS00]. Go [Bar01f]. GoAhead [Mas02]. Goes [Bar00e, Bar01h]. Going [Bar01h, Ste00b, Whi01]. Goldberg [Ano02b]. Gonna [Nis02l]. Good [Mit02b, Ste01e, Swa00d, Wil02c, Pat01]. Government [Bar01h, Coc03c]. Grant [TCM^{+}00]. Granting [TCM^{+}00]. Graph [Kel03, Rol02, SLL00]. Graphics [Abr00, Coc02l, Fom00b, HBT^{+}00, Bar00c]. Graphing [Bir00]. Graphs [Wal02b]. Green [Bar00m]. Grid [Coc02c]. Grinich [Bar01c]. Group [Gon00]. Guarantee [Hag02, Sto03g]. Guest [Bar00m]. GUI [Bri02b, Coc02j, Lai01b]. Guidelines [Jan01a]. Guild [Coc01b]. GUIs [Sny03]. Gumbie [Bri02b]. Gutenberg [Fis00].

Habit [Swa02t]. Hack [Swa00s]. Hacking [Coc02c]. Hall [Bar00l]. Handle [Bar01g, Cox01a]. handlers [Lin01]. Handling [BJ01a, SP00, Sco02]. Happy [Coc03g, VMNH03]. Harassment [TCM^{+}00]. Hard [Nis01h]. Hardware [Wil03a]. Harmful [Ams02, RCE^{+}00]. Hash [Get01b, Mit02b]. Headers [Tay01]. Healthcare [BTTF02, Eri01b]. Heap [Gab03]. Heart [Coc03c]. Heat [PWS^{+}02]. Hedy [Bar00h]. Hellman [ASW^{+}01]. Here [Bar00f]. Herriot [Coc03b]. Heuristic [Pal01b]. Hey [Eri00e]. High [Bar01c, Bar01h, BRVW00, Che01, Coc02b, DeC02, Eri03h, Wie00]. High-Definition [DeC02]. High-Dependability [Coc02b]. High-Level [BRVW00]. High-Performance [Che01]. High-Speed [Wie00]. High-Tech [Bar01h]. History [Bar01c, FMPL^{+}03, HBT^{+}00, Swa03d]. Hogging [Bar01b]. Holding [Swa03g]. Holzmann [Coc02c]. Home [MN00]. Honors [Coc03d]. Hood [Eri00k]. Hopes [Bar01d]. Horizon [Coc02k]. Hosted [Abu01, MKR^{+}01]. Hot [Lee01]. House [Ste01i]. HTML [Jan01b]. HTTP [Jon01, Zha00]. Hubbub [Swa02g]. Huffman [Bar00m]. Huge [BHP^{+}01]. Hull [Get01a]. Human [Bar00h]. Hungarians [Ste02d]. Hurd [Epp00a, LMM^{+}01]. Hussein [Swa01u].

IA [Jar01]. IA-64 [Jar01]. IBM [Bar00k, Coc02d, Coc03c, Woe00]. ICANN [Bar01e]. ICFP [Coc03e, Coc01b]. ICs [Bar00l, Bar00f]. Ideas [Bar00g, BHP^{+}01, Eri01g]. Identification [Smi03a, Dal01a]. Identifiers [Coc03a, RCE^{+}00]. Idle [Sto02k]. IEEE [Ale00]. If [Ste02e]. II [Ang00b, Car02b, WK00b]. IIS [Hal00]. I’m [Ste01g]. Image [Rie02, SEC01, Wit00]. Imaging [VF00]. Immersive [Coc01d]. Implement [Coh02, Gso00]. Implementation [DNM00]. implemented [Tor01]. Implementing [Hin02, KD02, Rou02]. Implicit [BGGT01]. Improving [Nak01]. IMSAI [Coc02j]. Incorrect [GOLD03]. Index [Mil00]. Indexing [Gun01b]. Industry [Bar00c, Bar01b]. Infantile [Bar00h]. Inferno [SM00]. Information [Bar01g, MPG^{+}02, MHM^{+}02, Sno01, Arn01]. Infrared [Smy00]. Infrastructure [Coc01e]. Ingalls [Ano02b]. Ingestible [Bar00g]. Initiative [Coc01d]. Ink [Bar00g]. Innovations [Sco00]. Insider [EJS^{+}01]. Install [Swa01c]. Installation [Tru02]. InstallShield [Ste02f]. Instant [Coc02l]. Instrument [Cox01d]. Integer [BK08]. Integrating [CM01, Cox00, Mas02]. Integrity [Eri01h]. Intel [Bar00c, BHP^{+}01, Coc02f, Nel02c]. Intelligent [Nel02d, Bar00g]. Interactive [Tra00, Müf01]. Interest [Ano02d, BL00a, BL00b, BL00c, BL00d, BL00e, BL00f, Coc00a, Coc00b, Coc00c, Coc00d, Coc00e, Coc00f, Coc01f, Coc01g, Coc01h, Coc01i, Coc01j, Coc01k, CS01a, CS01b, Coc03h, Coc03i, Coc03j, Son01a, Son01b, Son01c, Son01d, Son02a, Son02b, Son02c, Son02d, Son02e, Son02f, Son02g, Son02h, Son02i, Son02j, Son02k, Son03a, Son03b, Son03c, Son03d, Swa00h]. Interesting [Swa03e]. Interface [KSC^{+}00, Wil00a]. Interfaces [Alb03, Jan01b, O’B03, Sco00]. International [Coc03g].

Internationalization [Jac01]. Internet [Bar00h, Bar00k, Bar00f, Bar01b, Bar01e, Coc03b, Coc03a, Hou01, Müf01, Rit00, SZ01, TMVTdM02, Wil00a, Wil03b]. Interoperability [Bar00c, BES^{+}01, DHR^{+}01, IEZ01, TEM^{+}01]. Interrupt [Mas01, Pot01, VC01, Lin01]. Intersection [HM00]. Interview [Byk00]. Intrinsic [Woe02c]. Introduction [FV00]. Intrusion [FV01]. Invasion [ASW^{+}01]. Invention [Swa00s]. Inventor [Bar00m]. Inventors [Bar00l]. Inventory [Bar00g]. IOS [Esc02]. IP [KSC^{+}00, Räi01]. Iron [BW00]. ISAPI [Hal00, Man02]. ish [Sto02l]. ISO [MLDP02]. Issues [Coc03g]. Iterating [Ste00d]. Iteration [Qur01]. Iterator [Ana00]. Iterators [Kis00].

J.A.D.E. [Dau01]. J2EE [BTTF02, Jol01, Wal03a]. J2Me [Wri03]. Jack [Coc03b, Ste02g]. James [Kim01b]. Japan [Coc02k]. Java [Wal02a, Alb03, Ams00, Ams02, Ang01, BK08, Bar00j, Bar01b, Bar01d, Bar03, BC01, BHP^{+}01, Bru02, Cal00c, Coc02e, Coh02, Cox01a, Cox01b, Dau01, DHR^{+}01, Dwe00a, Dwe00b, Ead01, EKM00, Fox00a, Fox00b, Gib01, Gol01, GH00, Gso00, Hag02, Har00, HL00, Jac01, Jen00a, Jen00b, Jen02b, Jol01, Jon02a, Kan02, Kie01a, Kie02a, Lai01c, Li03a, Mar02b, McK01, Mos00, Nis02f, Nis02e, PH00, Pon03, Rao02, Rei00a, Rem01b, Roe00, Rot02b, Sea02, She01, Smi01b, Str01, Swa01a, TGV^{+}01, Tor01, Tra00, Tre02a, Tre02b, TCM^{+}00, Vil00, Wad00a, Whi03, Wil01a, Win01, Wra01, Yua02, dD01]. Java-Based [Li03a]. JavaScript [Jen02a, Joh00, McK01, TEM^{+}01]. JavaServer [Wal03b]. javax.crypto [Win01]. JComboBox [Wra01]. JDBC [Kie01a, Tan01]. JDBC-Based [Kie01a]. JDSL [TGV^{+}01]. JEFF [Coc02e]. Jelly [Gös03]. Jerry [Swa03h]. Jet [PEYS00]. Jim [Coc01d]. Jini [Pie01]. JJ [EKM00].

JMS [RG00, Rou02]. Job [Byk00, Sto03c]. Jobs [Bar01h, Lee01]. Johan [Coc02b]. John [Coc03b]. Johnny [Ste03b]. Joins [Bar00l]. Journalism [Swa00r]. JPEG [Nel02c, PTB^{+}02, SEC01]. Jukebox [MN00, Nel02d]. June [Ste02e]. JUnit [Alb03, Goe01]. Just [Bar01b, Coc02c, Swa03f]. JXTA [Li03b]. Jython [Bri02b].

Ka-Ching [Eri03e]. Katz [Bar00b]. KDE [EJS^{+}01, GWT^{+}01]. Keep [Bar01h]. Kemeny [Swa01h]. Kenneth [Coc01a]. Kerberos [Coc01d, Rub00]. Key [Bar00f, Bar00b]. keyword [VMNH03]. Kicking [Swa02t]. Killer [Bar01b, McR02]. King [Bar00j]. Knowledge [Eri02e]. Knows [Coc01c]. Konrad [Roj00]. Kramnik [Coc03e]. Kristen [Coc02b]. Kryptos [CNS^{+}00]. Kurtz [Swa01h]. Kurzweil [Bar00h]. Kylix [BJ01a].

Lab [Gol00]. Labor [Ell00, TCM^{+}00]. Laboratory [Coc02d]. Lakos [Pat01]. Lamarr [Bar00h]. Landing [Coc01b]. Language [Bri02a, Coc02h, Coc02l, Gös03, KSC^{+}00, MLDP02, McK01, MPG^{+}02, PWS^{+}02, Rat02, Rob01a, RWD02, Sas00, Ste00k, Wil00g, Pal01a, ZBKO01]. Languages [FMPL^{+}03, GOLD03, Kim01b, Kim02, MHG02, MHM^{+}02, PTB^{+}02, Wad00b, Cro01c]. LANs [Ste02d]. Larks [Ste00j]. Larry [PEYS00]. Last [Swa01r, Swa03a]. Latencies [Mas01, VC01]. latency [Lin01]. Latest [Sto03g]. Lattice [BJ01b]. Launched [Bar00i, Coc02b]. Launches [Coc01b]. Law [Bar00m, Wil03f, Bar01h]. Lawrence [Coc02d]. Layer [LPP01, Ram01]. Layout [Keo00]. LDAP [Cal00b]. Lead [Eri00d, RCE^{+}00, Coc02d]. Leads [Bar00f]. Leak [Ish01]. Leaks [HL00]. Leaves [Ste00j]. Lee [Coc02k]. Left [RSW^{+}01, Smi01a]. Legacy [BHP^{+}01].

Legal [Coc02f]. Legend [Eri03b]. Leighton [Rub00]. Leighton-Micali [Rub00]. Lemon [Bar00m]. Leo [Coc02l]. Leonard [Coc03b]. Lesson [FMPL^{+}03]. Lessons [Swa03d]. Let [Nis01c]. Letter [Eri01d, VMNH03]. Letters [ASW^{+}01, Ano00, Ano01, Ano02c, WDGO03, BKCD00, BES^{+}01, BDM^{+}03, BHP^{+}01, BTTF02, CDsJ^{+}00, CNS^{+}00, DHE03, DHR^{+}01, TLA^{+}02, EJS^{+}01, Eri01d, FMPL^{+}03, GLFO^{+}03, GDHI02, GF00, GOLD03, GWT^{+}01, GMTA02, HBT^{+}00, HWT^{+}01, KSC^{+}00, Kis00, KDG^{+}00, LMM^{+}01, MKR^{+}01, Mas01, MPG^{+}02, MHM^{+}02, PEYS00, PWS^{+}02, PGH^{+}00, PTB^{+}02, Rah01, RSW^{+}01, RCE^{+}00, RWD02, SSC^{+}00, SPRF02, Sta01b, TEM^{+}01, TCM^{+}00, VMNH03, WNL^{+}00, WWL^{+}02]. Level [BRVW00, Bui02]. Leveson [Bar00d]. Libraries [BHP^{+}01, GDHI02, Kie01c, MKR^{+}01, MHG02]. Library [Bra00, But02, Cun00, Nel02c, Pon03, TGV^{+}01, Woe00, SLL00]. License [Swa03l]. Lies [Swa01i]. Life [Nis01d]. Lifecycles [Ber01]. Lifetime [Coc01d]. Lifts [Bar00k]. Light [Bar00k, Coc02b]. Light-Emitting [Bar00k]. Lightweight [Jon03, Kim02, MHM^{+}02, RWD02, Tay02]. Like [Bar00h, Coc02f]. Limbo [SM00]. Line [Sno01]. Linear [Bar01d]. Lint [Zyz03]. Linux [Gri01, Ale00, ASW^{+}01, Bar00e, Bar01h, BJ01a, But02, Coc01c, Coc03f, Coc03d, DH02, Epp00b, Fis02, Min01, Nis01j, Nis03c, She03, Ste01d, Ste01i, Woe03, WM00]. LinuxWorld [Bar00c]. Lists [Mar00, Pig02, Sip00]. Literate [Coc02l, Dwe00a]. Little [GOLD03, MHM^{+}02, PTB^{+}02]. Live [Eri00f]. Livermore [Coc02d]. Living [Eri03b]. Load [Bar01a, EJS^{+}01, Fab01, Mon00a]. Loadbuilds [Smi03b]. Loader [BC01, BHP^{+}01]. Loaders [Roe00]. Loath [WDGO03]. Loathing [BDM^{+}03]. Local [Whi03]. Logging [Kis01, Mod01]. Logic [Wad00b, Wil03a]. Long [Coc02l, Ste01l]. Look [Bar00g, Ste02k]. Looking [ASW^{+}01, Swa02u]. Lore [GMW00]. Lost [Coc01b, Swa00j]. lot [Cro01c]. Love [Eri00g]. Lovelace [Coc02i]. Ltd. [Gab03]. Lyin’ [Swa01s].

Machine [GOLD03, Nis02j, Swa02g]. Machines [O’B03]. MacOS [Web00]. MacroSafe [SPRF02]. MACtimes [Far00]. Macworld [Swa03a]. Made [KDG^{+}00]. Madness [RCE^{+}00]. Magnetic [Bar00f]. Magnitude [Man00]. Mail [Bar01e, Swa00m, Kie01b]. Mail4Me [Ple02b]. mails [Swa02q]. Make [Bar00f, Kie02a, Sto03g]. Making [Dal01b, Müf01, GKM01]. Man [Swa03i]. Man-Month [Swa03i]. Manage [Jol01]. Management [Bar00n, GMW00, Keo00, Rei00a, Sim00b, Wil01b]. Manager [Mar03]. Managing [Gso01b, Kop00, Por03]. Manning [Wil03h]. Many [Mis00, Swa02a, Cro01c]. Map [Lev00, Wil02b, Yua02]. Mapping [Sav02]. Maps [Bug03]. Margination [Fis00, GF00]. Mark [Coc01a]. Market [Bar00b]. Markup [Kim01b]. Martin [Pat01]. Mass [Coc01c]. Masses [Pap02]. Matched [Qur03]. Matching [BJ01a, Dwe00b]. Math [Swa01d, Swa02e, Swa03i]. Mathematics [Rot02a]. Matters [Swa01j, Wil00d]. Mau [Swa01p]. Mau-mauing [Swa01p]. mauing [Swa01p]. Maximum [Coc01c]. May [Bar00m, Bar00f, Bar01h]. McConnell [Vli00]. McMaster [Bar00j]. Me [Bar00k, Eri00g, Ste01g, Swa01r]. Mean [Bar00l]. Meaning [Swa01k]. means [Whi01]. Measure [Mos00]. Measuring [TC00, Yun00]. Media [deC00a]. Meet [Bir00]. Meets [Bar00m, BW00, Bet02, McM00, Nis01j]. Membership [Coc01b]. Meme [Swa01w]. Memory [Bar00g, HL00, Ish01, Jol01, Mil02a, Nak01, Rob03]. MEMS [Coc02d]. Menuet [Tur01]. Merged [Coc03c]. Message [Coc02k, CNS^{+}00, Hum01, Rao02]. Message-Driver [Rao02]. Meta [Coc02d]. Method [Coh02, Sha02a]. Methodology [PB00]. Methods [Gin00, Pas00a]. Metis [PB00]. Mexican [Bar01h]. Micali [Rub00]. Microarchitecture [Bar00c]. Microcontrollers [Pas00a]. Microsoft [Bar01e, Coc03g, KDG^{+}00, McM00, Rei01a, WNL^{+}00]. Microvirus [Sha00g]. Midstream [Ste01b]. Migrating [Sut00]. Miles [Wil00b]. Milestone [Coc02c]. Million [Swa02q]. MIME [Dav01]. Mind [Bar01e]. Mingw32 [PGH^{+}00]. Minimal [WM00]. Minimizing [Lin01]. MiniRTL [WM00]. Minor [Swa03j]. Mint [Sha00h]. Minus [Ste01c, Ste01f]. Misinformation [Swa00g]. Mistakes [Bar00f]. Mithra [Fre03, GLFO^{+}03]. Mix [Nel02a, Nis02f]. Mixin [EBC01]. Mixin-Based [EBC01]. Mo [Coc01c]. Mobile [Coc02j, Jon02a, Kou01, Tre02b]. Model [Bar00m, Boe03, Coc01c, LPP01, Tor01]. Modeling [Pap02, PK02]. models [CM01]. Modem [Bar00k, Mis00]. Modern [Cha00]. Module [Jan01a]. Modules [Jac02]. Moka [dD01]. Molecular [Ree00]. Mondrian [SPM02]. Money [Eri00g, Swa00d]. Monitor [Bar00j]. Monitoring [Cox01d, Jor01, Kis01, Lar01b, Man02, McC02]. Monitors [Bec01, Coc02d]. Mono [dIJ02]. Monte [Gin00]. Month [Swa03i]. Moon [Coc01b]. Moore [Bar01h]. Mostly [Swa01d]. Mothballed [Bar00l]. Motive [Bar00a]. Motorists [Gut02]. Mouse [Eri00b]. Move [Bar00c]. Moves [Bar00k]. Moving [Rob01b, Ste01a]. MP3 [Bar01g, LMM^{+}01, Li03a, Nel02d, Ste00c]. MPEG [Wal02a, Woe02a]. MPEG-4 [Wal02a, Woe02a]. MPEGlets [Wal02a]. MQSeries [Woe00]. Mr. [Eri00e]. Mrs. [Coc02h]. MSXML [TEM^{+}01, Che01, Hei01]. Multicast [Gem00]. Multilanguage [Wen00]. Multiline [Cox01a]. Multiple [MHG02, Mit02b]. Multithreaded [CS02, HWZ00, Nak01, Zha00]. Multithreading [Nag01]. Museum [Bar01c, Coc02h]. Music [CDsJ^{+}00, FH02, Li03a, Ste01c, Ste01f]. Must [Ste02e]. My [Kie01a, Kie02a, Sea02, Swa02r, Swa03l]. MySQL [Kie00a, RZ00, YY02]. Mysty [Ste01g]. Myth [WNL^{+}00].

n [Sto02l]. Nag [Swa00t]. Name [Eri02k, Eri03d, PEYS00]. Names [Coc01d, Coc02h]. Namespaces [HWT^{+}01, Min01, Sut00]. Nancy [Bar00d]. Nanoengineering [Bar00f]. Nanoseconds [Bar00f, PGH^{+}00]. Nanostucture [Bar00g]. Nanotechnology [Coc02b]. NASA [RCE^{+}00]. National [Coc02d]. Navigation [TMVTdM02]. Near [Eri00f]. Nearing [Coc02c]. Need [Coc01d]. Needed [Bar00m]. Net [Bar00j, Jen00b]. Netscape [SSC^{+}00]. Network [BC01, Coc02e, Fab01, Kar00, Lac01, Tru00, Yun00, Coc01d, Pal01a]. Networking [How01, VR02, Tre02b]. Networks [Li03b, Ric00, Sil03, deC01a]. Neural [Bar00j]. Newell [Bar00d]. NewOS [TLA^{+}02, Gei01]. News [Bar00c, Bar00d, Bar00e, Bar00h, Bar00g, Bar00i, Bar00k, Bar00j, Bar00l, Bar00m, Bar00f, Bar00b, Bar01b, Bar01c, Bar01d, Bar01g, Bar01e, Bar01f, Bar01h, Coc01a, Coc01b, Coc01c, Coc01d, Coc01e, Coc02b, Coc02c, Coc02d, Coc02e, Coc02h, Coc02f, Coc02i, Coc02g, Coc02k, Coc02l, Coc02a, Coc02j, Coc03b, Coc03c, Coc03e, Coc03f, Coc03a, Coc03d, Coc03g]. Newt [Sha02k]. Newton [Sha02a]. Night [Ven00b]. NIST [Bar00c, Coc03d]. No [Swa00u]. Noah [Sha02h]. Noises [Coc02b]. Nominates [Bar01c]. Normal [Tay01]. Nose [Swa00l]. Notes [LMM^{+}01]. Nothing [SSC^{+}00].

Notification [Mod01]. NQL [Pal01a]. NSF [Bar01g]. NT [Met01, Jen00a, Str01]. Numbers [Coc02l, Dor02, Smi02]. Numbers2 [Sha02i]. Numeric [Che02, Wil03d]. Numerical [Eri02f, WWL^{+}02]. Nuns [Swa00u]. Nuon [Bet02]. Nygaard [Coc02b].

Object [Coh01, Dav00, Gon01, Gun01a, How01, Kie00b, McK01, Vil00, BJ01a]. Object-Oriented [Dav00, How01, Kie00b, McK01]. Objects [CNS^{+}00, SSC^{+}00, Smi01b, Yua02]. Observation [Wil03g]. ODBC [Kie02b]. Odds [DHE03]. Oeuvre [Swa00h]. Off [Bar00k, Coc02k]. Official [Coc02k]. Ogg [Li03a]. Oh [Swa00i]. Okay [Wil02c]. Old [Coc02l, Wad00b, Wil00c]. Olds [Rot00]. Ole [Coc02b]. Ole-Johan [Coc02b]. Omniheurist [Sha00j, Sha00k, Sha00i, Sha00b, Sha00d, Sha00g, Sha00l, Sha00f, Sha00e, Sha00a, Sha00h, Sha00c, Sha01l, Sha01i, Sha01a, Sha01e, Sha01c, Sha01f, Sha01b, Sha01g, Sha01k, Sha01d, Sha01h, Sha01j, Sha02h, Sha02f, Sha02g, Sha02k, Sha02e, Sha02j, Sha02i, Sha02b, Sha02c, Sha02d]. One [Eri02h, Nis03f, GMTA02, RWD02]. Online [Bar00m, Bar01f, O’M01]. Only [Bar00c, BDM^{+}03, Nis02g]. Oops [SPRF02]. Open [Bar00l, Bar01d, Coc01a, Coc01e, Coc02j, Coc03c, EJS^{+}01, Eri00i, Eri01d, Eri01i, SL01b, McC02, Bar00i, BES^{+}01, BW00, CDsJ^{+}00, Eri00i, SSC^{+}00]. Open-Source [Coc03c, SL01b, McC02]. OpenCard [HF00]. OpenCV [Bra00]. OpenML [Bar01b]. Opera [Cro01a]. Operating [TLA^{+}02, Eri01j, Gei01, Lac01, Tho00, Tur01]. Operation [Man00]. Operations [BFG01, RSW^{+}01]. Optical [Bar00d, Bar01g]. Optimal [Sco02]. Optimism [Bar00d]. Optimization [Pal01b, Qur01]. Optimizing [Fis02, Fom00b, Jar01]. Opts [Bar01e].

Oracle [Kie00a]. ORB [McC02]. Order [Swa01n]. Ordering [Win00]. Oriented [Dav00, Gro02, How01, Kie00b, McK01]. Orthogonal [Sta01a]. OS-Aware [Chr00]. Osborne [Swa03b]. Other [Swa00l, Swa01k, Swa01e, Wil00e, Wil03f]. Our [Cro01a]. Outhouse [Ste02c]. Outlawing [Coc02d]. Outsourcing [WDGO03, Eri03c, GLFO^{+}03]. Overflows [BK08]. Oversampling [Led00, RCE^{+}00]. Own [KDG^{+}00].

P2P [Coc02e]. Package [Win01]. packet [deC01a]. packet-based [deC01a]. PacketCable [deC01a]. Pad [Coc02d]. Padded [MPG^{+}02, MHM^{+}02]. Page [Hov02, Lat02]. Pages [Ang00a, Ang00b, Cox00, Kan02, Wal03b, Yue01, YY02]. Pairs [Mah03]. Palm [Swa00l, Cro01b, DeC01b, Ish01, Yam00]. Panamax [Sha01f]. Parade [Swa01q]. Paradigm [Ste02c]. Paradigms [Swa00b, Swa00a, Swa00c, Swa00d, Swa00e, Swa00f, Swa00g, Swa00h, Swa00l, Swa00j, Swa00i, Swa00k, Swa01a, Swa01b, Swa01c, Swa01k, Swa01d, Swa01e, Swa01h, Swa01f, Swa01g, Swa01i, Swa01j, Swa01l, Swa02c, Swa02e, Swa02f, Swa02g, Swa02h, Swa02i, Swa02d, Swa02j, Swa02k, Swa02l, Swa02b, Swa03c, Swa03b, Swa03d, Swa03e, Swa03f]. Parallelism [BGGT01, Gut00, HBT^{+}00, KDG^{+}00, Rol03]. Parasitic [Fre02]. PARC [Coc02c, Gol00]. Parkinson [Wil03f]. Parsing [Cox01c, GWT^{+}01, Gor00]. Part [Ang00a, Ang00b, Car02a, Car02b, She03, WK00b, WK00a]. Pascal [BJ01a]. Passes [Bar00g, Coc02g, Coc03b]. Passing [TLA^{+}02, MKR^{+}01]. Past [GDHI02, Whi01]. Patent [Coc01b, RCE^{+}00]. Patenting [Coc01c]. Patents [WNL^{+}00]. Pattern [Dwe00b]. Patterns [GS00, Mac00, Woe02a, Sto03d]. Pay [Eri02l, RCE^{+}00]. Paying [Ste02h].

PC [Coc02g, Kim01a, Smy00, Zyz03]. PC-Lint [Zyz03]. PCKS#7 [Dav01]. PCs [Bir00]. Peek [Coc02d]. Peek-a-Booty [Coc02d]. Peer [Li03b]. Peer-to-Peer [Li03b]. Embedded [DH02]. Redo [Kri01]. Server [KDG^{+}00]. Subscribe [Rou02]. Tk [TLA^{+}02, Lai01a, Nel02a]. Web [Phi01]. Penguin [Nis01e, Ste01j]. Pennsylvania [Bar00m]. People [ASW^{+}01]. Perchance [Eri01m]. Perfect [Get01b]. Performance [Che01, DeC00b, Jar01, Man02, Mos00, Pap02, Yun00]. Perimeters [Sha01g]. Peripherals [Cun00]. Perl [Cro01c, TLA^{+}02, GWT^{+}01, Gun01b, Jen02a, Kie00b, Kie01c, Lai01a, Lai01b, McM00, Sno01, Yue01, YY02, GP01]. Perl/ Tk [TLA^{+}02, Lai01a]. PerlCOM [McM00]. Persistence [Gon01, Hov02]. Personal [Coc01a, Coc02b]. Personalization [Jon02b]. PGP [Dav01]. Ph.D.s [Coc03a]. Phil [Bar00b]. Phone [Coc02b]. Phones [Kou01, RCE^{+}00]. Photomosaics [TLA^{+}02, MM01]. Photons [Bar00l]. PHP [Abu01, Cav02, MKR^{+}01, RZ00]. PHPLIB [RZ00]. physical [Pat01]. Physics [CL00]. Pickle [Swa02x]. Pining [Coc01e]. Pioneer [Bar00h, Bar01f, Coc03b]. Pipers [Ste02h]. Pitch [Eri03g]. Pitfalls [Mil02a]. PKWare [Bar00b]. Place [Swa02s]. Plan [Coc02j]. Plans [Coc03c]. Plastic [Coc02d]. Platform [Coc03d, Fra01, Gri02, Ma01, Sha03, VF00, Yam00, Zei01, Ste01k]. Platforms [Swa00l]. Play [Coc02d, LMM^{+}01, Ste01g]. Player [Li03a]. Playing [Coc02c]. PlayStations [Swa01u]. Plug [DHR^{+}01, HL00, Jen02b, Ste02i]. Plug-In [Jen02b, DHR^{+}01]. Plumbers [MPG^{+}02, MHM^{+}02]. Ply [Sha01d]. Pnuts [KSC^{+}00, McC00]. PocketPyro [DeC01b]. Point [Qur01]. Pointer [KSC^{+}00, TCM^{+}00]. Pointers [Tru01]. Points [Mah03, Nis01g, Nis02h]. Polymers [Bar00k]. Ponders [Coc03c]. Pool [BTTF02, Jol01, Wil00f]. Pooling [Vil00].

Poorly [Swa03j]. POP3 [Mic00]. popular [Hig02]. Popularity [Swa01k]. Portability [MW00]. Portable [Blo00, Kis01]. Porting [Cun00, Mon00b, SSC^{+}00, TCM^{+}00]. POSIX [Coc03c]. Postman [Eri00e]. Potential [Mou00]. Poverty [Swa01q]. Powers [Coc02h]. PPC [ASW^{+}01]. Practical [DNM00]. Practice [Eri00j]. Precompiled [Tay01]. Predication [Cha00]. Predicting [Pap01]. Prediction [Eri02f]. Predicts [Bar00h]. Preferences [TCM^{+}00]. Prepare [Bar00d]. Prepared [ASW^{+}01, FV01]. Present [Eri00l, GDHI02, Whi01]. Prime [Smi02]. Principles [TLA^{+}02, Eri00j, Eri01j, Pat01]. Print [Ste00e, Sto03c]. Printer [Bar00b]. Printout [Swa02x]. Prior [Coc01a, PTB^{+}02]. Priority [Fab01, Gut02, WWL^{+}02]. Privacy [Coc01e]. Private [Lar03a, Min01, Swa01s]. Prize [Bar01d, Bar01g, Coc02k]. Problem [Bar00c, Mah03]. Problems [Eth01]. Procedural [Dav00]. Process [WDGO03, Eri03c, GLFO^{+}03, GS00, Sto02k]. Processes [Tay02]. Processor [BDM^{+}03, Mil02b, PTB^{+}02]. Producing [Coc01c]. Professional [Gri01]. Professionals [Lee01]. Profit [Bar00a]. Program [Bar00g, BRVW00, Coc02b, TC00]. Programmable [Wil03a]. Programmer [Cro00a, Cro01a, Eri03f, Esc02, GS00, GP01, Gri01, Gri02, Hel02, Hig02, LS01, Lin02, MLDP02, McM00, Pap02, Rei00b, Rei01a, Rot00, Rot01b, Rot01a, Rot02a, Sta00b, SG01, Tay00, Tay02, Tho02, Tre02b, Wil00b, Wil00c, Wil00d, Wil00e, Wil00f, Wil01c, Wil01d, Wil01e, Wil02a, Wil02b, Wil02c, Wil02d, Wil02e, Wil03c, Wil03e, Wil03d, Wil03g, Wil03f, Wil03h, WG00, Woe01, Woe02b, Woe02c, Woe02a, Woe03, Coc01b]. Programmers [Bar00d, Bar01f, Coc00g, Coc01d, Coc03g, Zan01]. Programming [ASW^{+}01, Ano02b, Bar00f, Bar01f, Bjö01, Bri02a, Bui02, Coc01a, Coc01b, Coc02c, Coc02j, Coc03b, Coc03e, Dwe00a, EBC01, Eri00a, Eri01a, Eri03a, FMPL^{+}03, Gro02, Kim00b, Kim01a, KDG^{+}00, Kou01, MHG02, Müf01, Pum03, Rat02, Rot01b, Ste00a, Ste00b, Ste00d, Ste00k, Ste00e, Ste00f, Ste00j, Ste00i, Ste00h, Ste00g, Ste00l, Ste00c, Ste01d, Ste01j, Ste01b, Ste01c, Ste01e, Ste01a, Ste01f, Ste01g, Ste01h, Ste01i, Ste01l, Ste01k, Ste02b, Ste02c, Ste02d, Ste02f, Ste02g, Ste02i, Ste02h, Ste02j, Ste02k, Ste02a, Ste02e, Ste03a, Swa00b, Swa00a, Swa00c, Swa00d, Swa00e, Swa00f, Swa00g, Swa00h, Swa00l, Swa00j, Swa00i, Swa00k, Swa01a, Swa01b, Swa01c, Swa01k, Swa01d, Swa01e, Swa01h, Swa01f, Swa01g, Swa01i, Swa01j, Swa01l, Swa02c, Swa02e, Swa02f, Swa02g, Swa02h, Swa02i, Swa02d, Swa02j]. Programming [Swa02k, Swa02l, Swa02b, Swa03c, Swa03b, Swa03d, Swa03e, Swa03f, TH01, Wen00, Wil03i, Wil00g, Woe02b, Gri01, GP01]. Programs [CS02, Jan01b, Kie01a]. Project [Bar01d, Coc01d, Coc02k, Fis00, Ple02b]. Projects [Bar01g, Eri00i, Mar02a]. Promise [Ree00]. Promotes [Bar00i]. Properties [Bar00f]. Proposals [Coc01e]. Proposed [Bar01d, Coc02h, Coc02f]. Prose [Swa03k]. Prosecutions [Coc03f]. Protect [Coc01a]. Protection [WDGO03, Bar00g]. Protocol [Fre03, JRFH01, Rub00, Sha02i, Gun01a, Led01, deC00a]. Protocols [DeC00b, SZ01]. Proxies [Bar03]. Proxy [Eth01, Kat03]. Public [Bar00l, Bar01h]. Publish [Hou00, RG00, Rou02]. Publish/ Subscribe [Rou02]. Pumps [PWS^{+}02]. Purple [Ste00j]. Push [Coc02e]. Putting [Gun01a]. Puzzles [Swa01k]. PyGtk [CK00]. PyQt [Rem01a]. Python [Ano02a, Ang00a, Ang00b, Ple02a, Rem01a, Rem01b, Tan01, Wil00e]. Pythonic [Dal01b].

Q&A [Bru02, Cal00c, Coh02, Cox01a, EKM00, Fox00b, Gol01, Gso00, Hag02, HL00, Jac01, Jen00a, Jen00b, Jen02b, Jol01, Kie01a, Kie02a, Lai01c, McK01, Mos00, PH00, Rao02, Rei00a, Sea02, Smi01b, Str01, Tra00, Vil00, Win01, Wra01, Yua02, dD01]. qmail [Kie01b]. QNX [Van02a]. QT [Ste01g, DH02]. Qt/Embedded [DH02]. Quality [Räi01]. Quantum [Bar01g, Coc01b, Coc01c, Coc03b, CL00, Gro01, Swa02d]. QuantumBasic [Swa02k]. Query [Pal01a]. Querying [Nac02]. Quest [Ste01j]. Queues [Fab01, Gut02, WWL^{+}02]. Quicksort [Pig02]. Quincy [GDHI02, Ste00a, Ste00e, Ste00f].

R. [Eri01e]. Radix [Pig02]. Rage [Ste01h]. rainy [McC02]. Rambling [Ste00i]. Ramping [Coc02k]. Random [Mit02a]. Randomized [Rol00b]. Rant [Swa00x]. Rating [Nis01h, Nis01g, Nis01f]. Read [Swa01r]. Reading [Bar00d, Woe01]. Ready [Coc02d]. Real [Coc02i, Dou02, Fom00b, FH02, GB01, Har00, Jan01a, Nis01h, Nis01g, Nis01f, Nis01i, SSC^{+}00, She03, SPRF02, Swa06, Van01, WM00, Epp00b, Led01, Nis03c, She03]. Real-Time [Dou02, Fom00b, FH02, GB01, Har00, Jan01a, Van01, WM00, She03, Epp00b, Led01, Nis03c, She03]. Realities [HWT^{+}01]. Reality [Bar01g, Coc01d, WNL^{+}00]. Realizable [Eri01k, Swa01f, Swa01g]. Really [Coc02h, Dav01]. REBOL [Esc02, Sas00, Swa00m]. Receive [Coc03b]. Receives [Coc02c]. Recipients [Coc01d]. Record [BHP^{+}01, EJS^{+}01, Mah02, Sta01b, Rah01]. Records [Eri01i]. Recovered [Coc01b]. Recovery [Ven00a]. Recursion [Bay01, RCE^{+}00, Rob00]. Recycling [Bar00m]. Red [Bar00j]. Redirection [Tru01]. Reducing [All02]. Redux [CNS^{+}00]. reengineered [Hig02]. Reference [Coh01]. References [Ams00].

Reflective [Dwe00b]. Regained [Swa00j]. Regex [BTTF02]. Registries [Tre02a]. Regular [Mad01]. Relational [Oba03]. Relations [Pat01]. Relay [Sta00b, Zha00]. Relayed [Coc03e]. Relaying [Kie01b]. Release [Bar01d, Wil01b]. Released [Bar00c, Bar00l, Bar01b, Bar01e, Coc01c, Coc03f, Coc03d, Swa03f]. Releases [Bar00l]. Reliability [Ric00]. Relief [Bar01b, Sto03b]. Religions [Wil00e]. Renewing [Swa03l]. Repairing [Coc02d]. Repairs [Ste02k]. Replace [Bar00g]. Replacing [Car01]. Report [Zyz03]. Reports [Coc03g]. Repository [Bar00i]. Request [Ram01, Swa02v]. Rescue [ASW^{+}01]. Research [Bar00a, Bar00d, Bar01f, Gol00, Sch00]. Researcher [Bar00d, Coc01a]. Researchers [Bar00h, Coc02e]. Reservoir [HM01]. Resizable [Dem01]. Resonance [Jon02b]. Resource [Coc03a, Isr01, Jac01]. Resources [Tru00]. Responsibility [PGH^{+}00]. Resurrection [Swa00w]. Retreading [Ste00f]. Return [GDHI02]. Revenge [Sto03d]. Reverse [Coc02g]. Revisited [Swa00f]. Revolution [Kim01a]. Rewriters [TP02]. RF [VMNH03]. Rich [Eri00f]. Richard [Bar00g]. Ride [Ste02f]. Riding [SSC^{+}00, Ste02j]. Rights [Bar00b]. Rijndael [DR01]. Ringside [Eri01c]. Rising [Coc01l, Nis03d]. Rivest [Coc03b]. RMI [Mar02b]. Road [Lev00, Ste01h]. Roadmap [Coc02k]. Robocopter [Bar00k]. Robot [Bar01f, Coc02j, Coc03d, KDK00]. Robotic [Bar00f, Coc03c, Sny03]. Robotics [Bar00c]. Robots [Bar00h, Coc01d, Coc03e]. Roll [Swa01h]. ROM [Ale00]. Ronald [Coc03b]. Root [Car01]. Rotating [Ste00j]. Route [Lar03b]. Router [Coc01a]. Routines [Pon03]. RPC [Jon03, Por03]. RSA [Bar00l, Kal01, Wie00]. RTOS [Van02a, Van02b]. RTTI [Car02a, Car02b]. Ruby [MKR^{+}01, TH01, Tom02]. RubyCocoa [Tho02]. Rudy [Swa00w].

Rule [Coc02b]. Rules [Sha01k]. Run [Blo00, Bur00]. Run-Time [Blo00].

S [Eri01e, Dav01, KSC^{+}00, Ste00k]. S/MIME [Dav01]. SA [Cha03]. SA-C [Cha03]. Saddam [Swa01u]. Safety [Bar00d, Hag02]. Sale [Nis02h]. Sales [Eri03g]. Salesman [Bar01e, TCM^{+}00]. Sampling [HM01, Swa03k]. Sandia [Bar00j]. Satellite [TMVTdM02]. Satellites [Bar01h]. Saving [Swa01s]. SAX2 [TEM^{+}01, Hei01, Mus01]. Says [Bar01b, Coc03d]. Scaffolding [Swa00v]. Scalability [Nak01]. Scalable [Gem00, Tor01, Bar00c]. Scaling [Coc02k, Rie02]. Scan [Bar01h]. Scarcer [Coc03a]. Scene [Wal02b]. Scenes [Nis02d]. Scheduling [Pot01, Ram01]. Schema [Coc01a]. Scholarships [Bar01c]. School [Coc02f]. Schoolers [Bar01c]. Sci [Bar00g]. Sci-Fi [Bar00g]. Science [Bar01b, Bar01d, Bar01f, Bar01h, Coc02e, Coc03b, Coc03a, Cro00a, Eri01l, KDG^{+}00, LJ02, Coc02c]. Scientists [Coc01d]. SCRAM [Wil01b]. Scraping [Kat03]. Screen [Coc02c]. Script [Lai01c]. Scripting [Gös03, McC00, Rem01b, Rob01a, Sas00, Sav00, SL01b, Swa02d, Tra00]. Scripts [TLA^{+}02, Lai01a]. Sculpture [Bar01f]. SDK [Bar00l, Jon02a]. SDL [Bjö01]. Search [Ben00, Cal00b, Cal00c, Swa00c]. Searches [DNM00]. Searching [Coc02c, Gro01, Gun01b]. Seat [Eri01c]. Second [Tay00]. Secrecy [Swa02j]. Secrets [Bar00d, Swa01i, TEM^{+}01]. Secure [Ang01, Mar02b, deC01a]. Securing [Her02, Sil03]. Security [Bai01, Bar00h, Cal00a, Coc01d, Dal01a, DeC00b, KSC^{+}00, Ner00, PWS^{+}02, Rot02b, Sch00, Swa03m, Whi02, WWL^{+}02, Yud02, deC01a]. Seek [Coc01d]. Seems [Coc02f]. Select [Joh00]. Selected [Bar00l]. Selecting [GKM01]. Selection [Mah02]. Self [Coc02d]. Self-Repairing [Coc02d]. Semantic [GDHI02, Coc01d, Swa01k]. Semantics [Tay01]. Semiconductor [Coc02k, Bar01c]. sendmail [Kie01b]. Sense [CNS^{+}00, PEYS00]. Sequence [Bar01d]. Sequences [Sta01a]. Series [Coc02j]. Server [Ang00a, Ang00b, Bar01e, Bui02, Cro01a, Eth01, Jan01b, Jon01, Jon03, Kan02, Rei01a, Rei02, Yue01, YY02, Zha00]. Servers [GKM01, Mar02b, She01, TEM^{+}01]. Service [Coc01c, Pie01, Räi01, Coc02k, Mod01]. Service-Based [Pie01]. Services [Coc01a, Coc02i, Hum01, Jen00a, Jol03, Kan02, Pas00b, Sha03, Swa00m, Wal03a]. Servlet [Hin02]. Servlets [Cox01b, Kie02a, Rei00a]. Session [Rei00a]. Set [BFG01, JRFH01, RSW^{+}01, HWT^{+}01, Sta00c]. Set-Theoretic [BFG01, RSW^{+}01]. SETI [Bar01d]. Setting [BHP^{+}01, EJS^{+}01, Eri03h, Rah01, Sta01b]. SEU [Nis01j]. Sevenths [Swa01c]. Sexy [Sto03g]. SGI [Bar00l]. Shadow [Sav02]. Shamir [Coc03b]. Shannon [Bar01f]. Shapes [Bar01f]. Share [Mis00]. Shared [BES^{+}01, BHP^{+}01, Fox00a, GDHI02, MKR^{+}01, PGH^{+}00, TEM^{+}01]. Shareware [Ste00g]. Shift [RSW^{+}01, Smi01a]. Shift-Left [RSW^{+}01, Smi01a]. Shines [Coc02k]. Shirts [Bar00j]. Short [Hum01, Coc02k]. Shortage [Ano00, CDsJ^{+}00, CNS^{+}00, Ell00, Eri00m, HBT^{+}00, KSC^{+}00]. Shortchanging [Bar01h]. Should [Bar00k, Lai01c, Mis00]. Shuffling [Rol00b]. Shut [Coc02j]. Shuttle [RCE^{+}00]. sic [RSW^{+}01]. Sickness [Nis02i, SPRF02]. Sign [BTTF02, Dav01]. Sign-and-Encrypt [BTTF02, Dav01]. Signal [She03]. Signalling [BK08]. Signals [Bar00g]. Signatures [Kal01]. Signs [Bar00j]. Simple [Sha00i, Gun01a, Mus01]. Simplicity [Kim01b]. Simplify [Smi01b]. Simplifying [Gun01a]. Simulated [Pal01b]. Simulating [Roj00]. Simulation [Bar00m, Led01]. Simulator [Nis02i, SPRF02]. Single [Cha03, Coc03c]. Singularity [PEYS00]. Site [Gun01b, Coc02f]. Sites [Bar01a, Che01, EJS^{+}01]. Sixth [CNS^{+}00, PEYS00]. Size [Swa01j]. SKILL [Nel02a]. Skinned [Swa01o]. Skinny [Wil03e]. Skip [Mar00]. Slamming [Ste00h]. Sleep [Eri01m]. Small [Dem01, Sha02i, DH02]. Smaller [Bar00l]. Smalltalk [Coc02g, PGH^{+}00]. Smart [Bar00j, Bug03, Kis00, KD02, TCM^{+}00]. SmlMC [Boe03]. Smooth [SB02]. SMS [Coc02k]. SO KEEPALIVE [Fox00b]. SOAP [Cro01a, Gun01a]. Soapboxes [Ste01b]. Sociable [Bar01f]. Socket [Ang01, Bui02, But02]. Socket-Level [Bui02]. SocketCC [MPG^{+}02]. Software [Bar00d, Bar00i, Bar00f, Coc01a, Coc01c, Coc01d, Coc01l, Coc02c, Cro00a, Eri00j, Kap00, LMM^{+}01, Mar02a, Sin00, SA00, Ste01i, Whi02, Yun00, BES^{+}01, Wil01b]. Solved [Bar00c]. Some [Wil01e]. Sometimes [Wil02e]. Sort [Gon00]. Sorting [CDsJ^{+}00, Nil00, WNL^{+}00]. Sorts [CDsJ^{+}00, Pig02]. Sought [Coc02l]. Soundtrack [Sto03e]. Source [Bar00l, Bar01d, BES^{+}01, BHP^{+}01, Coc03c, EJS^{+}01, SL01b, TEM^{+}01, McC02, Bar00i, BES^{+}01, BW00, CDsJ^{+}00, Eri00i, SSC^{+}00]. Source-Code [BHP^{+}01]. Sourced [Coc01e]. SourceMonitor [Wan00]. Space [Bar00j, BES^{+}01, Coc03d, Nis01a, Nis01b, Nis01c, Nis01d, Nis01e, Nis01h, Nis01g, Nis01f, Nis01i, Nis01j, Nis02i, Nis02b, Nis02c, Nis02d, Nis02f, Nis02h, Nis02j, Nis02k, Nis02l, Nis02a, Nis02e, Nis02g, Nis03a, Nis03b, Nis03d, Nis03c, Nis03e, Nis03f, Nis03g, RCE^{+}00]. Spatial [ZBKO01]. Speaking [Bar00e, Ste00i]. Spec [Bar00c, Bar00l, Bar01b, WK00b]. Specific [PK02]. Specification [Bar01d, Coc03c, Coc03f, Coc03d, Har00]. Specs [Nis01c, Nis01e]. Spectral [Nis03e]. Spectrum [Bar00h]. Speculation [Cha00].

Speech [Bar01e, Coc02h]. Speed [Bar00k, Kie01a, Wie00]. Spell [Bar00e]. Spider [KD02]. Spinning [Ste00j]. Sponsors [Coc02c]. Sprawl [Sha01h]. Spread [Bar00h]. Sprinklers [Sha02f]. Spy [Bar01h]. Spyware [Ste00g]. SQL [Rei01a, TP02, Yua02]. sqrt [Rah01]. Square [Car01]. SquareList [GLFO^{+}03, Sam03]. Stack [Coh03]. Stall [Swa01c]. Standard [Bar00g, Bar00b, Coc01c, SEC01, Bar00l, Bar01g, Coc02f, Coc02k, DR01, Nag02, Sta00c]. Standardization [Bar01h]. Standards [Bar00k, Coc02k, Coc02l, Coc03g, Eri03h]. Standing [TLA^{+}02]. Stanford [Coc02i]. Starlets [Sha00j]. Stars [Sha00j]. State [Bar00h, BES^{+}01, GOLD03, Kis01, Mac00, Nis02j, O’B03, Rei00a]. Static [BJ01b]. Station [Bar00j]. Stick [Ste02b]. Sticks [Sha00k]. Still [Ste00l]. Stirring [Nis02e, Wil00f]. Stob [Sto01, Sto02a, Sto02b, Sto02c, Sto02d, Sto02e, Sto02f, Sto02g, Sto02h, Sto02i, Sto02j, Sto02k, Sto02l, Sto03b, Sto03a, Sto03d, Sto03g, Sto03c, Sto03e, Sto03f]. Stoblog [Sto03f]. Stock [Bar00b]. Stone [BDM^{+}03, Nis02g, Sha02j]. Storage [Coc03a]. Store [Bar01e]. Straight [BHP^{+}01, EJS^{+}01, Sta01b, Rah01]. Strangers [Ven00b]. Strategy [Coc03c, Swa02j]. Streaming [Gso01a, Lar03b]. Strike [Eri02h, GMTA02, RWD02]. Strikes [Eri00c]. String [DNM00]. Strings [Cox01a]. Strix [Gon01]. Structure [Sam03]. Structured [FMPL^{+}03]. Structures [CNS^{+}00, Gor00, SSC^{+}00, TGV^{+}01]. Student [Coc02i]. Studio [Cro00b, Hel02, Swa02d]. Studying [Coc03f]. Subscribe [Hou00, RG00]. Subscribers [Mis00]. Successes [Ste01f]. Suffix [Ben01, DNM00]. Suffix-Tree [DNM00]. Summit [Swa02w]. Super [Bar00e, Bar00h]. Supercomputer [Bar00h, Bar00k, Bar01h, Coc01d, Coc02i]. Supercomputers [Coc03c]. Supercomputing [Coc02c]. Support [Nis01d]. Surfing [Coc01a]. Surgeons [Bar00f, Coc03c]. Surveying [Zan01]. SVG [Bug03]. Swaine [GDHI02, Swa00n, Swa00o, Swa00p, Swa00w, Swa00q, Swa00r, Swa00s, Swa00t, Swa00u, Swa00x, Swa00v, Swa00y, Swa01m, Swa01n, Swa01o, Swa01t, Swa01p, Swa01u, Swa01q, Swa01r, Swa01s, Swa01v, Swa01x, Swa01w, Swa02m, Swa02o, Swa02p, Swa02r, Swa02s, Swa02v, Swa02n, Swa02w, Swa02t, Swa02u, Swa02q, Swa02x, Swa03i, Swa03l, Swa03g, Swa03h, Swa03j, Swa03k, Swa03m]. Swarm [Coc03d]. Swing [Wra01]. Swirling [Ste00j]. Swiss [Coc03a]. Synchronization [Bec01]. Synchronized [Bay01]. Synthesis [FH02]. System [Coc02h, Gei01, GMW00, McC02, Tho00, Tur01, Coc01d, Bar00n, TLA^{+}02, Lac01]. Systems [Coc03d, Eri01j, Jan01a, LS01, LJ01, Lin01, Nis03b, RSW^{+}01, Tho00, VC01, Wil03a, Wil03i, Woe02b, Wri03, DH02].

T [Eri02a]. Tables [Mit02b, Sea02, Yua02]. Tac [Sha01j]. Tackle [Coc02e]. Tags [Coc02h]. Tail [Eri00m]. Takes [Coc02c]. Takin [Nis02k]. Taking [Eri02i]. Talent [Bar01b]. Tales [Swa03m]. Talk [Nis01c]. Talking [SB02]. Tall [Eri00m]. Tape [Coc01b, Gib01]. Target [Bar00c, Ste01a]. Tariff [Bar00n]. Taxes [BES^{+}01]. Taxidermist [Swa01v]. TCL [Lai01c, Nel02a]. Tcl/Tk [Nel02a]. TCP [Fre02]. Teacher [DHE03]. Teaching [Eri02a]. Team [Bar00j, Gso01b, Bar00j]. Team-Based [Gso01b]. Tech [Bar01h, Coc03g]. Techniques [Lev00, Ven00a]. Technology [Bar01g, Eri02a]. telephony [deC01a]. Templates [Lat02]. Tenors [Swa00y]. Tense [Eri00l]. Terascale [Coc02h]. Test [Bar01d, Coc03f, MKR^{+}01, GKM01]. Testing [Alb03, Bar01a, Col00, DiM00, EJS^{+}01, Goe01, MLDP02, Nis03f, Pug02, RSW^{+}01, SL01b, Smi01a, Whi02]. Testpoint [Coh00]. Tests [Coc02e, HM00]. Texas [Eri02j]. Text [Gor00]. Theft [Coc02b]. Theoretic [BFG01, RSW^{+}01]. Theories [Swa01g]. Theory [Jon02b]. There [Bar00i]. Those [Swa00a]. Thought [Coc01a]. Thought-Controlled [Coc01a]. Thoughts [Sto02k]. Thrashing [Mil02a, MHM^{+}02]. Thread [Hag02, Whi03]. Thread-Local [Whi03]. Threading [DHR^{+}01, Rei01b]. Threads [TLA^{+}02]. Three [Eri00c, Wil03h]. Tic [Sha01j]. Tic-Tac-Toe [Sha01j]. Tides [Nis03d]. Time [Blo00, Dou02, Fom00b, FH02, GB01, Har00, Jan01a, Nis01h, Nis01g, Nis01f, Nis01i, SSC^{+}00, Tay00, Van01, Wil01d, WM00, She03, Epp00b, Led01, Nis03c, She03]. Times [Eri02f, Swa03e]. TINI [Wil00a]. Tiniest [Bar00e]. Tiny [Bar00i, Coc02h, Coc03f, Yue01, YY02, Bar00e]. Tip [KDG^{+}00]. Toast [Eri02j]. Toe [Sha01j]. Tombs [Sha02j]. Too [Wil00b]. Tool [Goe01, Jen02b, BK08]. Toolbox [Rit00]. Toolchest [Esc02, Hel02, Hig02, Lin02, MLDP02, McM00, Tho02, Tre02b]. Toolkit [CK00, Hig01, Rem01a, Wil01a, Cro00b, Fox00a]. Tools [Bar00i, GS00, Sav00]. toolset [Hig02]. Top [Coc00g]. Totally [DHR^{+}01]. Touch [Coc03e]. Tournament [Coc02i, Coc03a]. traceable [CM01]. Traces [Coh03]. Tracking [Mar02a, McR02]. Tracks [Bar00j]. Trade [Bar00d]. Trading [O’M01]. Training [Coc02f]. Traits [MC00]. Trampolines [Lin01, RSW^{+}01]. Transistor [Bar00e, Coc02f]. Transistors [Bar00i, Coc02h]. transport [Kie01b]. Traps [Coc02b]. Trash [Bar01e]. Traveling [Bar01e, TCM^{+}00]. Traversal [Ana00]. Traversals [Rol02]. Tree [Ana00, DNM00]. Trees [Rol00a]. Triangle [HM00, KDG^{+}00].

Tricks [EJS^{+}01]. Trifles [Wil03g]. Trio [SG01]. Triumph [Kim01b]. True [PEYS00, Swa01k]. trust [Dav01]. TSR [SPRF02]. Tundra [Sha01i]. Turing [Bar00i, Coc03a, Bar01f, Coc02b, Coc03b, Coc03f]. Turns [Coc02g]. TV [DeC02]. Twelve [Ste00l]. Two [Coc02j, Dou02, Nis03f, Rot00, GMTA02]. Two-Axis [Dou02]. Two-Year [Rot00]. Type [MC00, Mit00]. Type-Based [Mit00]. typename [Ste03b, VMNH03].

U. [Eri01e]. U.S. [Bar00f, Coc03d]. UCITA [Eri00h]. UDDI [Tre02a]. uh [Ste02b]. UI [Kie00b]. Ultimate [MN00, Sha01j]. UML [Bjö01, CM01, Coc01e]. Unbiasing [Mit02a]. Undercover [Ste00b]. Underdocumented [Ste03a]. Understanding [TLA^{+}02, MM01, Wal02b]. Undo [Kri01]. Undo/Redo [Kri01]. Unexpectedly [Bar00b]. Union [TCM^{+}00]. Universal [LPP01]. UNIX [Coc01b, Coc01e, Coc03c, DiM00, Mon00a, Mon00b, Kie02b]. Until [Ste00d]. Unveiled [Bar00b]. Update [MPG^{+}02, MHM^{+}02, TEM^{+}01, TCM^{+}00]. Updates [Fom00a]. Uploads [Bar00h]. Upstarts [Coc02e]. URL [TLA^{+}02]. Use [Bar01c, Jen02b, Win01]. Usenet [Coc01d]. USENIX [Coc01d]. User [Jan01b, O’B03, Rei00a, Sco00]. Using [Blo00, Cav02, Gso00, Met01, Tru00, Whi03]. Utilities [Coc01e]. Utility [Tuc00]. Utilization [BDM^{+}03, Mil02b, PTB^{+}02].

Value [BJ01b]. Values [GOLD03].

Vanishing [Bar01f]. Variables [Whi03]. variety [GKM01]. VB [WDGO03, BDM^{+}03, DHE03, VMNH03]. VB.NET [Hig02, PWS^{+}02]. Vecotr [Bar00c]. Verifying [Tru02]. Verity [Sto01, Sto02a, Sto02b, Sto02c, Sto02d, Sto02e, Sto02f, Sto02g, Sto02h, Sto02i, Sto02j, Sto02k, Sto02l, Sto03b, Sto03a, Sto03d, Sto03g, Sto03c, Sto03e, Sto03f].

Version [NN00, SSC^{+}00, Sim00b]. Versus [Che01, DHE03, Ead01, Gun02, Kie00a, Rub00, VMNH03, WWL^{+}02, ASW^{+}01, BES^{+}01, Coc03e, GWT^{+}01, HBT^{+}00, KDG^{+}00, RCE^{+}00, WNL^{+}00]. Victor [Bar01c]. Video [Coc02l, Gso01a, Lar01b, Lar03b]. Vie [Coc03g]. View [Bar00b]. Views [Bar00c, Bar00d, Bar00e, Bar00h, Bar00g, Bar00i, Bar00k, Bar00j, Bar00l, Bar00m, Bar00f, Bar00b, Bar01b, Bar01c, Bar01d, Bar01g, Bar01e, Bar01f, Bar01h, Coc01a, Coc01b, Coc01c, Coc01d, Coc01e, Coc02b, Coc02c, Coc02d, Coc02e, Coc02h, Coc02f, Coc02i, Coc02g, Coc02k, Coc02l, Coc02a, Coc02j, Coc03b, Coc03c, Coc03e, Coc03f, Coc03a, Coc03d, Coc03g]. Virtual [BFG01, Mil02a, RSW^{+}01, Sip00]. Virtualization [Swa06]. Virus [PGH^{+}00]. Viruses [MPG^{+}02]. Visas [Bar00m]. Visio [Tru00]. Visual [Lar01a, Zel01, Coc03g, Cro01b, Hel02, Lac01, Nag01, PWS^{+}02]. Visualizing [Tru00]. Vital [Bar00j]. VMware [NL00]. VNOS [Lac01]. Voice [Hou01, Phi01, Räi01, Yud02]. Voice-Driven [Hou01]. Voice-Over-IP [Räi01]. Voice/Web [Phi01]. VoiceXML [Hou01, Phi01]. Volcano [Jor01]. Volunteers [Bar00m]. Voyage [Coc02e]. vu [Eri01f]. VxWorks [Van02b].

W3C [Bar00k, Bar01h, Coc01b]. Waba [Wil01a]. Waiting [Swa02k]. Want [Wil02e]. Wanted [Bar01c]. WAP [JRFH01, Pas00b]. War [Sha02e]. Warps [Wil01d]. WaSP [Coc02k]. Waste [Coc02d]. Watermarking [Arn01, WWL^{+}02]. WAV [Li03a]. Wave [Ste02k]. Waves [Ste02f, Ste02j]. Way [Ste01l, Swa00p]. Wayback [Swa02g]. Ways [WDGO03, Nis01f]. Wear [Swa00u]. Weather [Bar00h, Bar00m, Eri02f, McC02, WWL^{+}02]. Web [Sha03, Bar01a, Che01, Coc01d, Coc02f, Coc02i, Coc02k, Coc03g, Cox00, Cox01b, EJS^{+}01, GDHI02, Gso01b, Gun01b, Hou00, Jol03, Kan02, Kat03, Lar01b, Man02, PA02, Swa01k, Wal03a]. Web-Based [Lar01b]. Webcasters [Coc02j]. Webcasting [Eri00l]. WebDAV [Hal00]. Webrelay [Zha00]. WebServer [Mas02]. WEP [Coc01e]. We’re [RSW^{+}01]. Whatever [Ste01c, Ste01f, Ste02b]. Where [Nis02l, Whi01]. Whistling [Swa01x]. White [Col00, Ste01i]. White-Box [Col00]. Who [Sha01k, Ste02g]. Wide [Coc03g]. Wildfires [Sha01l]. Will [Coc01a, Sto03g]. Willy [Swa00u]. Win [Bar01c, Bar01g]. Win32 [Bec01, Kar00, Mon00a, Mon00b, NN00, Sip00]. window [Rem01b, Tuc00]. Windows [GP01, Bar00c, Cal00a, Car02a, Car02b, Lar03a, Met01, Mil00, Pug02, Spu00, Sto03g, Tim03, Van01]. Winds [Coc02i]. Winner [Eri02d]. Winners [Bar01b, Coc01a]. Wins [Bar00i, Bar00j, Bar01f]. Wired [Coc01e, DHR^{+}01]. Wireless [Coc01a, GB01, LS01, Sil03, Tre02b, VR02]. Without [SP00, Sto02l, Sto03b]. Woes [BTTF02, Ste01c]. Wolfram [Swa02l]. Word [LMM^{+}01]. Words [Swa02r]. Wordsnakes [Sha00l]. Work [Rao02, Ste01e, Gun01a]. Worker [Ano00, CDsJ^{+}00, CNS^{+}00, Eri00m, HBT^{+}00, KSC^{+}00]. Works [Bar01h]. Workshop [MHM^{+}02, RWD02, Kim02]. World [Bar00k, DHE03, GWT^{+}01, Ste00c, VMNH03, WG00, Coc03g]. Worlds [Swa01e, Wil01c]. Worries [SSC^{+}00]. Worth [Sim00a]. Worthless [Swa01l]. Woz [Bar00l]. Wrap [Sta00b]. Wrap-up [Sta00b]. Wrapper [Tan01]. Write [Jen00a]. Writing [Jac02, Spu00]. Written [Nis03g]. WSDL [Bar00l, Bar01h]. WSP [Sha03]. wxWindows [Zei01].

Xbox [Abr00]. Xerox [Gol00]. XLink [Bar00k]. XML [Hei01, TEM^{+}01, Bar01d, Bro03, Coc01a, Cox01c, Cox01d, Dav01, Dwe00a, Dwe00b, FMPL^{+}03, GLFO^{+}03, GWT^{+}01, GMW00, Gös03, Her02, Jon03, Kim01b, MKR^{+}01, Mus01, Nac02, Oba03, Por03, SSC^{+}00, Sin00, Wil01e, Wil03i]. XML-Based [FMPL^{+}03, Gös03, Wil03i]. XML-RPC [Jon03, Por03]. XP [Lar03a, Woe03]. XrML [Bar00b]. XScale [Bar00c]. XSLT [Nac02, VMNH03].

Y2K [SSC^{+}00, Ste00d]. Y5K [Bar00d]. YAPP [Ste01k]. Year [Rot00]. Years [Ste00l]. Yesterday [Coc02f]. Young [Coc03g]. You’re [Coc01c, Eri00c]. You’ve [Ste01l].

Zappa [Swa01w]. Zealanders [RSW^{+}01]. Zebras [Swa02u]. Zimbabwe [Sha02j]. Zombies [Swa00a]. Zones [Nis01i]. Zooming [Wit00]. Zope [Lat02]. Zurich [Ste02e]. Zuse [BHP^{+}01, KDG^{+}00, Roj00].

References

Abrash:2000:IXG

[Abr00] Michael Abrash. Inside Xbox graphics. Dr. Dobb’s Journal of Software Tools, 25(8):21–22, 24–26, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Abualsamid:2001:PHA

[Abu01] Ahmad Abualsamid. PHP and hosted applications. Dr. Dobb’s Journal of Software Tools, 26(1):56, 58, 60–63, January 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_01/php.txt.

Albrecht:2003:TJI

[Alb03] Matt Albrecht. Testing Java interfaces with JUnit. Dr.

Dobb’s Journal of Software Tools, 28(2):24, 26, 28, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7790/ddj0302c/.

Alexander:2000:LIC

[Ale00] William F. Alexander. A Linux IEEE 1394 configuration ROM decoder. Dr. Dobb’s Journal of Software Tools, 25(8): 80, 82, 84, August 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_08/
linx1394.txt;   http://www.
ddj.com/ftp/2000/2000_08/
linx1394.zip.

Allen:2002:RDN

[All02] Scott Allen. Reducing dependencies in .Net development. Dr. Dobb’s Journal of Software Tools, 27(1): 36–40, January 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_01/
gdepends.txt;   http://www.
ddj.com/ftp/2002/2002_01/
gdepends.zip.

Amsterdam:2000:JR

[Ams00] Jonathan Amsterdam. Java references. Dr. Dobb’s Journal of Software Tools, 25(2):42, 44, 46, 48, 50, 52, February 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_02/
jrefer.txt.

Amsterdam:2002:JNC

[Ams02] Jonathan Amsterdam. Java’s new considered harmful. Dr. Dobb’s Journal of Software Tools, 27(4):19–20, 22, 24, 26, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Ananiev:2000:AAG

[Ana00] Alexander Ananiev. Algorithm alley: A generic iterator for tree traversal. Dr. Dobb’s Journal of Software Tools, 25(11):149–152, 154, November 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_11/aa0011.
txt.

Angell:2000:PSPa

[Ang00a] Kirby W. Angell. Python server pages: Part 1. Dr. Dobb’s Journal of Software Tools, 25(1):44, 46–47, 50, January 2000. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_01/
psp1.txt;   http://www.ddj.
com/ftp/2000/2000_01/psp100.
zip.

Angell:2000:PSPb

[Ang00b] Kirby W. Angell. Python server pages: Part II. Dr. Dobb’s Journal of Software Tools, 25(2):54, 57–61, February 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ftp/2000/2000_
02/psp2.txt.

Angell:2001:JSS

[Ang01] Kirby W. Angell. The Java Secure Socket Extensions. Dr. Dobb’s Journal of Software Tools, 26(2):21–22, 24, 26, 28, February 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_02/
                                        jsse.txt;   http://www.ddj.
                                        com/ftp/2001/2001_02/jsse.
                                        zip.

Anonymous:2000:LWS

[Ano00] Anonymous. Letters: Worker shortage. Dr. Dobb’s Journal of Software Tools, 25(9): 12, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:2001:LC

[Ano01] Anonymous. Letters: Correction. Dr. Dobb’s Journal of Software Tools, 26(4):12, April 2001. CODEN DDJOEB. ISSN 1044-789X. See [Pal01c].

Anonymous:2002:CPF

[Ano02a]  Anonymous.  Correction:

“Python and Finite Elements”. Dr. Dobb’s Journal of Software Tools, 27(4–4):10, April 2002. CODEN DDJOEB. ISSN 1044-789X. See [Ple02a].

Anonymous:2002:DDE

[Ano02b] Anonymous. Dr. Dobb’s Excellence in Programming Awards: Adele Goldberg and Dan Ingalls. Dr. Dobb’s Journal of Software Tools, 27(5):18–19, May 2002. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:2002:Lb

[Ano02c] Anonymous. Letters. Dr.

Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Anonymous:2002:I

[Ano02d] Anonymous. Of interest. Dr.

Dobb’s Journal of Software Tools, 27(12):83, December 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7718/ddj0212o/.

Arnold:2001:AW

[Arn01] Michael Arnold. Audio watermarking: Burying information in the data. Dr. Dobb’s Journal of Software Tools, 26(11):21–22, 24–26, 28, November 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_11/watermk.
txt;   http://www.ddj.com/
ftp/2001/2001_11/watermk.
zip.

Allison:2001:LLE

[ASW^{+}01] Dennis Allison, Randy Schrickel, Reid Womack, Jeremy C. Reed, Ashley Tate, and Paul Munsey. Letters: Looking for early PPC [People’s Computing Company] people; being prepared for invasion; BetterBASIC; Linux versus BSD; Diffie–Hellman to the rescue; the future of programming. Dr. Dobb’s Journal of Software Tools, 26(6):10, 12, June 2001. CODEN DDJOEB. ISSN 1044-789X.

Bailey:2001:IES

[Bai01] Daniel V. Bailey. Inside eBook security. Dr. Dobb’s Journal of Software Tools, 26(11):42, 44– 45, November 2001. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:ERP

[Bar00a] Nicholas Baran. Editorial: Research and the profit motive. Dr. Dobb’s Journal of Software Tools, 25(12 (supplement)):4, December 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVg

[Bar00b] Nicholas Baran. News and views: 108-bit elliptic curve cryptographic key found; new algorithm cracks the stock market; first complete Babbage printer unveiled; XrML view to be digital rights standard; PKWare founder [phil katz] dies unexpectedly. Dr. Dobb’s Journal of Software Tools, 25(7):18, July 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVC

[Bar00c] Nicholas Baran. News and views: Computing problem solved; Scalable Vecotr Graphics spec released; Intel announces XScale microarchitecture; Windows not the only target at LinuxWorld; NIST, robotics industry move towards interoperability. Dr.

Dobb’s Journal of Software Tools, 25(11):18, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVD

[Bar00d] Nicholas Baran. News and views: Does “reading” have a future?; trade secrets compromising academic research?; optical optimism; call to all programmers: Prepare for Y5K; software safety researcher [nancy G. leveson] gets ACM’s Newell Award. Dr. Dobb’s Journal of Software Tools, 25(5):18, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVE

[Bar00e] Nicholas Baran. News and views: ERP can spell disaster; Linux goes super; dot-com companies bet the farm; the tiniest transistor; and speaking of tiny. … Dr. Dobb’s Journal of Software Tools, 25(2): 18, February 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ieee.org/
conference/iedm.

Baran:2000:NVf

[Bar00f] Nicholas Baran. News and views: Freenet: More anarchy for the Internet?; magnetic properties key to nanoengineering; nanoseconds not fast enough? here come femtoseconds; Caltech leads U.S. field in ACM programming contest; robotic surgeons may make fewer mistakes; free software for designing ICs. Dr. Dobb’s Journal of Software Tools, 25 (6):18, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVIb

[Bar00g] Nicholas Baran. News and views: “Intelligent” ink dries to form nanostucture; Europeans look to sci-fi for new ideas; computer program can act on brain signals; inventory of automatic memory protection [richard bloch] passes away; ingestible camera could replace standard endoscope. Dr. Dobb’s Journal of Software Tools, 25(8):18, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVIa

[Bar00h] Nicholas Baran. News and views: Internet security in “infantile state”; Hedy Lamarr: Spread spectrum pioneer; new weather supercomputer not so super; researchers like armies of cheap robots; Kurzweil predicts human brain uploads. Dr. Dobb’s Journal of Software Tools, 25(4):18, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVM

[Bar00i] Nicholas Baran. News and views: More on tiny transistors; Open Source repository launched; design contest promotes new software tools; and then there’s a decryption contest; Fred Brooks wins ACM Turing Award. Dr.

Dobb’s Journal of Software Tools, 25(3):18, March 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://

sourceforge.net/.

Baran:2000:NVNb

[Bar00j] Nicholas Baran. News and views: Neural net crowned king; smart shirts monitor vital signs; McMaster team wins college design contest; Java applet tracks space station; Sandia Red Team batting 1000; new digital geometry compression algorithm. Dr. Dobb’s Journal of Software Tools, 25 (10):18, October 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVNa

[Bar00k] Nicholas Baran. News and views: New modem standards should speed up Internet access; robocopter: AI lifts off; feet don’t fail me now; IBM claims world’s fastest supercomputer; new color displays based on light-emitting polymers; W3C moves forward with XLink. Dr. Dobb’s Journal of Software Tools, 25(9): 18, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVR

[Bar00l] Nicholas Baran. News and views: RSA algorithm in the public domain; Woz joins the Inventors Hall of Fame; entangled photons mean faster, smaller ICs; BEHEMOTH mothballed; Advanced Encryption Standard selected; SGI releases SDK as open source; WSDL spec released. Dr. Dobb’s Journal of Software Tools, 25(12):18, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:NVV

[Bar00m] Nicholas Baran. News and views: Volunteers needed for weather simulation model; Bill Gates meets Bill Blass — not!; computer lemon law may debut in Pennsylvania; inventor of Huffman encoding dies; green cards, not guest visas; online forums on electronic recycling. Dr. Dobb’s Journal of Software Tools, 25(1):18, January 2000. CODEN DDJOEB. ISSN 1044-789X.

Barker:2000:BBA

[Bar00n] John S. Barker. Building the Bell Atlantic Tariff Management System. Dr. Dobb’s Journal of Software Tools, 25(5):90, 92–95, 97, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2001:LTW

[Bar01a] Nicholas Baran. Load testing Web sites. Dr. Dobb’s Journal of Software Tools, 26(3):112, 114, 116, 118–119, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2001/2001_03/
                                        webload.txt.

Baran:2001:NVA

[Bar01b] Nicholas Baran. News and views: Anonymity and the Internet; is industry hogging computer science talent?; relief from acronyms; OpenML spec released; C# not just a Java killer, says anders; and the winners are …. Dr. Dobb’s Journal of Software Tools, 26 (7):18, July 2001. CODEN DDJOEB. ISSN 1044-789X.

Baran:2001:NVCa

[Bar01c] Nicholas Baran. News and views: Cofounder [victor grinich] of Fairchild Semiconductor dies; Computer History Museum nominates new fellows; high schoolers win scholarships with new algorithm; examples of fair use wanted. Dr. Dobb’s Journal of Software Tools, 26(2):18, February 2001. CODEN DDJOEB. ISSN 1044-789X.

Baran:2001:NVCb

[Bar01d] Nicholas Baran. News and views: College Board to add Java test; astronomy project beats out algorithm for science prize; consortium to release XML business specification; open source developer agreement proposed; consortium hopes to eliminate computer failures; new linear algorithm for sequence analysis; distributing computing à la SETI catches on. Dr. Dobb’s Journal of Software Tools, 26(3):18, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://publicscience.net/
                                        ;   http://www.ddj.com/.

Baran:2001:NVM

[Bar01e] Nicholas Baran. News and views: Mail server benchmark released; new traveling salesman algorithm; ACLU challenges ICANN on freedom of speech; Microsoft opts for decaf; store your data on the Internet; and mind your electronic trash. Dr. Dobb’s Journal of Software Tools, 26(4):18, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL

                                        http://www.cs.wustl.edu/
                                        ;   http://www.ddj.com/;
                                        http://www.spec.org/osg/
                                        mail2001/docs/FAQ.html;
                                        http:/www.research.att.com/
                                        ~alb/ALENEX01.

Baran:2001:NVP

[Bar01f] Nicholas Baran. News and views: Programming the shapes of sculpture; computation expert wins Turing Award; Cobol programmers a vanishing breed?; science controversies go online; MIT builds sociable robot; research finds flaw in DNA; binary pioneer [Claude Shannon] dies at 84. Dr. Dobb’s Journal of Software Tools, 26(5):18, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Baran:2001:NVE

[Bar01g] Nicholas Baran. News and views: The explosion of digital information; new algorithm can handle genetic data; MP3 creators win German Future Prize; optical breakthrough in quantum computing; Fibre Channel Standard closer to reality; NSF funds information technology projects. Dr. Dobb’s Journal of Software Tools, 26(1):18, January 2001. CODEN DDJOEB. ISSN 1044-789X.

Baran:2001:NVW

[Bar01h] Nicholas Baran. News and views: WSDL goes to W3C for standardization; shortchanging science; EUVL may keep Moore’s Law going; spy satellites to generate high-tech jobs; Mexican government adopts Linux; supercomputer on a chip in the works; brain scan database goes public. Dr. Dobb’s Journal of Software Tools, 26(6):18, June 2001. CODEN DDJOEB. ISSN 1044-789X.

Barrett:2003:DPJ

[Bar03] Tom Barrett. Dynamic proxies in Java and .NET. Dr. Dobb’s Journal of Software Tools, 28(7):18, 20, 22, 24, 26, July 2003. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2003/2003_07/dyproxy.
zip.

Bayern:2001:AAS

[Bay01] Shawn Bayern. Algorithm alley: Synchronized recursion. Dr. Dobb’s Journal of Software Tools, 26(6):151– 155, June 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_06/aa0601.
                                        txt.

Bettini:2001:JNC

[BC01] Lorenzo Bettini and Donato Cappetta. A Java 2 network class loader. Dr. Dobb’s Journal of Software Tools, 26(2):58, 60, 62, 64, February 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_02/
                                        jload.txt;   http://www.
                                        ddj.com/ftp/2001/2001_02/
                                        jload.zip.

Bloss:2003:LLV

[BDM^{+}03] Wayne Bloss, DDJ, Ian MacKinnel, Jeff Adams, Michael Mahon, and Henri Socha. Letters: Loathing VB; determining processor utilization; only stone endures. Dr. Dobb’s Journal of Software Tools, 28(1):10, January 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7758/ddj0301l/.

Becker:2001:SMW

[Bec01] Thomas Becker. Synchronization monitors for Win32. Dr. Dobb’s Journal of Software Tools, 26(12):46, 48, 50–52, 54, December 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_12/monitor.

txt;   http://www.ddj.com/
ftp/2001/2001_12/monitor.
zip.

Bentley:2000:AAB

[Ben00] Jon Bentley. Algorithm alley: Binary search: Algorithm, code, and caching. Dr. Dobb’s Journal of Software Tools, 25(4):111–112, 114, 116, April 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_04/
aa400.txt.

Bentley:2001:AAS

[Ben01] Jon Bentley. Algorithm alley: Suffix arrays. Dr. Dobb’s Journal of Software Tools, 26 (4):145–147, April 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_04/
aa0401.txt.

Bernardin:2000:AC

[Ber00] Laurent Bernardin. Analytical computing. Dr. Dobb’s Journal of Software Tools, 25(9): 42, 44, 46–48, September 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_09/
analytic.txt.

Berg:2001:DCC

[Ber01] Cliff Berg. Distributed computing component lifecycles. Dr. Dobb’s Journal of Software Tools, 26(9):32, 34, 36, 39–40, September 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_09/
lifecycl.zip.

Bauer:2001:LSY

[BES^{+}01] Barr Bauer, Jonathan Erickson, Richard Stallman, Mike Pentney, and Frank C. Earl. Letters: The state of your taxes; embedded space; more Open Source versus Free Software; shared source?; CORBA interoperability. Dr. Dobb’s Journal of Software Tools, 26 (9):10, 12, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Betz:2002:BMN

[Bet02]  David Betz.  Bob meets

Nuon. Dr. Dobb’s Journal of Software Tools, 27(2):60, 62– 64, February 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_02/bob.zip;
 http://www.ddj.com/ftp/
2002/2002_02/bobnuon.txt.

Begelman:2001:CST

[BFG01] Gregory Begelman, Lev Finkelstein, and Evgeniy Gabrilovich. C++ set-theoretic operations on virtual containers. Dr. Dobb’s Journal of Software Tools, 26(10):40, 42, 46, 48, October 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_10/
setcpp.txt;   http://www.
ddj.com/ftp/2001/2001_10/

setcpp.zip. See correction [RSW^{+}01].

Bik:2001:AEI

[BGGT01] Aart J. C. Bik, Milind Girkar, Paul M. Grey, and Xinmin Tian. Automatically exploiting implicit parallelism. Dr. Dobb’s Journal of Software Tools, 26(7):28, 30, 32–34, July 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_07/
parallel.txt.

Bonzini:2001:LHG

[BHP^{+}01] Paolo Bonzini, Stuart Halloway, John Penry, Oluseyi Sonaiya, Bruce E. Hogman, Greg Bissell, Michael Hobbs, and Ben Laurie. Letters: Huge GCC executables; Java class loader; Department of Dumb Ideas; setting the record straight; the legacy of C#; DHTML source-code correction; shared libraries aren’t all bad; Zuse and Intel. Dr. Dobb’s Journal of Software Tools, 26(8):10, 12, August 2001. CODEN DDJOEB. ISSN 1044-789X.

Birkett:2000:GCM

[Bir00] David S. Birkett. Graphing calculators meet PCs. Dr. Dobb’s Journal of Software Tools, 25(9):52, 54, 56–58, 60, 62, September 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_09/maxwell.
zip.

Boling:2001:EHK

[BJ01a] Eli Boling and Chuck Jazdzewski. Exception handling in Kylix: Matching up Object Pascal, Linux, and exception handling. Dr. Dobb’s Journal of Software Tools, 26(11):66, 68–71, November 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_11/
                                        kylix.txt.

Brew:2001:VLS

[BJ01b] William Brew and Maggie Johnson. Value lattice static analysis. Dr. Dobb’s Journal of Software Tools, 26(3):30, 32, 36, 38, March 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_03/lattice.
                                        txt.

Bjorkander:2001:PSU

[Bjö01]  Morgan Björkander.  Pro-

gramming in SDL and UML. Dr. Dobb’s Journal of Software Tools, 26(6):93, 96–99, June 2001. CODEN DDJOEB. ISSN 1044-789X.

Bapst:2008:SIO

[BK08] Frederic Bapst and François Kilchoer. Signalling integer overflows in Java: A tool for checking overflows in Java code. Dr. Dobb’s Journal of Software Tools, 33(9):54– 58, September 2008. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://home.hefr.ch/
                                        bapst/cojac.

Barros:2000:LDD

[BKCD00] Paulo Barros, Tim Kientzle, Paul Condon, and Tony Duarte. Letters: Database debates. Dr. Dobb’s Journal of Software Tools, 25(9): 10, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:Ia

[BL00a] Nicholas Baran and Amy Lincicum. Of interest. Dr. Dobb’s Journal of Software Tools, 25(1):126–127, January 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:Ib

[BL00b] Nicholas Baran and Amy Lincicum. Of interest. Dr. Dobb’s Journal of Software Tools, 25(2):142–143, February 2000. CODEN DDJOEB. ISSN 1044-789X.

Baran:2000:Ic

[BL00c] Nicholas Baran and Amy Lincicum. Of interest. Dr. Dobb’s Journal of Software Tools, 25 (3):142–143, March 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Baran:2000:Id

[BL00d] Nicholas Baran and Amy Lincicum. Of interest. Dr. Dobb’s Journal of Software Tools, 25 (4):134–135, April 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Baran:2000:Ie

[BL00e] Nicholas Baran and Amy Lincicum. Of interest. Dr. Dobb’s Journal of Software Tools, 25 (5):134–135, May 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Baran:2000:If

[BL00f] Nicholas Baran and Amy Lincicum. Of interest. Dr. Dobb’s Journal of Software Tools, 25 (6):150–151, June 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Bloom:2000:UAP

[Blo00] Ryan Bloom. Using Apache portable run-time. Dr. Dobb’s Journal of Software Tools, 25(10):100, 102, 104– 106, October 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_10/
aprt.txt.

Boeke:2003:SMC

[Boe03] Wouter Boeke. The SmlMC model checker. Dr. Dobb’s Journal of Software Tools, 28 (3):48, 50–53, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7826/
ddj0303f/.

Bradski:2000:OL

[Bra00] Gary Bradski. The OpenCV library. Dr. Dobb’s Journal of Software Tools, 25(11):120, 122–125, November 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_11/
opencv.txt.

Bright:2002:DPL

[Bri02a] Walter Bright. The D programming language. Dr. Dobb’s Journal of Software Tools, 27(2):36–40, February 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ftp/2002/2002_
02/d.txt.

Brinkmann:2002:GGG

[Bri02b] Peter Brinkmann. Gumbie:

A GUI generator for Jython. Dr. Dobb’s Journal of Software Tools, 27(4):46–50, April 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2002/2002_04/
gumbie.txt;   http://www.
ddj.com/ftp/2002/2002_04/
gumbie.zip.

Brookes:2003:XDB

[Bro03] Allen Brookes. XML data binding. Dr. Dobb’s Journal of Software Tools, 28(3): 26, 28, 30, 32, 35–36, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7826/ddj0303c/.

Bruno:2002:JQ

[Bru02] Eric J. Bruno. Java Q&A:

So what is a Java event agent? Dr. Dobb’s Journal of Software Tools, 27(7): 86–91, July 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_07/jqa0702.
txt;   http://www.ddj.com/

                                        ftp/2002/2002_07/jqa0702.
                                        zip.

Bloom:2000:HLP

[BRVW00] Bard Bloom, Jim Russell, John Vlissides, and Mark Wegman. High-level program development. Dr. Dobb’s Journal of Software Tools, 25(12 (supplement)):17–21, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Boyer:2002:LDS

[BTTF02] John Boyer, Andrew D. Todd, Jason Trenough, and Doug Farrell. Letters: Defective sign-and-encrypt and healthcare woes and J2EE cache and pool and Regex++. Dr. Dobb’s Journal of Software Tools, 27 (2):10, February 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Bugg:2003:SSM

[Bug03] Keith Bugg. SVG and smart maps. Dr. Dobb’s Journal of Software Tools, 28(3):38, 40–41, March 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7826/
                                        ddj0303d/.

Buis:2002:SLS

[Bui02] Paul Buis. Socket-level server programming and .NET. Dr. Dobb’s Journal of Software Tools, 27(9):25–26, 28, 30, 32, September 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/

ftp/2002/2002_09/tcpecho.
txt;   http://www.ddj.com/
ftp/2002/2002_09/tcpecho.
zip.

Burton:2000:GCR

[Bur00] Joshua W. Burton. Garbage collection on the run. Dr. Dobb’s Journal of Software Tools, 25(4):46, 48–49, 51– 53, April 2000. CODEN DDJOEB. ISSN 1044-789X.

But:2002:CSL

[But02] Jason But. A C++ socket library for Linux. Dr. Dobb’s Journal of Software Tools, 27 (6):19–22, 24, June 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_06/
socketcc.txt;   http://www.
ddj.com/ftp/2002/2002_06/
socketcc.zip.

Beckman:2000:OSM

[BW00] Pete Beckman and Gregory V.

Wilson. Open Source meets big iron. Dr. Dobb’s Journal of Software Tools, 25(6):30, 32, 34–35, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Bykov:2000:AYJ

[Byk00] Alex Bykov. Ace your job interview. Dr. Dobb’s Journal of Software Tools, 25(4):??, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Calbucci:2000:WSD

[Cal00a] Marcelo Calbucci. Windows 2000 security descriptors. Dr. Dobb’s Journal of Software Tools, 25(11):57–58, 60, 63, 66, November 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_11/
sdw2k.txt;   http://www.
ddj.com/ftp/2000/2000_11/
sdw2k.zip.

Calbucci:2000:LSF

[Cal00b] Marcelo A. F. Calbucci. LDAP search filters. Dr. Dobb’s Journal of Software Tools, 25(5):44, 46, 48, 50–51, May 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_05/
ldapsrch.txt;   http://www.
ddj.com/ftp/2000/2000_05/
ldapsrch.zip.

Calsavara:2000:JQH

[Cal00c] Alexandre Pereira Calsavara. Java Q&A: How can I extend Java’s search capabilities? Dr. Dobb’s Journal of Software Tools, 25(12):141–142, 144, 146, December 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_12/jqa0012.
zip.

Cargill:2001:AA

[Car01] Tom Cargill. Algorithm alley: Replacing a dictionary with a square root. Dr. Dobb’s Journal of Software Tools, 26(10): 123–125, October 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_10/
aa1001.txt.

Carles:2002:AERa

[Car02a]  Dani Carles.  Adding ex-

ceptions and RTTI to the Windows Ce compiler: Part I. Dr. Dobb’s Journal of Software Tools, 27(8): ??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_08/tcu_mod.
txt;   http://www.ddj.com/
ftp/2002/2002_08/tcu_mod.
zip.

Carles:2002:AERb

[Car02b] Dani Carles. Adding exceptions and RTTI to the Windows Ce compiler: Part II. Dr. Dobb’s Journal of Software Tools, 27(9):52, 54, 56, 58, 60, September 2002. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/2002/2002_09/tcu_
pt2.txt.

Cavalier:2002:DPU

[Cav02] Forrest J. Cavalier III. Debugging PHP using a C++ compiler. Dr. Dobb’s Journal of Software Tools, 27(3): 42, 44–46, March 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_03/
phpcpp.txt;   http://www.
ddj.com/ftp/2002/2002_03/
phpcpp.zip.

Collins:2000:LMW

[CDsJ^{+}00] Terry Collins, Geoff Davis, Tarjei Tjxstheim [sic] Jensen, Raj Dash, DDJ, Rick Box, Dimitrios Souflis, James K. Yun, Guy Hammond, and DDJ. Letters: More worker shortage; music city; sorting through sorts; buy the book; E-address change; Open Source debate. Dr. Dobb’s Journal of Software Tools, 25 (6):10, 12, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Chasin:2000:PSM

[Cha00]  Andrew Chasin.  Predica-

tion, speculation, and modern CPUs. Dr. Dobb’s Journal of Software Tools, 25(5):58, 60– 61, 63–64, May 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_05/
ps_cpu.txt.

Chawathe:2003:SCS

[Cha03] Monica Chawathe. SA-C: Single assignment C. Dr. Dobb’s Journal of Software Tools, 28 (5):60, 62–64, May 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7927/
ddj0305f/.

Chester:2001:HPW

[Che01] Tim Chester. High-performance Web sites: ADO versus MSXML. Dr. Dobb’s Journal of Software Tools, 26(10):81– 84, 86, 88, October 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_10/
enroll_txt.zip.

Cheng:2002:CNC

[Che02] Harry H. Cheng. C99 and numeric computing. Dr. Dobb’s Journal of Software Tools, 27 (3):28, 30–34, March 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2002/2002_03/
c99.txt.

Christie:2000:BAD

[Chr00] Stewart Christie. Building an OS-aware debugger. Dr. Dobb’s Journal of Software Tools, 25(12):94, 96, 98, 100, 102, 104, 106, December 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_12/
osdebug.txt.

Chapman:2000:EPT

[CK00] Mitch Chapman and Brian Kelley. Examining the PyGtk toolkit. Dr. Dobb’s Journal of Software Tools, 25(4):82, 84, 86, 88, April 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_04/
pygtk.txt.

Cory:2000:CQP

[CL00] David Cory and Raymond Laflamme. Computing with quantum physics. Dr. Dobb’s Journal of Software Tools, 25(12 (supplement)):29–32, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Coats:2001:ICU

[CM01] Mark Coats and Terry Mellon. Integrating CMOS with UML: Building fully traceable models. Dr. Dobb’s Journal of Software Tools, 26(6):21– 22, 24, 26, 28, 30, 32–33, 35–36, June 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_06/cmosuml.
                                        zip.

Cromwell:2000:LWS

[CNS^{+}00] Jeff Cromwell, Jim North, J. Stephen Riley Silber, Graham M. Sherrington, and John Gegus. Letters: Worker shortage; the sixth sense, redux; kryptos message; data structures as objects. Dr. Dobb’s Journal of Software Tools, 25 (5):10, 12, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2000:Ia

[Coc00a] Shannon Cochran. Of interest. Dr. Dobb’s Journal of Software Tools, 25(7):150–151, July 2000. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2000:Ib

[Coc00b] Shannon Cochran. Of interest. Dr. Dobb’s Journal of Software Tools, 25(8):150–151, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2000:Ic

[Coc00c] Shannon Cochran. Of interest. Dr. Dobb’s Journal of Software Tools, 25(9):138–139, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2000:Id

[Coc00d] Shannon Cochran. Of interest.

Dr. Dobb’s Journal of Software Tools, 25(10):166–167, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2000:Ie

[Coc00e] Shannon Cochran. Of interest.

Dr. Dobb’s Journal of Software Tools, 25(11):166–167, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2000:If

[Coc00f] Shannon Cochran. Of interest.

Dr. Dobb’s Journal of Software Tools, 25(12):166–167, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2000:TCP

[Coc00g] Shannon Cochran. Top 10 cities for programmers. Dr. Dobb’s Journal of Software Tools, 25(10):??, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:NV

[Coc01a] Shannon Cochran. News and views: Curl programming contest winners announced; prior art database will protect open software; AI researcher [Kenneth Mark Colby] dies; personal router for surfing wireless services; XML schema formalized; thought-controlled computers. Dr. Dobb’s Journal of Software Tools, 26(8):18, August 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:NVI

[Coc01b]  Shannon Cochran.  News

and views: ICFP programming contest concludes; lost moon landing tape recovered; W3C patent controversy; the UNIX epoch; connecting the quantum dots; Programmer’s Guild launches membership campaign. Dr. Dobb’s Journal of Software Tools, 26(12): 20, December 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:NVL

[Coc01c] Shannon Cochran. News and views: Linux Standard Base 1.0 released; now everyone knows you’re a dog; maximum fiberoptic capacity calculated; mass producing quantum chips; mo’ better batteries; patenting the software service model. Dr. Dobb’s Journal of Software Tools, 26(10): 18, October 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:NVS

[Coc01d] Shannon Cochran. News and views: Scientists seek immersive reality; USENIX names lifetime achievement recipients [the GNU Project and the Kerberos network authentication system]; robots need programmers; evangelizing the Semantic Web; get your supercomputer software free; Usenet creator Jim Ellis dies; DARPA funds FreeBSD security initiative. Dr. Dobb’s Journal of Software Tools, 26(9): 18, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:NVW

[Coc01e] Shannon Cochran. News and views: WEP [wired equivalent privacy]: Pining for the fjords?; bye bye Be; UNIX utilities open sourced; cosmology computer calculates creation; UML 2.0 infrastructure proposals. Dr. Dobb’s Journal of Software Tools, 26(11): 18, November 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:Ia

[Coc01f] Shannon Cochran. Of interest. Dr. Dobb’s Journal of Software Tools, 26(1):206– 207, January 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:Ib

[Coc01g] Shannon Cochran. Of interest.

Dr. Dobb’s Journal of Software Tools, 26(2):166–167, February 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:Ic

[Coc01h] Shannon Cochran. Of interest. Dr. Dobb’s Journal of Software Tools, 26(3):150– 151, March 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:Id

[Coc01i] Shannon Cochran. Of interest. Dr. Dobb’s Journal of Software Tools, 26(4):158– 159, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:Ie

[Coc01j] Shannon Cochran. Of interest. Dr. Dobb’s Journal of Software Tools, 26(5):158–159, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:If

[Coc01k] Shannon Cochran. Of interest. Dr. Dobb’s Journal of Software Tools, 26(6):166–167, June 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2001:RCS

[Coc01l] Shannon Cochran. The rising costs of software complexity. Dr. Dobb’s Journal of Software Tools, 26(4):??, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVb

[Coc02a] Shannon Cochran. News and views. Dr. Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVCa

[Coc02b] Shannon Cochran. News and views: Chains for nanotechnology; high-dependability computing program launched; personal computers rule; combating cellular phone theft; alien noises; crystal traps light; 2001 Turing Award [Ole-Johan Dahl and Kristen Nygaard] announced. Dr. Dobb’s Journal of Software Tools, 27(4):

16, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVCb

[Coc02c] Shannon Cochran. News and views: China takes ACM programming contest; supercomputing Science Grid nearing birth; hacking the glow of the screen; PARC searching for sponsors; milestone in DNA computing; Holzmann receives ACM Software Award; games aren’t just for playing around. Dr. Dobb’s Journal of Software Tools, 27(6):14, June 2002. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVP

[Coc02d] Shannon Cochran. News and views: Peek-a-Booty ready to play; back to basics [IBM Meta Pad]; better batteries [Lawrence Livermore National Laboratory MEMS fuel cell]; self-repairing plastic [Automend]; outlawing E-waste [lead in computer monitors]. Dr. Dobb’s Journal of Software Tools, 27(5):14, May 2002. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVR

[Coc02e]  Shannon Cochran.  News

and views: Researchers tackle P2P network; Java upstarts push JEFF; computer science tests canceled; the continuing voyage. Dr. Dobb’s Journal of Software Tools, 27(12): 14, December 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7718/
                                        ddj0212n/.

Cochran:2002:NVSb

[Coc02f]  Shannon Cochran.  News

and views: School of Adaptive Computer Training; it seems like yesterday… [10th anniversary of the first Web site]; double-gate transistor breakthrough; 802.11g Standard proposed; 30th anniversary of the [Intel] 4004; DeCSS legal decisions. Dr. Dobb’s Journal of Software Tools, 27 (2):18, February 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVSd

[Coc02g] Shannon Cochran. News and views: Smalltalk turns 30; the billionth PC; reverse challenge concludes; Edsger Dijkstra passes away. Dr. Dobb’s Journal of Software Tools, 27 (10):14, October 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVSa

[Coc02h] Shannon Cochran. News and views: Speech application language tags proposed; Computer Museum names new Fellows; terascale computing system powers up; the enigma of Mrs. BB; really, really tiny transistors developed. Dr. Dobb’s Journal of Software Tools, 27(1):18, January 2002. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVSc

[Coc02i] Shannon Cochran. News and views: Stanford student winds collegiate tournament; Web services get real; new supercomputer on the block; 2001 Lovelace Award announced. Dr. Dobb’s Journal of Software Tools, 27(7):14, July 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVc

[Coc02j] Shannon Cochran. News and views: The programming alphabet expands; IMSAI, series two; webcasters shut down; open mobile alliance; GUI for robot control; business plan boneyard. Dr. Dobb’s Journal of Software Tools, 27(9): 14, September 2002. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVW

[Coc02k] Shannon Cochran. News and views: WaSP [Web Standards Project] buzzes off; Eclipse Project on the horizon; semiconductor roadmap: Ramping up, scaling down; AES [Advanced Encryption Standard]: Its finally official; SMS [Short Message Service] shines on; Berners-Lee awarded Japan Prize. Dr. Dobb’s Journal of Software Tools, 27(3): 14, March 2002. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2002:NVY

[Coc02l] Shannon Cochran. News and views: Yet another graphics language; instant on boot cycles; fourth and long: It’s all in the numbers; literate leo; AI standards sought for video games; out with the old, in with the new. Dr. Dobb’s Journal of Software Tools, 27 (11):14, November 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7644/
ddj0211n/.

Cochran:2003:NVd

[Coc03a] Shannon Cochran. News and views: Caltech “Turing Tournament”; science and engineering Ph.D.s getting scarcer; E. Coli for data storage; Swiss accept Internet ballots; extensible resource identifiers. Dr. Dobb’s Journal of Software Tools, 28(4):10, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7827/
ddj0304o/.

Cochran:2003:NVC

[Coc03b]  Shannon Cochran.  News

and views: Cryptographers [Ronald Rivest, Adi Shamir, and Leonard Adleman] receive Turing Award; computerscience pioneer [John G. “Jack” Herriot] passes away; programming quantum computers; the demography of the Internet. Dr. Dobb’s Journal of Software Tools, 28(7):14, July 2003. CODEN DDJOEB. ISSN 1044-789X.

Cochran:2003:NVGb

[Coc03c]  Shannon Cochran.  News

and views: Government ponders open-source strategy; POSIX, Single UNIX specification merged; IBM plans new supercomputers; robotic surgeons have a heart. Dr. Dobb’s Journal of Software Tools, 28 (2):14, February 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7790/
ddj0302o/.

Cochran:2003:NVe

[Coc03d] Shannon Cochran. News and views: MIT honors builder of robot swarm; developing space; Embedded Linux platform specification released; biometric systems at U.S. borders, says NIST. Dr. Dobb’s Journal of Software Tools, 28(5):14, May 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7927/
ddj0305o/.

Cochran:2003:NVR

[Coc03e] Shannon Cochran. News and views: The robots are coming; Fortran 2000 drafted; touch relayed across Atlantic; Kramnik versus Deep Fritz; 2002 ICFP Programming Content concludes. Dr. Dobb’s Journal of Software Tools, 28(1):14, January 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7758/ddj0301n/.

Cochran:2003:NVT

[Coc03f] Shannon Cochran. News and views: Tiny executable contest concludes; DeCSS, DMCA prosecutions fail; studying for the Turing test; Embedded Linux specification released. Dr. Dobb’s Journal of Software Tools, 28(3):14, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7826/ddj0303n/.

Cochran:2003:NVf

[Coc03g] Shannon Cochran. News and views: Young programmers vie in contests; happy birthday World Wide Web; Microsoft issues Visual C++ bug alert; etiquette and bug reports; Commerce Department fights international tech standards. Dr. Dobb’s Journal of Software Tools, 28(6):14, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        8213/ddj0306n/.

Cochran:2003:Ia

[Coc03h] Shannon Cochran. Of interest. Dr. Dobb’s Journal of Software Tools, 28(5):86, May 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7927/ddj0305p/.

Cochran:2003:Ib

[Coc03i] Shannon Cochran. Of interest. Dr. Dobb’s Journal of Software Tools, 28(6):78, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
8213/ddj0306o/.

Cochran:2003:Ic

[Coc03j] Shannon Cochran. Of interest.

Dr. Dobb’s Journal of Software Tools, 28(7):78, July 2003. CO-DEN DDJOEB. ISSN 1044-789X.

Cohen:2000:BTF

[Coh00] Noam Cohen. Building a testpoint framework. Dr. Dobb’s Journal of Software Tools, 25 (3):46, 48–50, March 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_03/
testpont.txt;   http://www.
ddj.com/ftp/2000/2000_03/
testpont.zip.

Cohen:2001:COR

[Coh01] Noam Cohen. COM object reference counting. Dr. Dobb’s Journal of Software Tools, 26 (3):58, 60–61, 63, March 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_03/
refcatch.txt;   http://www.
ddj.com/ftp/2001/2001_03/
refcatch.zip.

Cohen:2002:JQH

[Coh02] Tal Cohen. Java Q&A: How do I correctly implement the equals() method? Dr. Dobb’s Journal of Software Tools, 27 (5):83–84, 86, May 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_05/
                                        jqa0502.txt.

Cohen:2003:CST

[Coh03] Noam Cohen. C++ stack traces. Dr. Dobb’s Journal of Software Tools, 28(2): 34, 36, February 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7790/
                                        ddj0302d/.

Cole:2000:WBT

[Col00] Oliver Cole. White-box testing. Dr. Dobb’s Journal of Software Tools, 25(3):23– 26, 28, March 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_03/
                                        wboxtest.txt.

Cox:2000:IWP

[Cox00]  David Cox.  Integrating

Web pages with databases. Dr. Dobb’s Journal of Software Tools, 25(9):94, 96– 98, September 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_09/
                                        ceres.txt;   http://www.
                                        ddj.com/ftp/2000/2000_09/
                                        ceres.zip.

Cox:2001:JQH

[Cox01a]  Brad Cox.  Java Q&A:

How do I handle multiline strings? Dr. Dobb’s Journal of Software Tools, 26(6):

147–149, June 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_06/jqa0601.
txt.

Cox:2001:WAJ

[Cox01b] Brad Cox. Web applications as Java servlets. Dr. Dobb’s Journal of Software Tools, 26(5): 97–100, 102–104, May 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_05/
webapp.txt;   http://www.
ddj.com/ftp/2001/2001_05/
webapp.zip.

Cox:2001:PX

[Cox01c] David Cox. Parsing XML.

Dr. Dobb’s Journal of Software Tools, 26(1):96, 98, 100, January 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_01/minixml.
txt;   http://www.ddj.com/
ftp/2001/2001_01/minixml.
zip.

Cox:2001:XIC

[Cox01d] David Cox. XML for instrument control and monitoring. Dr. Dobb’s Journal of Software Tools, 26(11):83– 85, November 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_11/
ixml.txt.

Cromwell:2000:PBA

[Cro00a]  Jeff Cromwell.  Program-

mer’s bookshelf: The art and science of software architecture. Dr. Dobb’s Journal of Software Tools, 25(6):143–144, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Crooks:2000:EGS

[Cro00b] Clayton E. Crooks II. Examining the 3D Game Studio Toolkit. Dr. Dobb’s Journal of Software Tools, 25(8):106, 108– 111, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Cromwell:2001:PBD

[Cro01a] Jeff Cromwell. Programmer’s bookshelf: Days of our data: A server SOAP opera. Dr. Dobb’s Journal of Software Tools, 26 (7):138, 140, July 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Crooks:2001:AVB

[Cro01b] Clayton E. Crooks II. App-Forge: Visual Basic for the Palm Os. Dr. Dobb’s Journal of Software Tools, 26(8):56, 58, 60, August 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_08/appforg.
                                        txt;   http://www.ddj.com/
                                        ftp/2001/2001_08/appforg.
                                        zip.

Crowell:2001:CP

[Cro01c] Talbott Crowell. C# and Perl: C# has a lot in common with many languages. Dr. Dobb’s Journal of Software Tools, 26 (10):50, 52, 54, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_10/
csharpperl.txt.

Cochran:2001:Ig

[CS01a] Shannon Cochran and Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 26 (7):142–143, July 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Cochran:2001:Ih

[CS01b] Shannon Cochran and Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 26 (8):126–127, August 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Carothers:2002:CMP

[CS02] Christopher D. Carothers and Boleslaw K. Szymanski. Checkpointing multithreaded programs. Dr. Dobb’s Journal of Software Tools, 27(8): ??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_08/checkpt.
txt.

Cunningham:2000:BPL

[Cun00] Tom Cunningham. A BSP porting library for embedded peripherals. Dr. Dobb’s Journal of Software Tools, 25(7):80, 82, 84, 86–87, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
bspseral.zip.

Dale:2001:BSA

[Dal01a] Richard Dale. Biometric security: It’s all about identification and authentication. Dr. Dobb’s Journal of Software Tools, 26(11):93–94, 96, November 2001. CODEN DDJOEB. ISSN 1044-789X.

Dalke:2001:MCE

[Dal01b] Andrew Dalke. Making C extensions more Pythonic. Dr. Dobb’s Journal of Software Tools, 26(1):68, 70, 72, 74, 76, January 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_01/cpython.
txt;   http://www.ddj.com/
ftp/2001/2001_01/cpython.
zip.

Dautelle:2001:JDJ

[Dau01] Jean-Marie Dautelle. J.A.D.E.: The Java Addition to the Default Environment. Dr. Dobb’s Journal of Software Tools, 26 (2):52, 54, 56, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_02/
jade.txt.

Davis:2000:OOD

[Dav00] Thomas E. Davis. Objectoriented design in procedural environments. Dr. Dobb’s Journal of Software Tools, 25 (6):68, 70–72, June 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_06/
oodesign.txt.

Davis:2001:DSE

[Dav01] Don Davis. Defective signand-encrypt: Can you really trust S/MIME, PCKS#7, PGP, and XML? Dr. Dobb’s Journal of Software Tools, 26 (11):30, November 2001. CO-DEN DDJOEB. ISSN 1044-789X.

deDinechin:2001:JQW

[dD01] Christophe de Dinechin. Java Q&A: What is Moka? Dr. Dobb’s Journal of Software Tools, 26(7):125–126, 128, 130–131, July 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_07/jqa0701.
txt.

deCarmo:2000:MGC

[deC00a] Linden deCarmo. The Media Gateway Control Protocol. Dr. Dobb’s Journal of Software Tools, 25(5):23–24, 26–28, 32, May 2000. CODEN DDJOEB. ISSN 1044-789X.

DeCarmo:2000:SPP

[DeC00b] Linden DeCarmo. Security protocols and performance. Dr. Dobb’s Journal of Software Tools, 25(11):40, 44, 46, 48, November 2000. CODEN DDJOEB. ISSN 1044-789X.

deCarmo:2001:PSB

[deC01a] Linden deCarmo. PacketCable security: Building secure packet-based telephony networks. Dr. Dobb’s Journal of Software Tools, 26(11):73–76, 78, November 2001. CODEN DDJOEB. ISSN 1044-789X.

DeCarmo:2001:PPD

[DeC01b]  Linden DeCarmo.  Pock-

etPyro for Palm development. Dr. Dobb’s Journal of Software Tools, 26(3):82, 84–87, March 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_03/
palmpyro.txt.

DeCarmo:2002:DAH

[DeC02] Linden DeCarmo. Digital, analog, and high-definition TV. Dr. Dobb’s Journal of Software Tools, 27(11):46, 48, 50, 52–54, November 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7644/
ddj0211e/.

Demaine:2001:AAF

[Dem01] Erik Demaine. Algorithm alley: Fast and small resizable arrays. Dr. Dobb’s Journal of Software Tools, 26(7): 132–134, July 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_07/
aa0701.txt.

Dalheimer:2002:EDQ

[DH02] Matthias Kalle Dalheimer and Steffen Hansen. Embedded development with Qt/ Embedded: Linux for small systems. Dr. Dobb’s Journal of Software Tools, 27(3):48, 50, 52–54, March 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_03/qtembed.
txt;   http://www.ddj.com/
ftp/2002/2002_03/qtembed.
zip.

Diener:2003:LVV

[DHE03] Edward Diener, Kirt Haden, and Arpad Elo, Jr. Letters: VB versus the world; teacher, teacher; odds & ends. Dr. Dobb’s Journal of Software Tools, 28(3):10, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7826/
ddj0303l/.

Duncan:2001:LPD

[DHR^{+}01] Ray Duncan, Duncan Harris, Douglas Reilly, Craig Rodrigues, Michael Birken, and Paul S. Person. Letters: Plugin desupport; threading and the .Net framework; CORBA interoperability; game over for Java; totally wired. Dr. Dobb’s Journal of Software Tools, 26 (11):10, 12, November 2001. CODEN DDJOEB. ISSN 1044-789X.

deIcaza:2002:MNF

[dIJ02] Miguel de Icaza and Brian Jepson. Mono and the .Net framework. Dr. Dobb’s Journal of Software Tools, 27(1): 21–24, 26, January 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_01/
mono.txt.

DiMaggio:2000:TUD

[DiM00] Len DiMaggio. Testing UNIX daemons. Dr. Dobb’s Journal of Software Tools, 25(3):38, 40, 42, 44, March 2000. CODEN DDJOEB. ISSN 1044-789X.

Dorohonceanu:2000:AAP

[DNM00] Bogdan Dorohonceanu and Craig Nevill-Manning. Algorithm alley: A practical suffixtree implementation for string searches. Dr. Dobb’s Journal of Software Tools, 25(7): 133–136, 140, July 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_07/
                                        aa700.txt;   http://www.
                                        ddj.com/ftp/2000/2000_07/
                                        aa700.zip.

Dorobonceanu:2002:CFN

[Dor02] Bogdan Dorobonceanu. Comparing fuzzy numbers. Dr. Dobb’s Journal of Software Tools, 27(12):38, 40, 42, 44– 45, December 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7718/
                                        ddj0212e/.

Dougan:2002:TAR

[Dou02] Cort Dougan. Two-axis, realtime camera control. Dr. Dobb’s Journal of Software Tools, 27(10):60, 62, 64– 67, October 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_10/
                                        rtlcam.txt;   http://www.

ddj.com/ftp/2002/2002_10/
rtlcam.zip.

Daemen:2001:AAR

[DR01] Joan Daemen and Vincent Rijmen. Algorithm alley: Rijndael: The Advanced Encryption Standard. Dr. Dobb’s Journal of Software Tools, 26 (3):137–139, March 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_03/
aa0301.txt;   http://www.
esat.kuleuven.ac.be/~rijmen/
rijndael/.

Dwelly:2000:JXL

[Dwe00a] Andrew Dwelly. Java, XML, and literate programming. Dr. Dobb’s Journal of Software Tools, 25(2):62, 64–66, 68, February 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_02/litjava.
txt;   http://www.ddj.com/
ftp/2000/2000_02/litjava.
zip.

Dwelly:2000:XRP

[Dwe00b] Andrew Dwelly. XML, reflective pattern matching, and Java. Dr. Dobb’s Journal of Software Tools, 25(6):46, 49–52, 54, June 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_06/
marius05.zip;   http://www.
ddj.com/ftp/2000/2000_06/
xmljava.txt.

Eaddy:2001:CVJ

[Ead01] Marc Eaddy. C# versus Java. Dr. Dobb’s Journal of Software Tools, 26(2):74, 76, 78, 80, 82, February 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_02/csharpj.
                                        txt.

Eisenecker:2001:MBP

[EBC01] Ulrich W. Eisenecker, Frank Blinn, and Krzysztof Czarnecki. Mixin-based programming in C++. Dr. Dobb’s Journal of Software Tools, 26 (1):111, 112, 114, 116, 118, 120–122, January 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_01/
                                        mixin.txt.

Ellis:2001:LOS

[EJS^{+}01] Steven Ellis, David Johnson, Mike Schmit, Jenny Jones, Simon Cooke, and Kurt Granroth. Letters: Open source Cobol; setting the Debian record straight; back to basics; load testing Web sites; open source hat tricks; KDE insider. Dr. Dobb’s Journal of Software Tools, 26(7):10, 12, July 2001. CODEN DDJOEB. ISSN 1044-789X.

Epstein:2000:JQ

[EKM00] David Epstein, Joseph Kiniry, and John Motil. Java Q&A: What is “JJ”? Dr. Dobb’s Journal of Software Tools, 25 (4):101–106, April 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_04/
jqa400.txt.

Ellis:2000:ILS

[Ell00] Richard Ellis. The IT labor shortage: Fact or fiction? Dr. Dobb’s Journal of Software Tools, 25(4):??, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Epplin:2000:IDH

[Epp00a] Jerry Epplin. Inside Debian Hurd. Dr. Dobb’s Journal of Software Tools, 25(12):21–22, 24, 26, December 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Epplin:2000:IRT

[Epp00b] Jerry Epplin. Inside Real-Time Linux. Dr. Dobb’s Journal of Software Tools, 25(3):72, 74, 76, 78, March 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_03/rtlinux.
txt;   http://www.ddj.com/
ftp/2000/2000_03/rtlinux.
zip.

Erickson:2000:DDE

[Eri00a] Jonathan Erickson. Dr. Dobb’s Excellence In Programming Award. Dr. Dobb’s Journal of Software Tools, 25(4): 20–21, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:ECM

[Eri00b] Jonathan Erickson. Editorial: Cat and mouse games.

Dr. Dobb’s Journal of Software Tools, 25(11):8, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:EEE

[Eri00c] Jonathan Erickson. Editorial: Extend, embrace, extinguish: Three strikes and you’re out. Dr. Dobb’s Journal of Software Tools, 25(8):8, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:EGL

[Eri00d] Jonathan Erickson. Editorial: Getting the lead out. Dr. Dobb’s Journal of Software Tools, 25(4):8, April 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:2000:EHM

[Eri00e] Jonathan Erickson. Editorial: Hey, Mr. Postman. Dr. Dobb’s Journal of Software Tools, 25 (5):8, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:EBR

[Eri00f] Jonathan Erickson. Editorial: It’s better to be rich and famous, than to live near an airport. Dr. Dobb’s Journal of Software Tools, 25(1):8, January 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:EMC

[Eri00g] Jonathan Erickson. Editorial: Money can’t buy me love. Dr. Dobb’s Journal of Software Tools, 25(6):10, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:ENY

[Eri00h] Jonathan Erickson. Editorial: Now you UCITA, now you don’t. Dr. Dobb’s Journal of Software Tools, 25(9): 8, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:EOS

[Eri00i] Jonathan Erickson. Editorial:

Open Source, open projects. Dr. Dobb’s Journal of Software Tools, 25(3):10, March 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:ESD

[Eri00j] Jonathan Erickson. Editorial:

Software development: Principles and practice. Dr. Dobb’s Journal of Software Tools, 25 (10):8, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:EUH

[Eri00k] Jonathan Erickson. Editorial:

Under the hood. Dr. Dobb’s Journal of Software Tools, 25 (7):8, July 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:EWP

[Eri00l] Jonathan Erickson. Editorial:

Webcasting in present and future tense. Dr. Dobb’s Journal of Software Tools, 25(12): 8, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2000:EWS

[Eri00m] Jonathan Erickson. Editorial:

Worker shortage a tall tail?

Dr. Dobb’s Journal of Software Tools, 25(2):10, February 2000. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2001:DDE

[Eri01a] Jonathan Erickson. Dr. Dobb’s Excellence In Programming Award. Dr. Dobb’s Journal of Software Tools, 26(5):20–21, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2001:EWA

[Eri01b] Jonathan Erickson. Editorial: R_{XML} for what ails healthcare. Dr. Dobb’s Journal of Software Tools, 26(11):8, November 2001. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2001:ERS

[Eri01c] Jonathan Erickson. Editorial: A ringside seat. Dr. Dobb’s Journal of Software Tools, 26 (8):8, August 2001. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2001:EOL

[Eri01d] Jonathan Erickson. Editorial: An open letter on open letters. Dr. Dobb’s Journal of Software Tools, 26(5):8, May 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:2001:EBU

[Eri01e] Jonathan Erickson. Editorial: Back in the U. S. S. R. Dr. Dobb’s Journal of Software Tools, 26(10):8, October 2001. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2001:EDV

[Eri01f] Jonathan Erickson. Editorial: Déjà vu all over again. Dr. Dobb’s Journal of Software Tools, 26(2):8, February 2001. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2001:EDD

[Eri01g] Jonathan Erickson. Editorial: From the Department of Dumb Ideas. Dr. Dobb’s Journal of Software Tools, 26 (3):8, March 2001. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2001:EAA

[Eri01h] Jonathan Erickson. Editorial: It’s all about integrity. Dr. Dobb’s Journal of Software Tools, 26(7):8, July 2001. CODEN DDJOEB. ISSN 1044-789X. Comments on threat from the Recording Industry Association of America (RIAA) and the Secure Digital Music Initiative (SDMI) Foundation that prevented Ed Felton from presenting a conference paper on the defeat of SDMI’s digital watermarking, and on Microsoft’s calling open source “unAmerican”.

Erickson:2001:EOR

[Eri01i] Jonathan Erickson. Editorial:

Open records, closed doors. Dr. Dobb’s Journal of Software Tools, 26(6):8, June 2001. CODEN DDJOEB. ISSN 1044-789X. Essay on open source code issues in government records.

Erickson:2001:EOS

[Eri01j] Jonathan Erickson. Editorial: Operating systems and operating principles. Dr. Dobb’s Journal of Software Tools, 26 (12):8, December 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:2001:ERF

[Eri01k] Jonathan Erickson. Editorial: Realizable fantasies. Dr. Dobb’s Journal of Software Tools, 26(1):8, January 2001. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2001:ESF

[Eri01l] Jonathan Erickson. Editorial: Science fiction in the Electronic Era. Dr. Dobb’s Journal of Software Tools, 26(9): 8, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2001:ESP

[Eri01m] Jonathan Erickson. Editorial: To sleep, perchance to dream. Dr. Dobb’s Journal of Software Tools, 26(4):8, April 2001. CO-DEN DDJOEB. ISSN 1044-789X. Comments on ridiculous trademarks and patents: the frowny face “:-(” has been now trademarked!

Erickson:2002:ETT

[Eri02a] Jon Erickson. Editorial: T^{3}: Technology to teaching. Dr. Dobb’s Journal of Software Tools, 27(12):8, December 2002. CODEN DDJOEB.

ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7718/ddj0212k/.

Erickson:2002:Eb

[Eri02b] Jonathan Erickson. Editorial. Dr. Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2002:EAI

[Eri02c] Jonathan Erickson. Editorial:

Adding it all up. Dr. Dobb’s Journal of Software Tools, 27 (7):8, July 2002. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2002:EWI

[Eri02d] Jonathan Erickson. Editorial: And the winner is …. Dr. Dobb’s Journal of Software Tools, 27(5):8, May 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:2002:EDD

[Eri02e] Jonathan Erickson. Editorial: Deciphering the doors of knowledge. Dr. Dobb’s Journal of Software Tools, 27 (4):8, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2002:EET

[Eri02f] Jonathan Erickson. Editorial:

Early times and numerical weather prediction. Dr. Dobb’s Journal of Software Tools, 27 (1):8, January 2002. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2002:EGC

[Eri02g] Jonathan Erickson. Editorial: Getting a charge out of the future. Dr. Dobb’s Journal of Software Tools, 27 (3):8, March 2002. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2002:ESO

[Eri02h] Jonathan Erickson. Editorial: Strike one …. Dr. Dobb’s Journal of Software Tools, 27 (2):8, February 2002. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2002:ETC

[Eri02i] Jonathan Erickson. Editorial: Taking care of business. … Dr. Dobb’s Journal of Software Tools, 27(10):8, October 2002. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2002:TT

[Eri02j] Jonathan Erickson. Editorial: Texas toast. Dr. Dobb’s Journal of Software Tools, 27 (11):8, November 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7644/
ddj0211k/.

Erickson:2002:Ec

[Eri02k] Jonathan Erickson. Editorial: The name game. Dr. Dobb’s Journal of Software Tools, 27 (9):8, September 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Erickson:2002:EYD

[Eri02l] Jonathan Erickson. Editorial: You don’t always get what you pay for. Dr. Dobb’s Journal of Software Tools, 27(6):8, June 2002. CODEN DDJOEB. ISSN 1044-789X.

Erickson:2003:DDE

[Eri03a] Jon Erickson. 2003 Dr. Dobb’s Excellence in Programming Award. Dr. Dobb’s Journal of Software Tools, 28 (4):16–17, April 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7827/
ddj0304b/.

Erickson:2003:ELL

[Eri03b] Jonathan Erickson. Editorial:

A living legend. Dr. Dobb’s Journal of Software Tools, 28 (1):8, January 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7758/
ddj0301k/.

Erickson:2003:EBPa

[Eri03c] Jonathan Erickson. Editorial: Bidness process outsourcing. Dr. Dobb’s Journal of Software Tools, 28(2):8, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7790/ddj0302l/.

Erickson:2003:EAN

[Eri03d] Jonathan Erickson. Editorial: It’s all in a name. Dr. Dobb’s Journal of Software Tools, 28(6):8, June 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.

                                        ddj.com/documents/s=8213/
                                        ddj0306k/.

Erickson:2003:EC

[Eri03e] Jonathan Erickson. Editorial: Ka-ching, ka-ching, ka-ching. Dr. Dobb’s Journal of Software Tools, 28(5):8, May 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7927/
                                        ddj0305l/.

Erickson:2003:EBPb

[Eri03f] Jonathan Erickson. Editorial: On being a programmer. Dr. Dobb’s Journal of Software Tools, 28(4):8, April 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7827/
                                        ddj0304l/.

Erickson:2003:ESP

[Eri03g] Jonathan Erickson. Editorial: Sales pitch. Dr. Dobb’s Journal of Software Tools, 28(3):8, March 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7826/
                                        ddj0303k/.

Erickson:2003:E

[Eri03h] Jonathan Erickson. Editorial: Setting high standards. Dr. Dobb’s Journal of Software Tools, 28(7):8, July 2003. CO-DEN DDJOEB. ISSN 1044-789X.

Esch:2002:PT

[Esc02]  James Esch.  Program-

mer’s toolchest: The RE-BOL IOS distributed filesystem. Dr. Dobb’s Journal of Software Tools, 27(9):62, 64, 66, September 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_09/
rebolios.txt.

Ethington:2001:DPS

[Eth01] Russ Ethington. Diagnosing proxy server problems. Dr. Dobb’s Journal of Software Tools, 26(3):104, 106, 108, 110, March 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_03/
proxy.txt;   http://www.
ddj.com/ftp/2001/2001_03/
proxy.zip.

Fabian:2001:BNL

[Fab01] Frank Fabian. Balancing network load with priority queues. Dr. Dobb’s Journal of Software Tools, 26(12):75–76, 78, 80, December 2001. CODEN DDJOEB. ISSN 1044-789X.

Farmer:2000:WM

[Far00] Dan Farmer. What are MACtimes? Dr. Dobb’s Journal of Software Tools, 25(10):68, 70– 74, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Farmer:2001:BYD

[Far01] Dan Farmer. Bring out your dead. Dr. Dobb’s Journal of Software Tools, 26(1):102, 104– 105, 107–108, January 2001. CODEN DDJOEB. ISSN 1044-789X.

Fomitchev:2002:RTM

[FH02] Max I. Fomitchev and Joe Hershberger. Real-time music synthesis and embedded applications. Dr. Dobb’s Journal of Software Tools, 27(1):55– 58, 60, January 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_01/
                                        synth.txt;   http://www.
                                        ddj.com/ftp/2002/2002_01/
                                        synth.zip.

Fishburne:2000:MPG

[Fis00] William Fishburne. Margination and Project Gutenberg. Dr. Dobb’s Journal of Software Tools, 25(5):52, 54– 56, May 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_05/gutenbg.
                                        txt;   http://www.ddj.com/
                                        ftp/2000/2000_05/gutenbg.

zip. See letter [GF00].

Fischer:2002:OEL

[Fis02] Todd Fischer. Optimizing Embedded Linux. Dr. Dobb’s Journal of Software Tools, 27 (5):51–57, May 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_05/emlinux.
                                        txt.

Fialli:2003:LXD

[FMPL^{+}03] Joe Fialli, Mike Morley, Eduardo Pelegri-Llopart, Arpad Elo, Jr., and Joseph Roy D. North. Letters: XML data binding; XML-based programming; block-structured languages; history lesson. Dr. Dobb’s Journal of Software Tools, 28(6):10, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=8213/
ddj0306l/.

Fomitchev:2000:AUD

[Fom00a] Max I. Fomitchev. Automatic updates for distributed applications. Dr. Dobb’s Journal of Software Tools, 25(7): 129–131, July 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_07/aupdate.
txt;   http://www.ddj.com/
ftp/2000/2000_07/aupdate.
zip.

Fomitchev:2000:ORT

[Fom00b] Max I. Fomitchev. Optimizing 3Dnow! real-time graphics. Dr. Dobb’s Journal of Software Tools, 25(8): 40, 42–46, August 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_08/
3dnow.txt;   http://www.
ddj.com/ftp/2000/2000_08/
3dnow.zip.

Fox:2000:CAJ

[Fox00a] Joshua Fox. Collaborative applications and the Java Shared Data Toolkit. Dr. Dobb’s Journal of Software Tools, 25(2): 32, 34–40, February 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_02/
                                        jsdt.txt;   http://www.ddj.
                                        com/ftp/2000/2000_02/jsdt.
                                        zip.

Fox:2000:JQW

[Fox00b] Joshua Fox. Java Q&A: So

                                        what is SO KEEPALIVE? Dr.

Dobb’s Journal of Software Tools, 25(9):123–126, 128, 130, September 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_09/jqa0900.
                                        txt;   http://www.ddj.com/
                                        ftp/2000/2000_09/jqa0900.
                                        zip.

Frazier:2001:CPC

[Fra01] George F. Frazier. Crossplatform coroutines in C++. Dr. Dobb’s Journal of Software Tools, 26(3):72, 74, 76, 78–80, March 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_03/
                                        cppco.txt;   http://www.
                                        ddj.com/ftp/2001/2001_03/
                                        cppco.zip.

Freeh:2002:APC

[Fre02] Vincent W. Freeh. Anatomy of a parasitic computer: Computing with the TCP checksum. Dr. Dobb’s Journal of Software Tools, 27(1):63–64, 66–67, January 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_01/
                                        parasite.txt.

Fremberg:2003:MAP

[Fre03] Daniel Fremberg. The Mithra authentication protocol. Dr. Dobb’s Journal of Software Tools, 28(5):44, 46–48, May 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7927/ddj0305d/.

Farmer:2000:FCA

[FV00] Dan Farmer and Wietse Venema. Forensic computer analysis: An introduction. Dr. Dobb’s Journal of Software Tools, 25(9):70, 72–75, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Farmer:2001:BPI

[FV01] Dan Farmer and Wietse Venema. Being prepared for intrusion. Dr. Dobb’s Journal of Software Tools, 26(4):78, 80, 82, 84–85, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Gabrilovich:2003:HL

[Gab03] Evgeniy Gabrilovich. Heap Ltd. Dr. Dobb’s Journal of Software Tools, 28(6):54, 56–57, June 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=8213/
ddj0306f/.

Gonzales:2001:RTD

[GB01] David Ruimy Gonzales and Brian Branson. Real-time debugging and wireless devices. Dr. Dobb’s Journal of Software Tools, 26(3):92, 94, 96– 98, 100–101, March 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Galkowski:2002:LSS

[GDHI02] Jan Theodore Galkowski, Steve Donovan, Roy Hann, and David Irving. Letters: Swaine’s semantic Web; the return of Quincy; DBMS: Present, past, and future; shared libraries aren’t all bad. Dr. Dobb’s Journal of Software Tools, 27(11):10, November 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7644/ddj0211l/.

Geiselbrecht:2001:NOS

[Gei01] Travis K. Geiselbrecht. The NewOS operating system. Dr. Dobb’s Journal of Software Tools, 26(12):33, 35, 38, 40, 42, 44, December 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_12/
newos.txt;   http://www.
ddj.com/ftp/2001/2001_12/

newos.zip. See correction [TLA^{+}02].

Gemmell:2000:SMF

[Gem00] Jim Gemmell. Scalable multicast file distribution. Dr. Dobb’s Journal of Software Tools, 25(5):82, 84–89, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Gettys:2001:AAC

[Get01a] Thomas Gettys. Algorithm alley: Computational geometry: Finding the convex hull. Dr. Dobb’s Journal of Software Tools, 26(5):141– 144, 146, May 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_05/
aa0501.txt.

Gettys:2001:AAG

[Get01b] Thomas Gettys. Algorithm alley: Generating perfect hash functions. Dr. Dobb’s Journal of Software Tools, 26(2):151, 153, 155, February 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_02/
aa0201.txt.

Gantz:2000:LM

[GF00] Howard Gantz and William Fishburne. Letters: Margination. Dr. Dobb’s Journal of Software Tools, 25(9):10, 12, September 2000. CODEN DDJOEB. ISSN 1044-789X. See [Fis00].

Green:2000:JC

[GH00] Marcus Green and David Hecksel. Java certification. Dr. Dobb’s Journal of Software Tools, 25(10):??, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Gibbons:2001:TDJ

[Gib01] Chad Gibbons. Tape devices and Java. Dr. Dobb’s Journal of Software Tools, 26(2):44, 46, 48–50, February 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.

                                        ddj.com/ftp/2001/2001_02/
                                        javatape.txt;   http://www.
                                        ddj.com/ftp/2001/2001_02/
                                        javatape.zip.

Ginsberg:2000:MCM

[Gin00] Matthew Ginsberg. Monte Carlo methods. Dr. Dobb’s Journal of Software Tools, 25 (4):34, 36–37, April 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_04/
                                        mc.txt.

Ghaly:2001:SEA

[GKM01] Ragae Ghaly, Krishna Kothapalli, and Uma Meyyappan. Selecting EJB application servers: Benchmark and test a variety of EJB servers before making a decision. Dr. Dobb’s Journal of Software Tools, 26 (9):96, 98, 100–101, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Gabeler-Lee:2003:LMM

[GLFO^{+}03] Matthew Gabeler-Lee, Daniel Fremberg, J. G. Owen, Gregory A. Janie, and Mike Moretti. Letters: Mithra; more business process outsourcing; SquareList; XML data binding. Dr. Dobb’s Journal of Software Tools, 28(7):10, July 2003. CODEN DDJOEB. ISSN 1044-789X.

Guy:2002:LFI

[GMTA02] Peter M. Guy, Eric McRae, Andrew W. Todd, and Curt Adalbert. Letters: Fix it; strike one…; strike two. … Dr. Dobb’s Journal of Software Tools, 27(6):10, June 2002. CODEN DDJOEB. ISSN 1044-789X.

Goldman:2000:LDM

[GMW00] Roy Goldman, Jason McHugh, and Jennifer Widom. Lore: A database management system for XML. Dr. Dobb’s Journal of Software Tools, 25(4): 76, 78–80, April 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_04/
lore.txt.

Goeschl:2001:JTT

[Goe01] Siegfried Goeschl. The JUnit++ testing tool. Dr. Dobb’s Journal of Software Tools, 26 (2):34, 36–38, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_02/
junitpp.txt;   http://www.
ddj.com/ftp/2001/2001_02/
junitpp.zip.

Gold:2000:XPI

[Gol00] Rich Gold. Xerox PARC at 30: Inside a research lab. Dr. Dobb’s Journal of Software Tools, 25(12 (supplement)):42– 46, 48, December 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Goldman:2001:JQW

[Gol01] Oliver Goldman. Java Q&A:

What are checked exceptions? Dr. Dobb’s Journal of Software Tools, 26(11):111–112, 114, November 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_11/jqa1101.
                                        txt.

Garbet:2003:LSM

[GOLD03] Shawn Garbet, J. G. Owen, Georg Lohrer, and DDJ. Letters: State of the machine; more on little languages; incorrect values. Dr. Dobb’s Journal of Software Tools, 28 (2):10, February 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7790/
                                        ddj0302m/.

Gong:2000:AAA

[Gon00] Mingfu Gong. Algorithm alley: An N log N adaptive group sort algorithm. Dr. Dobb’s Journal of Software Tools, 25(3):123–124, 126– 127, March 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_03/
                                        aa300.txt.

GonzalezPerez:2001:SOP

[Gon01] César A. Gonzalez Perez. The Strix Object Persistence Engine. Dr. Dobb’s Journal of Software Tools, 26(8):40, 42, 44, 46, 48, August 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_08/
                                        strix.txt.

Gorman:2000:PCT

[Gor00]  Ian E. Gorman.  Pars-

ing complex text structures.

Dr. Dobb’s Journal of Software Tools, 25(6):90, 92– 98, June 2000. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_06/
parse.txt;   http://www.
ddj.com/ftp/2000/2000_06/
parse.zip.

Goschl:2003:JXB

[Gös03] Siegfried Göschl. Jelly: An XML-based scripting language. Dr. Dobb’s Journal of Software Tools, 28 (7):66–67, July 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2003/2003_07/
jelly.zip.

Grinzo:2001:PBB

[GP01] Lou Grinzo and Trudy Pelzer.

Programmer’s bookshelf: Beginning Perl and Windows 2000 Programming. Dr. Dobb’s Journal of Software Tools, 26 (1):197–198, January 2001. CODEN DDJOEB. ISSN 1044-789X.

Grinzo:2001:PBP

[Gri01] Lou Grinzo. Programmer’s bookshelf: Professional Linux Programming. Dr. Dobb’s Journal of Software Tools, 26 (4):151–152, April 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Grinzo:2002:PBC

[Gri02] Lou Grinzo. Programmer’s bookshelf: C# and the .NET platform. Dr. Dobb’s Journal of Software Tools, 27(1): 108, January 2002. CODEN DDJOEB. ISSN 1044-789X.

Grover:2001:SQC

[Gro01] Lov K. Grover. Searching with quantum computers. Dr. Dobb’s Journal of Software Tools, 26(4):34, 36, 38–40, 42–43, April 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_04/quantum.
                                        txt.

Grosso:2002:AOP

[Gro02]  William Grosso.  Aspect-

oriented programming and AspectJ. Dr. Dobb’s Journal of Software Tools, 27(8): ??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_08/aop.txt;
                                         http://www.ddj.com/ftp/
                                        2002/2002_08/aop.zip.

Grinzo:2000:PBG

[GS00] Lou Grinzo and Jacques Surveyer. Programmer’s bookshelf: GNU tools and process patterns. Dr. Dobb’s Journal of Software Tools, 25(12):159– 160, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Gsoedl:2000:JQC

[Gso00] Jacob Gsoedl. Java Q&A: Can you implement COM components using Java? Dr. Dobb’s Journal of Software Tools, 25(6):119–120, 122, 124, 126, June 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_06/jqa600.
txt.

Gsoedl:2001:ASV

[Gso01a] Jacob Gsoedl. Automated streaming video. Dr. Dobb’s Journal of Software Tools, 26 (5):38, 40–48, May 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_05/
video.txt;   http://www.
ddj.com/ftp/2001/2001_05/
video.zip.

Gsoedl:2001:MTB

[Gso01b] Jacob Gsoedl. Managing teambased Web development. Dr. Dobb’s Journal of Software Tools, 26(1):148, 150, 152, 154, 156, January 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_01/webteam.
txt;   http://www.ddj.com/
ftp/2001/2001_01/webteam.
zip.

Gunton:2001:SSD

[Gun01a] Neil Gunton. SOAP: Simplifying distributed development: Putting the Simple Object Access Protocol to work. Dr. Dobb’s Journal of Software Tools, 26(9):89, 91–92, 95, September 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_09/
soap.txt.

Gunton:2001:WSS

[Gun01b] Neil Gunton. Web site searching and indexing in Perl. Dr. Dobb’s Journal of Software Tools, 26(4):106, 108– 109, 111–112, April 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_04/
                                        mysearch.txt;   http://www.
                                        ddj.com/ftp/2001/2001_04/
                                        mysearch.zip.

Gunion:2002:NVC

[Gun02] Robert Gunion. .NET versus COM. Dr. Dobb’s Journal of Software Tools, 27(10):49–52, 54, 56, 58, October 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_10/
                                        netvscom.txt;   http://www.
                                        ddj.com/ftp/2002/2002_10/
                                        netvscom.zip.

Gutman:2000:AAE

[Gut00] Ron Gutman. Algorithm alley: Exploiting 64-bit parallelism. Dr. Dobb’s Journal of Software Tools, 25(9):133–134, 136, September 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_09/
                                        aa0900.txt.

Gutman:2002:AAP

[Gut02] Ron Gutman. Algorithm alley: Priority queues for motorists. Dr. Dobb’s Journal of Software Tools, 27(9):89–90, 92–94, September 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_09/
aa0902.txt.

Gohel:2001:LGK

[GWT^{+}01] Himanshu Gohel, Thomas Weidenfeller, Andrew D. Todd, Ze’ev Atlas, and Dave Paris. Letters: GNOME, KDE clarification; parsing XML; CPRM frenzy; deadlock detecting; Perl versus the world. Dr. Dobb’s Journal of Software Tools, 26(5):10, 12, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Haggar:2002:JQD

[Hag02] Peter Haggar. Java Q&A: Does Java guarantee thread safety? Dr. Dobb’s Journal of Software Tools, 27(6):91–83, June 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2002/2002_06/
jqa0602.txt. Comments on

lack of atomic-update guarantee in Java for objects larger than 32 bits, such as long and double, with sample code to exhibit the failure.

Hallerdal:2000:WII

[Hal00]  Martin Hallerdal.  Web-

DAV, IIS, and ISAPI filters. Dr. Dobb’s Journal of Software Tools, 25(11):114, 116– 119, November 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_11/
webdav.txt;   http://www.
ddj.com/ftp/2000/2000_11/
webdav.zip.

Hardin:2000:RTS

[Har00]  David Hardin.  The real-

time specification for Java. Dr. Dobb’s Journal of Software Tools, 25(2):78, 80–82, 84, February 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_02/rtjava.
                                        txt.

Havewala:2000:AB

[Hav00] Aspi Havewala. Automated builds. Dr. Dobb’s Journal of Software Tools, 25(8):60, 62– 66, 68, August 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_08/
                                        autbuild.txt.

Hadzic:2000:LGA

[HBT^{+}00] Mirza Hadzic, Phil Bagwell, Andrew D. Todd, Berney Williams, and Murray Lesser. Letters: Graphics algorithms; exploiting 64-bit parallelism; worker shortage; computer history; analog versus digital. Dr. Dobb’s Journal of Software Tools, 25(11):10, 12, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Heijl:2001:DXS

[Hei01] Danny Heijl. The Delphi XML SAX2 component and MSXML 3.0. Dr. Dobb’s Journal of Software Tools, 26 (9):42, 46, 48, 50, 52, 54, September 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/

ftp/2001/2001_09/xmlsax2.
txt;   http://www.ddj.com/
ftp/2001/2001_09/xmlsax2.

zip. See correction [TEM^{+}01].

Heller:2002:PT

[Hel02] Martin Heller. Programmer’s toolchest: Examining Visual Studio .NET. Dr. Dobb’s Journal of Software Tools, 27 (7):71–73, July 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_07/
vsnet.txt;   http://www.
ddj.com/ftp/2002/2002_07/
vsnet.zip.

Herzberg:2002:SX

[Her02]  Amir Herzberg.  Securing

XML. Dr. Dobb’s Journal of Software Tools, 27(3):56, 59–62, March 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_03/
secxml.txt.

Hassler:2000:OFA

[HF00] Vesna Hassler and Oliver Fodor. OpenCard Framework application development. Dr. Dobb’s Journal of Software Tools, 25(2):70, 72, 74– 76, February 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_02/ocfjava.
txt;   http://www.ddj.com/
ftp/2000/2000_02/ocfjava.
zip.

Hightower:2001:EFT

[Hig01] Lauren Hightower. Examining the FDF toolkit. Dr. Dobb’s Journal of Software Tools, 26(8):87, 89–90, 92, 94– 95, August 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_08/fdf.txt;
                                         http://www.ddj.com/ftp/
                                        2001/2001_08/fdf.zip.

Hightower:2002:PTE

[Hig02] Lauren Hightower. Programmer’s toolchest: Examining VB.NET: A popular toolset is reengineered. Dr. Dobb’s Journal of Software Tools, 27 (3):65–66, 68–69, March 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2002/2002_03/
                                        vbnet.txt.

Hinke:2002:ICS

[Hin02]  John Hinke.  Implement-

ing C++ servlet containers. Dr. Dobb’s Journal of Software Tools, 27(4):52, 54– 56, April 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_04/cppserv.
                                        txt.

Henry:2000:JQH

[HL00] Ethan Henry and Ed Lycklama. Java Q&A: How do you plug Java memory leaks? Dr. Dobb’s Journal of Software Tools, 25(2):115– 119, 121, February 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_02/
jqa220.txt.

Haines:2000:TIT

[HM00] Eric Haines and Tomas Möller.

Triangle intersection tests. Dr. Dobb’s Journal of Software Tools, 25(8):32, 34, 36, August 2000. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/2000/2000_08/tri_
test.rtf;   http://www.ddj.
com/ftp/2000/2000_08/tri_
test.zip.

Hultquist:2001:AA

[HM01] Paul F. Hultquist and William R.

Mahoney. Algorithm alley: Reservoir sampling. Dr. Dobb’s Journal of Software Tools, 26(1):189–190, January 2001. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_01/
aa0101.txt;   http://www.
ddj.com/ftp/2001/2001_01/
aa0101.zip.

Houlding:2000:PSC

[Hou00] David Houlding. Publish and subscribe with CORBA Web events. Dr. Dobb’s Journal of Software Tools, 25(7):88, 90–93, July 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_07/pscorba.
txt;   http://www.ddj.com/
ftp/2000/2000_07/pscorba.
zip.

Houlding:2001:VVD

[Hou01] David Houlding. VoiceXML and the voice-driven Internet. Dr. Dobb’s Journal of Software Tools, 26(4):88, 90– 92, 94–95, April 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_04/
                                        voicexml.zip;   http://www.
                                        ddj.com/ftp/2001/2001_04/
                                        voicexmltxt.zip.

Hovel:2002:ANP

[Hov02] David Hovel. ASP.NET page persistence and extended attributes. Dr. Dobb’s Journal of Software Tools, 27(12):30, 32, 34–36, December 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7718/
                                        ddj0212d/.

Howard:2001:OOD

[How01]  Michael Howard.  Object-

oriented device networking. Dr. Dobb’s Journal of Software Tools, 26(8):63, 66– 69, August 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_08/
                                        oodevice.txt.

Hume:2001:SMS

[Hum01] Ron Hume. Short message services. Dr. Dobb’s Journal of Software Tools, 26(10):62, 64– 68, 70, October 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.

com/ftp/2001/2001_10/sms.
txt.

Humphrey:2003:GCC

[Hum03] William F. Humphrey. Generalized callbacks: C++ and C#. Dr. Dobb’s Journal of Software Tools, 28(3):42–43, 46–47, March 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7826/
ddj0303e/.

Herzberg:2001:LSR

[HWT^{+}01] Amir Herzberg, Jerry Whiting, Andrew D. Todd, Jose Melo de Assis Fonseca, Eddie Edwards, and Øyvind Teig. Letters: SET realities and Cue-Cat and C++ namespaces and C++ aliasing. Dr. Dobb’s Journal of Software Tools, 26 (1):10, 12, January 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Horwood:2000:DMA

[HWZ00] Peter Horwood, Shlomo Wygodny, and Martin Zardecki. Debugging multithreaded applications. Dr. Dobb’s Journal of Software Tools, 25(3): 32, 34–37, March 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_03/
dbgmulti.txt.

Ironside:2001:ECI

[IEZ01] Eric Ironside, Letha Etzkorn, and David Zajac. Examining CORBA interoperability.

Dr. Dobb’s Journal of Software Tools, 26(6):111, 113–114, 116–117, 120, 122, June 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_06/
corba.txt.

Ishaq:2001:PML

[Ish01] Jeff Ishaq. Palm OS memory leak detection. Dr. Dobb’s Journal of Software Tools, 26(7):62, 66, 68, 70, July 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_07/palmmem.
txt;   http://www.ddj.com/
ftp/2001/2001_07/palmmem.
zip.

Israel:2001:ACR

[Isr01] Lillian Israel. The ACM Career Resource Center. Dr. Dobb’s Journal of Software Tools, 26 (4):??, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Jackson:2001:JQW

[Jac01] Jared Jackson. Java Q&A: What are resource bundles and what do they have to do with internationalization? Dr. Dobb’s Journal of Software Tools, 26(9):117–119, September 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_09/jqa0901.
txt.

Jackson:2002:WAM

[Jac02]  L. Blunt Jackson.  Writ-

ing Apache modules. Dr.

Dobb’s Journal of Software Tools, 27(9):42, 44–48, 50, September 2002. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_09/
apachmod.txt.

Janik:2001:MDG

[Jan01a] David Janik. Module design guidelines for real-time systems. Dr. Dobb’s Journal of Software Tools, 26(2):108, 110– 114, 117, February 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_02/
module.txt.

Janulaitis:2001:CHU

[Jan01b] Michael Janulaitis. Creating HTML user interfaces for server programs. Dr. Dobb’s Journal of Software Tools, 26(6):101–102, 104, 106, 108–109, June 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_06/
htmlui.txt;   http://www.
ddj.com/ftp/2001/2001_06/
htmlui.zip.

Jarp:2001:OIP

[Jar01] Sverre Jarp. Optimizing IA-64 performance. Dr. Dobb’s Journal of Software Tools, 26(7): 21–22, 24, 26, July 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_07/
ia64.txt;   http://www.ddj.
com/ftp/2001/2001_07/ia64.
zip.

Jennings:2000:JQC

[Jen00a] Mike Jennings. Java Q&A: Can you write NT services in Java? Dr. Dobb’s Journal of Software Tools, 25(3):113– 116, 118, March 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_03/
                                        jqa300.txt;   http://www.
                                        ddj.com/ftp/2000/2000_03/
                                        jqa300.zip.

Jennings:2000:JQH

[Jen00b] Mike Jennings. Java Q&A: How do I establish a net connection? Dr. Dobb’s Journal of Software Tools, 25(11):141– 142, 144–147, November 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2000/2000_11/
                                        jqa0011.txt;   http://www.
                                        ddj.com/ftp/2000/2000_11/
                                        jqa0011.zip.

Jenkins:2002:GJP

[Jen02a] Stephen B. Jenkins. Generating JavaScript from Perl. Dr. Dobb’s Journal of Software Tools, 27(5):59–63, May 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2002/2002_05/
                                        jscript.txt.

Jennings:2002:JQ

[Jen02b] Mike Jennings. Java Q&A: How do I use the Java plugin tool? Dr. Dobb’s Journal of Software Tools, 27(1): 93–94, January 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_01/jqa0102.
zip.

Johnson:2000:DSC

[Joh00] Steve Johnson. A dynamic select component for JavaScript. Dr. Dobb’s Journal of Software Tools, 25(1):36, 38– 42, January 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_01/dselect.
txt;   http://www.ddj.com/
ftp/2000/2000_01/dselect.
zip.

Jolin:2001:JQC

[Jol01] Art Jolin. Java Q&A: Can J2EE manage cache and pool memory? Dr. Dobb’s Journal of Software Tools, 26(10):117– 120, October 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_10/jqa1001.
txt;   http://www.ddj.com/
ftp/2001/2001_10/jqa1001.
zip.

Jolitz:2003:WSD

[Jol03] William F. Jolitz. Web services and datacenter environments. Dr. Dobb’s Journal of Software Tools, 28(4):60, 62–65, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7827/ddj0304g/.

Jones:2001:EHS

[Jon01] M. Tim Jones. An embeddable HTTP server. Dr. Dobb’s Journal of Software Tools, 26(10):73–74, 77–79, October 2001. CODEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_10/
                                        emhttp.txt;   http://www.
                                        ddj.com/ftp/2001/2001_10/
                                        emhttp.zip.

Jones:2002:JMA

[Jon02a] M. Tim Jones. Java mobile agents and the Aglets SDK. Dr. Dobb’s Journal of Software Tools, 27(1):42, 44, 46–48, January 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_01/
                                        aglet.txt.

Jones:2002:PAR

[Jon02b] M. Tim Jones. Personalization and adaptive resonance theory. Dr. Dobb’s Journal of Software Tools, 27(10): 36–39, October 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_10/
                                        aa1002.zip.

Jones:2003:ELX

[Jon03] M. Tim Jones. An embeddable lightweight XML-RPC server. Dr. Dobb’s Journal of Software Tools, 28(6):60, 62–64, 66–67, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
8213/ddj0306g/.

Jordi:2001:VM

[Jor01] Steve Jordi. Volcano monitoring. Dr. Dobb’s Journal of Software Tools, 26(3):64, 66–70, March 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_03/volcano.
txt;   http://www.ddj.com/
ftp/2001/2001_03/volcano.
zip.

Jin:2001:WCS

[JRFH01] Liang Jin, Shi Ren, Liang Feng, and Gao Zheng Hua. WAP clients and set protocol. Dr. Dobb’s Journal of Software Tools, 26(6):85, 87– 89, 91, June 2001. CODEN DDJOEB. ISSN 1044-789X.

Kaliski:2001:RDS

[Kal01] Burton S. Kaliski, Jr. RSA digital signatures. Dr. Dobb’s Journal of Software Tools, 26 (5):30, 32–33, 35–36, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Kanalakis:2002:WSJ

[Kan02] John M. Kanalakis, Jr. Web services and Java server pages. Dr. Dobb’s Journal of Software Tools, 27(1):28, 30– 34, January 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_01/
wsjsp.txt;   http://www.
ddj.com/ftp/2002/2002_01/
wsjsp.zip.

Kaplan:2000:FBS

[Kap00] John Kaplan. Finding the best software developers. Dr. Dobb’s Journal of Software Tools, 25(10):??, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Karim:2000:WNC

[Kar00] Jawed Karim. A Win32 network crawler. Dr. Dobb’s Journal of Software Tools, 25(5): 38, 40, 42–43, May 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_05/
                                        voyeur.txt;   http://www.
                                        ddj.com/ftp/2000/2000_05/
                                        voyeur.zip.

Katseff:2003:WSP

[Kat03] Howard P. Katseff. Web scraping proxy. Dr. Dobb’s Journal of Software Tools, 28(6): 46, 48, 50, 52, June 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=8213/
                                        ddj0306e/.

Konchady:2002:ISS

[KD02] Manu Konchady and Ray D’Amore. Implementing a smart spider. Dr. Dobb’s Journal of Software Tools, 27 (8):??, August 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_08/
                                        spider.zip.

Kiselyov:2000:LZA

[KDG^{+}00] Oleg Kiselyov, Charles Dorcey, Ron Gutman, Eric Haines, Tomas Möller, Peter Varhol, Edward Harned, and M. Jojje. Letters: Zuse accolades; exploiting 64-bit parallelism; triangle correction; computer science versus programming; Microsoft made its own bed; client/server tip. Dr. Dobb’s Journal of Software Tools, 25(12):10–12, December 2000. CODEN DDJOEB. ISSN 1044-789X. See [].

Konolige:2000:RCC

[KDK00] Kurt Konolige, Jeanne Dietsch, and William Kennedy. Robot control and Colbert. Dr. Dobb’s Journal of Software Tools, 25(11):90, 94, 96, 98, November 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_11/
robot.txt.

Kelley:2003:GC

[Kel03] Brian Kelley. Graph canonicalization. Dr. Dobb’s Journal of Software Tools, 28(5):66–69, May 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7927/ddj0305g/.

Keogh:2000:AAL

[Keo00] John Keogh. Algorithm alley: Layout management. Dr. Dobb’s Journal of Software Tools, 25(6):131–134, 136, June 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_06/
aa600.txt;   http://www.
ddj.com/ftp/2000/2000_06/
aa600.zip.

Kientzle:2000:DEM

[Kie00a] Tim Kientzle. Database engines: MySQL versus Oracle. Dr. Dobb’s Journal of Software Tools, 25(7):98, 100, 102, 104, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
mysqlorl.txt.

Kiesling:2000:OOU

[Kie00b] Robert Kiesling. An objectoriented UI for Perl. Dr. Dobb’s Journal of Software Tools, 25(6):58, 60–61, June 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ftp/2000/2000_
06/perlooui.txt;   http:/
/www.ddj.com/ftp/2000/2000_
06/worksp.zip.

Kientzle:2001:JQH

[Kie01a] Tim Kientzle. Java Q&A: How can I speed up my JDBC-based programs? Dr. Dobb’s Journal of Software Tools, 26(12):111– 112, December 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_12/jqa1201.
txt;   http://www.ddj.com/
ftp/2001/2001_12/jqa1201.
zip.

Kiesling:2001:MR

[Kie01b] Robert Kiesling. E-mail relaying: Coping with the sendmail and qmail mail transport agents. Dr. Dobb’s Journal of Software Tools, 26(11):87– 91, November 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_11/
relay.txt;   http://www.
ddj.com/ftp/2001/2001_11/
relay.zip.

Kiesling:2001:EPL

[Kie01c] Robert Kiesling. Exploring Perl libraries. Dr. Dobb’s Journal of Software Tools, 26(2): 84, 86–88, February 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_02/
perllib.txt;   http://www.
ddj.com/ftp/2001/2001_02/
perllib.zip.

Kientzle:2002:JQH

[Kie02a] Tim Kientzle. Java Q&A: How can I make my Java servlets faster? Dr. Dobb’s Journal of Software Tools, 27(4): 97–99, April 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_04/jqa0402.
txt;   http://www.ddj.com/
ftp/2002/2002_04/jqa0402.
zip.

Kiesling:2002:OUE

[Kie02b] Robert Kiesling. ODBC in Unix environments. Dr. Dobb’s Journal of Software Tools, 27 (12):16, 18–20, 22, December 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7718/ddj2012b/.

Kim:2000:CDE

[Kim00a] Eugene Eric Kim. A conversation with Doug Engelbart. Dr. Dobb’s Journal of Software Tools, 25(9):21–22, 24, 26, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Kim:2000:FP

[Kim00b] Eugene Eric Kim. The future of programming. Dr. Dobb’s Journal of Software Tools, 25 (12 (supplement)):6–9, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Kim:2001:PPR

[Kim01a] Eugene Eric Kim. Programming and the PC revolution. Dr. Dobb’s Journal of Software Tools, 26(1):21–22, 24, 30, 32, 34, 36, 38, January 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Kim:2001:TSJ

[Kim01b] Eugene Eric Kim. A triumph of simplicity: James Clark on markup languages and XML. Dr. Dobb’s Journal of Software Tools, 26(7):56, 58–60, July 2001. CODEN DDJOEB. ISSN 1044-789X.

Kim:2002:LLW

[Kim02] Eugene Eric Kim. The MIT Lightweight Languages Workshop. Dr. Dobb’s Journal of Software Tools, 27(2):21– 22, 24, 26, February 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/technetcast.ddj.com/tnc_
catalog.html?item_id=1295.

Kiselyov:2000:LSI

[Kis00] Oleg Kiselyov. Letters: Smart iterators. Dr. Dobb’s Journal of Software Tools, 25(9): 10, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Kiselev:2001:PDE

[Kis01] Ivan Kiselev. A portable distributed event-logging facility: Monitoring the state of applications. Dr. Dobb’s Journal of Software Tools, 26(9):26, 28– 30, September 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_09/logging.
txt;   http://www.ddj.com/
ftp/2001/2001_09/logging.
zip.

Kopp:2000:MCC

[Kop00] Carlo Kopp. Managing cluster computers. Dr. Dobb’s Journal of Software Tools, 25(7): 21–26, July 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_07/cluster.
txt.

Kougiouris:2001:PMP

[Kou01] Panos Kougiouris. Programming mobile phones. Dr. Dobb’s Journal of Software Tools, 26(9):73, 75, 77, 78– 79, September 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_09/mphones.
txt.

Kripac:2001:AAE

[Kri01] Jiri Kripac. Algorithm alley: An efficient undo/redo algorithm. Dr. Dobb’s Journal of Software Tools, 26(11): 117–119, 122–123, November 2001. CODEN DDJOEB. ISSN 1044-789X.

Kautz:2000:LLI

[KSC^{+}00] Frederick Kautz, Dimitrios Souflis, Robert Carbonari, DDJ, Al Gettier, Tina Femia, Andrew D. Todd, Anonymous, Alan Horn, Anonymous, Brad Clawsie, and Samuel Pearlman. Letters: The S language; IP security; Pnuts pointer; flyby; worker shortage?; design by interface. Dr. Dobb’s Journal of Software Tools, 25(4): 10, 12, April 2000. CODEN DDJOEB. ISSN 1044-789X. See [Ste00k].

Lacas:2001:VVN

[Lac01] Mark Lacas. VNOS: The Visual Network Operating System. Dr. Dobb’s Journal of Software Tools, 26(12):89– 93, December 2001. CODEN DDJOEB. ISSN 1044-789X.

Laird:2001:CPT

[Lai01a] Cameron Laird. Compiling Perl/Tk scripts. Dr. Dobb’s Journal of Software Tools, 26 (5):60, 63, 65–66, May 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_05/
perltk.txt.

Laird:2001:GCP

[Lai01b] Cameron Laird. GUI construction with Perl. Dr. Dobb’s Journal of Software Tools, 26 (1):80, 82, 84, 86, January 2001. CODEN DDJOEB. ISSN 1044-789X.

Laird:2001:JQW

[Lai01c] Cameron Laird. Java Q&A:

When should you script Java with TCL? Dr. Dobb’s Journal of Software Tools, 26(3): 133–134, March 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_03/jqa0301.
txt.

Lardieri:2001:ACE

[Lar01a] Steve Lardieri. ActiveX controls for Embedded Visual Basic. Dr. Dobb’s Journal of Software Tools, 26(7):72, 76, 78– 84, 86, July 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_07/embedvb.
txt;   http://www.ddj.com/
ftp/2001/2001_07/embedvb.
zip.

Larson:2001:WBV

[Lar01b]  Michael Larson.  Web-

based video monitoring. Dr. Dobb’s Journal of Software Tools, 26(7):88, 90, 93– 94, 96, July 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_07/
                                        video.txt;   http://www.
                                        ddj.com/ftp/2001/2001_07/
                                        video.zip.

Lardieri:2003:PDW

[Lar03a] Stephen Lardieri. Private desktops and Windows XP. Dr. Dobb’s Journal of Software Tools, 28(1):26, 28, 30, 32, January 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7758/ddj0301c/.

Larson:2003:RCS

[Lar03b] Michael Larson. Route control and streaming video. Dr. Dobb’s Journal of Software Tools, 28(6):18, 20, 22, 24, 26, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        8213/ddj0306b/.

Latteier:2002:ZPT

[Lat02] Amos Latteier. Zope page templates. Dr. Dobb’s Journal of Software Tools, 27(2):67, 69– 70, 73–75, February 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_02/
                                        zope.zip.

Ledin:2000:DFO

[Led00] Jim Ledin. Digital filtering and oversampling. Dr. Dobb’s Journal of Software Tools, 25(4):66, 68–70, 72– 74, April 2000. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_04/
filter.txt;   http://www.
ddj.com/ftp/2000/2000_04/
filter.zip.

Ledin:2001:RTS

[Led01]  Jim Ledin.  The Real-

Time Simulation Protocol. Dr. Dobb’s Journal of Software Tools, 26(5):83, 86–88, 90, 91, May 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_05/
rtsp.txt;   http://www.ddj.
com/ftp/2001/2001_05/rtsp.
zip.

Lee:2001:HJI

[Lee01] Katherine Spencer Lee. Hot jobs for IT professionals in 2001. Dr. Dobb’s Journal of Software Tools, 26(4):??, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Levitin:2000:NRM

[Lev00] Anany Levitin. A new road map of algorithm design techniques. Dr. Dobb’s Journal of Software Tools, 25(4):23–24, 26–28, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Li:2003:JBM

[Li03a] Sing Li. A Java-based music player for MP3, Ogg, and WAV. Dr. Dobb’s Journal of Software Tools, 28(2):38, 40, 42, 44–45, February 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7790/
                                        ddj0302e/.

Li:2003:JPP

[Li03b] Sing Li. JXTA and peer-to-peer networks. Dr. Dobb’s Journal of Software Tools, 28 (6):30, 32–34, June 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=8213/
                                        ddj0306c/.

Link:2001:TES

[Lin01] Joseph M. Link. Trampolines for embedded systems: Minimizing interrupt handlers latency. Dr. Dobb’s Journal of Software Tools, 26(9):81, 84– 87, September 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_09/
                                        tramp.txt.

Linwood:2002:PT

[Lin02] Jeff Linwood. Programmer’s toolchest. Dr. Dobb’s Journal of Software Tools, 27 (8):??, August 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_08/
                                        jmeter.txt;   http://www.
                                        ddj.com/ftp/2002/2002_08/
                                        jmeter.zip.

Lischner:2002:BCC

[Lis02]  Ray Lischner.  Borland’s

CLX component framework.

Dr. Dobb’s Journal of Software Tools, 27(2):77, 79–81, February 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_02/clx.txt.

Larsson:2001:DS

[LJ01] Andreas Larsson and Henrik Jeppsson. Designing 3G systems. Dr. Dobb’s Journal of Software Tools, 26(5):68, 70– 73, 75, 77–78, May 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_05/
3g.txt.

Lamper:2002:SC

[LJ02] David Lamper and Neil F.

Johnson. The science of complexity. Dr. Dobb’s Journal of Software Tools, 27(10):16, 18– 20, 22, October 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_10/complex.
txt;   http://www.ddj.com/
ftp/2002/2002_10/complex.
zip.

Laurie:2001:LAA

[LMM^{+}01] Ben Laurie, James Metzger, Maury Markowitz, Edward Harnsed, Matthew Saltzman, Claude Brown, and Robert Wuest. Letters: Analyzing algorithms; child’s play; Hurd, Hurd, Hurd — Hurd is the word; software in the 21st Century; analyzing analytic computing; giving Boost a boost; MP3 notes. Dr. Dobb’s Journal of Software Tools, 26(2):10, 12, February 2001. CODEN DDJOEB. ISSN 1044-789X.

Leibundguth:2001:UMD

[LPP01] Jonathan Leibundguth, Michael Pease, and Andrei Povodyrev. A universal model for designing the entity EJB layer. Dr. Dobb’s Journal of Software Tools, 26(6):38, 40–43, 46, June 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_06/ejb.txt;
                                         http://www.ddj.com/ftp/
                                        2001/2001_06/ejb.zip.

Lach-Smith:2001:PBD

[LS01] Barbara Lach-Smith. Programmer’s bookshelf: Developing for wireless systems. Dr. Dobb’s Journal of Software Tools, 26(10):130–131, October 2001. CODEN DDJOEB. ISSN 1044-789X.

Ma:2001:CPD

[Ma01] Charlie Ma. Cross-platform DHTML. Dr. Dobb’s Journal of Software Tools, 26(2):120, 122–128, February 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_02/
                                        dhtml.txt.

Macri:2000:SPC

[Mac00] Julian Macri. State patterns and C++. Dr. Dobb’s Journal of Software Tools, 25(6):36, 38, 40, 42, 44–45, June 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.

ddj.com/ftp/2000/2000_06/
statepat.txt.

Maddock:2001:REC

[Mad01]  John Maddock.  Regular

expressions in C++. Dr. Dobb’s Journal of Software Tools, 26(10):21–22, 24, 26, October 2001. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_10/
regexpp3.txt;   http://www.
ddj.com/ftp/2001/2001_10/
regexpp3.zip.

Mahoney:2002:AAY

[Mah02] William R. Mahoney. Algorithm alley: Yet another record selection algorithm. Dr. Dobb’s Journal of Software Tools, 27(2):98– 99, February 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_02/aa0202.
txt.

Mahoney:2003:APC

[Mah03] William R. Mahoney. The “allpairs closest points” problem. Dr. Dobb’s Journal of Software Tools, 28(1):48–49, 51, January 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7758/ddj0301f/.

Manivannan:2000:EAM

[Man00] S. Manivannan. An efficient algorithm for magnitude operation. Dr. Dobb’s Journal of Software Tools, 25(4):

54, 565, April 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ftp/2000/2000_04/mag.
                                        txt.

Manley:2002:MWA

[Man02] Kevin T. Manley. Monitoring Web application performance and ISAPI. Dr. Dobb’s Journal of Software Tools, 27 (7):65–68, July 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_07/
                                        isapi.txt;   http://www.
                                        ddj.com/ftp/2002/2002_07/
                                        isapi.zip.

Marrie:2000:AAA

[Mar00]  Laurence Marrie.  Algo-

rithm alley: Alternating skip lists. Dr. Dobb’s Journal of Software Tools, 25(8):133– 136, August 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_08/
                                        aa800.zip.

Marasco:2002:TSD

[Mar02a] Joe Marasco. Tracking software development projects. Dr. Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Marques:2002:BSJ

[Mar02b] Paulo Marques. Building secure Java RMI servers. Dr. Dobb’s Journal of Software Tools, 27(11):36, 38, 40–42, 44, November 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7644/
ddj0211d/.

Martin:2003:CBM

[Mar03] Ken Martin. The CMake build manager. Dr. Dobb’s Journal of Software Tools, 28(1):40, 42, 44, 46, January 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7758/
ddj0301e/.

Masta:2001:LIL

[Mas01] Robert Masta. Letters: Interrupt latencies. Dr. Dobb’s Journal of Software Tools, 26 (4):10, 12, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Massa:2002:IGW

[Mas02] Anthony J. Massa. Integrating GoAhead WebServer and eCos. Dr. Dobb’s Journal of Software Tools, 27(11):56, 58, 60, 62, 64, November 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7644/
ddj0211f/.

Maddock:2000:CTT

[MC00] John Maddock and Steve Cleary. C++ type traits. Dr. Dobb’s Journal of Software Tools, 25(10):38, 40, 42, 44, October 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_10/
cpptt.txt.

McCoy:2000:SP

[McC00] John H. McCoy. Scripting for Pnuts. Dr. Dobb’s Journal of Software Tools, 25(1):21–22, 24, 26, January 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_01/
                                        pnutdemo.zip;   http://www.
                                        ddj.com/ftp/2000/2000_01/
                                        pnuts.txt.

McClellen:2002:WDD

[McC02] Chris McClellen. Weather data distribution and system monitoring: An opensource ORB for a rainy day. Dr. Dobb’s Journal of Software Tools, 27(1):69–70, 72– 74, January 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_01/weather.
                                        txt.

McKenzie:2001:JQJ

[McK01]  Nadine McKenzie.  Java

Q&A: Is JavaScript an objectoriented language? Dr. Dobb’s Journal of Software Tools, 26 (8):115–116, August 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_08/
                                        jqa0801.txt.

McMillan:2000:PTE

[McM00] Mike McMillan. Programmer’s toolchest: Examining Perl-COM: Perl meets Microsoft’s COM. Dr. Dobb’s Journal of Software Tools, 25(1):84, 86– 87, January 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_01/perlcom.
txt.

McRae:2002:TKB

[McR02] Eric McRae. Tracking down killer bugs. Dr. Dobb’s Journal of Software Tools, 27(4):58, 60–61, 63–64, April 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Metzger:2001:UWN

[Met01] James Metzger. Using Windows NT Embedded 4.0. Dr. Dobb’s Journal of Software Tools, 26(4):98, 100, 102, 104– 105, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Martin:2002:CLM

[MHG02] Ken Martin, William Hoffman, and Berk Geveci. Creating libraries for multiple programming languages. Dr. Dobb’s Journal of Software Tools, 27 (2):54, 56–59, February 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2002/2002_02/
multilan.txt;   http://www.
ddj.com/ftp/2002/2002_02/
multilan.zip.

Milewski:2002:LDT

[MHM^{+}02] Bartosz Milewski, Jeff Hubbach, Richard Mickelsen, Patrick Traill, Douglas Lyon, and J. G. Owen. Letters: Disk thrashing; equals update; discarding information; padded cells and plumbers; the lightweight languages workshop; little languages. Dr. Dobb’s Journal of Software Tools, 27(7):10, July 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Michaels:2000:PC

[Mic00] Laura Michaels. POP3 for C++. Dr. Dobb’s Journal of Software Tools, 25(12):109– 110, December 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_12/cpppop3.
                                        zip.

Milewski:2000:WCI

[Mil00] Bartosz Milewski. The Windows 2000 content index. Dr. Dobb’s Journal of Software Tools, 25(12):40, 42, 44–46, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Milewski:2002:DTP

[Mil02a] Bartosz Milewski. Disk thrashing and the pitfalls of virtual memory. Dr. Dobb’s Journal of Software Tools, 27(5):34, 36, 38–40, May 2002. CODEN DDJOEB. ISSN 1044-789X.

Miller:2002:DPU

[Mil02b] Byron Miller. Determining processor utilization. Dr. Dobb’s Journal of Software Tools, 27(7):56–57, 60–62, July 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/ftp/2002/2002_
                                        07/processr.txt.

Minnich:2001:PNL

[Min01] Ronald G. Minnich. Private namespaces for Linux. Dr. Dobb’s Journal of Software Tools, 26(12):23–24, 26, 28, 30, December 2001. CODEN DDJOEB. ISSN 1044-789X.

Missaghi:2000:AA

[Mis00]  Moheb Missaghi.  Algo-

rithm alley: How many subscribers should share a modem? Dr. Dobb’s Journal of Software Tools, 25(5): 119–121, May 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_05/
aa500.zip.

Mitchell:2000:TBA

[Mit00] Mark Mitchell. Type-based alias analysis. Dr. Dobb’s Journal of Software Tools, 25 (10):54, 56, 58–60, October 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ftp/2000/2000_
10/alias.txt.

Mitzenmacher:2002:AAU

[Mit02a] Michael Mitzenmacher. Algorithm alley: Unbiasing random bits. Dr. Dobb’s Journal of Software Tools, 27(4): 101–104, April 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_04/
aa0402.txt;   http:/www.
ciphergoth.org/software/
unbiasing.

Mitzenmacher:2002:GHT

[Mit02b] Michael Mitzenmacher. Good hash tables and multiple hash functions. Dr. Dobb’s Journal of Software Tools, 27(5): 28, 30–32, May 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_05/
hash.txt.

Martin:2001:LXC

[MKR^{+}01] Larry Martin, Eugene Kim, Toby Reyelts, Al Stevens, Rob Chamberlin, Michael Brinkley, Michael Wojcik, and Jeff Duntemann. Letters: XML and the 21st Century; passing the C++ test; Ruby, Ruby; PHP and hosted applications; shared libraries aren’t all bad. Dr. Dobb’s Journal of Software Tools, 26(3): 10, 12, March 2001. CODEN DDJOEB. ISSN 1044-789X.

Malloy:2002:PTT

[MLDP02] Brian A. Malloy, Scott A. Linde, Edward B. Duffy, and James F. Power. Programmer’s toolchest: Testing C and C++ compilers for ISO language conformance. Dr. Dobb’s Journal of Software Tools, 27(6):71– 74, 76–78, June 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_06/
isocpp.txt;   http://www.
ddj.com/ftp/2002/2002_06/
isocpp.zip.

Michelone:2001:UP

[MM01] Manuel Lopez Michelone and Marcelo Perez Medel. Understanding photomosaics. Dr. Dobb’s Journal of Software Tools, 26(11):58, 60– 63, November 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_11/
mosaic.zip.

Munro:2000:UHJ

[MN00] Charlie Munro and Mark Nelson. The ultimate home jukebox. Dr. Dobb’s Journal of Software Tools, 25(1):76, 78– 82, January 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_01/jukebox.
txt.

Modi:2001:DLC

[Mod01] Tarak Modi. Dynamic logging and the CORBA Notification Service. Dr. Dobb’s Journal of Software Tools, 26(3):42, 44, 46–49, March 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_03/
dynlog.txt;   http://www.
ddj.com/ftp/2001/2001_03/
dynlog.zip.

Mondal:2000:LBU

[Mon00a] Abdul Sakib Mondal. Load balancing for UNIX and Win32. Dr. Dobb’s Journal of Software Tools, 25(7):32, 34, 36, 38–40, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
load.txt;   http://www.ddj.
com/ftp/2000/2000_07/load.
zip;   http://www.ddj.com/
ftp/2000/2000_07/xyalb.zip.

Mondal:2000:PAU

[Mon00b] Abdul Sakib Mondal. Porting across UNIX and Win32. Dr. Dobb’s Journal of Software Tools, 25(12):30, 32, 34, 36, December 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_12/
portuw.txt;   http://www.
ddj.com/ftp/2000/2000_12/
portuw.zip.

Moss:2000:JQ

[Mos00] Karl Moss. Java Q&A: How can I measure Java code performance? Dr. Dobb’s Journal of Software Tools, 25 (10):135–136, 138, 140–142, 144, October 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_10/
jqa010.txt.

Mount:2000:ADP

[Mou00] John Mount. Automatic detection of potential deadlock. Dr. Dobb’s Journal of Software Tools, 25(12):64, 66–70, 72, December 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_12/
deadlock.txt;   http://www.

ddj.com/ftp/2000/2000_12/
deadlock.zip.

Mickelsen:2002:LDI

[MPG^{+}02] Richard Mickelsen, Justin Perkins, Tom Groves, Patrick Traill, and Jason But. Letters: Discarding information; ASP.NET; the D language; computer viruses, padded cells, and plumbers; SocketCC update. Dr. Dobb’s Journal of Software Tools, 27(9):10, September 2002. CODEN DDJOEB. ISSN 1044-789X.

Muffke:2001:CPE

[Müf01] Friedger Müffke. The Curl programming environment: Making the Internet interactive. Dr. Dobb’s Journal of Software Tools, 26(9):66, 68, 70– 71, September 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_09/
curllist.zip.

Musayev:2001:SSA

[Mus01] Eldar A. Musayev. SAX2:

The Simple API for XML. Dr. Dobb’s Journal of Software Tools, 26(2):130, 132– 133, February 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_02/
sax2.txt;   http://www.ddj.
com/ftp/2001/2001_02/sax2.
zip.

Metzger:2000:PC

[MW00] James Metzger and William Wright. Portability in C.

Dr. Dobb’s Journal of Software Tools, 25(10):92, 94, 96–98, October 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_10/
                                        cport.txt.

Naccarato:2002:XQX

[Nac02] Giuseppe Naccarato. XSLT querying and XML documents. Dr. Dobb’s Journal of Software Tools, 27(12):24, 26, 28, December 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7718/
                                        ddj0212c/.

Nagle:2001:MFV

[Nag01]  Dan Nagle.  Multithread-

ing, Fthreads, and Visual Fortran. Dr. Dobb’s Journal of Software Tools, 26(7): 36, 38, 40, July 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_07/
                                        fthreads.zip.

Nagle:2002:FS

[Nag02] Dan Nagle. The Fortran 2000 Standard. Dr. Dobb’s Journal of Software Tools, 27(2): 48, 50–52, February 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_02/
                                        fortn2k.txt.

Nakhimovsky:2001:ISM

[Nak01] Greg Nakhimovsky. Improving scalability of multithreaded dynamic memory allocation. Dr. Dobb’s Journal of Software Tools, 26(7):44, 46, 48– 50, 52, 54, July 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_07/
mthot.txt;   http://www.
ddj.com/ftp/2001/2001_07/
mthot.zip.

Nelson:2002:TTS

[Nel02a]  Christopher Nelson.  Tcl/

Tk and SKILL mix it up. Dr. Dobb’s Journal of Software Tools, 27(2):42, 44– 46, February 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_02/
tclskill.txt;   http://www.
ddj.com/ftp/2002/2002_02/
tclskill.zip.

Nelson:2002:AA

[Nel02b] Mark R. Nelson. Algorithm alley. Dr. Dobb’s Journal of Software Tools, 27 (8):??, August 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_08/
aa0802.txt;   http://www.
ddj.com/ftp/2002/2002_08/
aa0802.zip.

Nelson:2002:IIJ

[Nel02c] Mark R. Nelson. Inside Intel’s JPEG library. Dr. Dobb’s Journal of Software Tools, 27 (7):38, 40, 42, 44, July 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2002/2002_07/
                                        slide.zip.

Nelson:2002:IMJ

[Nel02d] Mark R. Nelson. An intelligent MP3 jukebox. Dr. Dobb’s Journal of Software Tools, 27(10):24, 26, 28, 32, 34, October 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_10/
                                        llmp3.txt;   http://www.
                                        ddj.com/ftp/2002/2002_10/
                                        llmp3.zip.

Nerurkar:2000:SAD

[Ner00] Uttara Nerurkar. Security analysis and design. Dr. Dobb’s Journal of Software Tools, 25(11):50, 52, 54, 56, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Nilsson:2000:FSA

[Nil00] Stefan Nilsson. The fastest sorting algorithm? Dr. Dobb’s Journal of Software Tools, 25(4):38, 40, 42, 44– 45, April 2000. CODEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_04/
                                        fastsort.txt;   http://www.
                                        ddj.com/ftp/2000/2000_04/
                                        fastsort.zip.

Nisley:2001:ESD

[Nis01a] Ed Nisley. Embedded space: Dynamic attention. Dr. Dobb’s Journal of Software Tools, 26 (12):107–109, December 2001. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2001:ESE

[Nis01b] Ed Nisley. Embedded space:

Embedded elections. Dr. Dobb’s Journal of Software Tools, 26(4):13500137, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2001:ESLa

[Nis01c] Ed Nisley. Embedded space:

Let’s talk about specs. Dr. Dobb’s Journal of Software Tools, 26(5):129–131, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2001:ESLb

[Nis01d] Ed Nisley. Embedded space:

Life support. Dr. Dobb’s Journal of Software Tools, 26(11): 107–109, November 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Nisley:2001:ESP

[Nis01e] Ed Nisley. Embedded space:

Penguin specs. Dr. Dobb’s Journal of Software Tools, 26 (6):139–141, June 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Nisley:2001:ESRc

[Nis01f] Ed Nisley. Embedded space:

Rating real time: Count the ways. Dr. Dobb’s Journal of Software Tools, 26(9):113– 115, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2001:ESRb

[Nis01g] Ed Nisley. Embedded space:

Rating real time: Design points. Dr. Dobb’s Journal of Software Tools, 26(8):111– 112, August 2001. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2001:ESRa

[Nis01h] Ed Nisley. Embedded space: Rating real time: Hard and brittle. Dr. Dobb’s Journal of Software Tools, 26(7):121–123, July 2001. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2001:ESRd

[Nis01i] Ed Nisley. Embedded space: Real time zones. Dr. Dobb’s Journal of Software Tools, 26 (10):111–112, 115, October 2001. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2001:ESS

[Nis01j] Ed Nisley. Embedded space: SEU meets Embedded Linux. Dr. Dobb’s Journal of Software Tools, 26(3):129–131, March 2001. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2002:ESae

[Nis02a] Ed Nisley. Embedded space. Dr. Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2002:ESA

[Nis02b] Ed Nisley. Embedded space: All your drop are belong to us. Dr. Dobb’s Journal of Software Tools, 27(9):82, 85– 86, September 2002. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2002:ESCb

[Nis02c] Ed Nisley. Embedded space:

Chemical attraction. Dr. Dobb’s Journal of Software Tools, 27(7):83–85, July 2002. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2002:ESE

[Nis02d] Ed Nisley. Embedded space:

ESC scenes. Dr. Dobb’s Journal of Software Tools, 27(1): 87–89, January 2002. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2002:ESb

[Nis02e] Ed Nisley. Embedded space:

Java: Stirring the cup. Dr. Dobb’s Journal of Software Tools, 27(2):90, 92–93, February 2002. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2002:ESJ

[Nis02f] Ed Nisley. Embedded space:

Java: The mix. Dr. Dobb’s Journal of Software Tools, 27 (3):83–84, 86, March 2002. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2002:OSE

[Nis02g] Ed Nisley. Embedded space:

Only stone endures. Dr. Dobb’s Journal of Software Tools, 27(11):77–79, November 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7644/ddj0211i/.

Nisley:2002:ESP

[Nis02h] Ed Nisley. Embedded space:

Points of sale. Dr. Dobb’s Journal of Software Tools, 27 (5):77–78, 80, May 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Nisley:2002:ES

[Nis02i] Ed Nisley. Embedded space: Simulator sickness. Dr. Dobb’s Journal of Software Tools, 27 (10):76, 78–79, October 2002. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2002:ESS

[Nis02j] Ed Nisley. Embedded space: State of the machine. Dr. Dobb’s Journal of Software Tools, 27(12):73–74, 76, December 2002. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7718/
ddj0212i/.

Nisley:2002:EST

[Nis02k] Ed Nisley. Embedded space: Takin’ care of business. Dr. Dobb’s Journal of Software Tools, 27(6):87–89, June 2002. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2002:ESW

[Nis02l] Ed Nisley. Embedded space: Where you gonna get ’em? Dr. Dobb’s Journal of Software Tools, 27(4):93–95, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2003:ES

[Nis03a] Ed Nisley. Embedded space. Dr. Dobb’s Journal of Software Tools, 28(7):71–73, July 2003. CODEN DDJOEB. ISSN 1044-789X.

Nisley:2003:ESB

[Nis03b] Ed Nisley. Embedded space:

Boomin’ systems. Dr. Dobb’s Journal of Software Tools, 28 (1):67–69, January 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7758/
ddj0301i/.

Nisley:2003:ESRb

[Nis03c] Ed Nisley. Embedded space:

Real-Time Linux. Dr. Dobb’s Journal of Software Tools, 28 (4):79–81, April 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7827/
ddj0304j/.

Nisley:2003:ESRa

[Nis03d] Ed Nisley. Embedded space:

Rising tides. Dr. Dobb’s Journal of Software Tools, 28(3): 65–67, March 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7826/
ddj0303i/.

Nisley:2003:ESS

[Nis03e] Ed Nisley. Embedded space:

Spectral analysis. Dr. Dobb’s Journal of Software Tools, 28 (2):71–73, February 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7790/
ddj0302j/.

Nisley:2003:EST

[Nis03f] Ed Nisley. Embedded space: Testing one two. Dr. Dobb’s Journal of Software Tools, 28 (5):80–83, May 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7927/
                                        ddj0305j/.

Nisley:2003:ESW

[Nis03g] Ed Nisley. Embedded space: Written in blood. Dr. Dobb’s Journal of Software Tools, 28 (6):72, 74–75, June 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=8213/
                                        ddj0306i/.

Nieh:2000:EV

[NL00] Jason Nieh and Ozgur Can Leonard. Examining VMware. Dr. Dobb’s Journal of Software Tools, 25(8):70, 72–74, 76, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Ni:2000:WVC

[NN00] Ping Ni and Mark Nelson. Win32 version control. Dr. Dobb’s Journal of Software Tools, 25(6):62, 64–66, June 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_06/
                                        win32ver.txt.

OByrne:2003:SMU

[O’B03] Brian O’Byrne. State machines and user interfaces. Dr. Dobb’s Journal of Software Tools, 28(1):34, 36, 38, January 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7758/ddj0301d/.

Obasanjo:2003:XRD

[Oba03] Dare Obasanjo. XML and relational databases. Dr. Dobb’s Journal of Software Tools, 28 (5):54, 56–58, May 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7927/
ddj0305e/.

Okmianski:2002:ED

[Okm02] Anton Okmianski. Embedded databases. Dr. Dobb’s Journal of Software Tools, 27 (12):46, 48, 50, 52–53, December 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7718/ddj0212f/.

OMalley:2001:AAO

[O’M01] Kevin O’Malley. Agents and automated online trading. Dr. Dobb’s Journal of Software Tools, 26(5):23–24, 26–28, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Poliashenko:2002:GWA

[PA02] Max Poliashenko and Chip Andrews. Globalized Web applications and ASP.NET. Dr. Dobb’s Journal of Software Tools, 27(4):67–70, 72, April 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ftp/2002/2002_
04/aspnet.txt.

Pallman:2001:NNQ

[Pal01a] David Pallman. NQL: The Network Query Language. Dr. Dobb’s Journal of Software Tools, 26(1):158, 160, 162, 164, January 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_01/nql.txt.

Palshikar:2001:AAS

[Pal01b] Girish Keshav Palshikar. Algorithm alley: Simulated annealing: A heuristic optimization algorithm. Dr. Dobb’s Journal of Software Tools, 26(9): 121–124, September 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_09/
                                        aa0901.txt;   http://www.
                                        ddj.com/ftp/2001/2001_09/
                                        aa0901.zip.

Palshikar:2001:CBD

[Pal01c] Girish Keshav Palshikar. Consistencybased diagnosis. Dr. Dobb’s Journal of Software Tools, 26 (3):50, 52, 54–56, March 2001. CODEN DDJOEB. ISSN 1044-789X. See correction [Ano01].

Papaefstathiou:2001:PCB

[Pap01] Stathis Papaefstathiou. Predicting communications bottlenecks. Dr. Dobb’s Journal of Software Tools, 26(4):58, 60, 62–64, 66, April 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_04/

ccmod.txt;   http://www.
ddj.com/ftp/2001/2001_04/
ccmod.zip.

Papaefstathiou:2002:PBP

[Pap02] Stathis Papaefstathiou. Programmer’s bookshelf: Performance modeling for the masses. Dr. Dobb’s Journal of Software Tools, 27(4): 108, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Pashea:2000:MCM

[Pas00a] Michael Pashea. Microcontrollers and CORDIC methods. Dr. Dobb’s Journal of Software Tools, 25(9):86, 88– 90, 92, September 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_09/
cordic.txt;   http://www.
ddj.com/ftp/2000/2000_09/
cordic.zip.

Passani:2000:CWS

[Pas00b] Luca Passani. Creating WAP services. Dr. Dobb’s Journal of Software Tools, 25(7):70, 73– 75, 78, July 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_07/wapserv.
txt;   http://www.ddj.com/
ftp/2000/2000_07/wapserv.
zip.

Paton:2001:EER

[Pat01] Keith Paton. Extraction and examination of relations in C++: Principles of good physical design, courtesy of Lakos and Martin. Dr. Dobb’s Journal of Software Tools, 26(10): 28, 30, 32, 34, October 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_10/
extract.txt.

Perez:2000:MDM

[PB00] César A. Gonzalez Perez and Pablo Criado Boado. The Metis design methodology. Dr. Dobb’s Journal of Software Tools, 25(6):21–22, 24–26, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Patten:2000:LLE

[PEYS00] Michael Patten, Jonathan Erickson, Eliezer S. Yudkowsky, and Larry Sollman. Letters: Larry Ellison’s jet; the true name of the singularity; the sixth sense. Dr. Dobb’s Journal of Software Tools, 25(3): 12, 14, March 2000. CODEN DDJOEB. ISSN 1044-789X.

Pollard:2000:LNF

[PGH^{+}00] W. Thomas Pollard, Jan Theodore Galkowski, John Herbster, Scott Maxwell, and Al Stevens. Letters: Nanoseconds not fast enough? why not Smalltalk? shared responsibility for virus attacks; Mingw32. Dr. Dobb’s Journal of Software Tools, 25 (8):10, 12, August 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Perelman-Hall:2000:JQ

[PH00] David Perelman-Hall. Java Q&A: How do I build a find? Dr. Dobb’s Journal of Software Tools, 25(7):119–120, 122, 124, 126, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
jqa700.txt;   http://www.
ddj.com/ftp/2000/2000_07/
jqa700.zip.

Phillips:2001:VVW

[Phi01] Lee Anne Phillips. VoiceXML and the voice/Web environment. Dr. Dobb’s Journal of Software Tools, 26(10):91–92, 94, 96, October 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_10/
voicexml.zip;   http://www.
ddj.com/ftp/2001/2001_10/
voicexml_txt.zip.

Pierce:2001:BSB

[Pie01] Bill Pierce. Building servicebased architectures with Jini. Dr. Dobb’s Journal of Software Tools, 26(6):50, 52– 58, June 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_06/sba.txt.

Pigeon:2002:AAQ

[Pig02] Steven Pigeon. Algorithm alley: Quicksort and radix sorts on lists. Dr. Dobb’s Journal of Software Tools, 27(5):89– 90, 92, 94, May 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_05/
aa0502.txt;   http://www.

                                        ddj.com/ftp/2002/2002_05/
                                        aa0502.zip.

Pohjonen:2002:DSM

[PK02] Risto Pohjonen and Steven Kelly. Domain-specific modeling. Dr. Dobb’s Journal of Software Tools, 27(8): ??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_08/dsm.txt;
                                         http://www.ddj.com/ftp/
                                        2002/2002_08/dsm.zip.

Pletzer:2002:PFE

[Ple02a] Alexander Pletzer. Python and finite elements. Dr. Dobb’s Journal of Software Tools, 27 (3):36, 38–40, March 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2002/2002_03/
                                        ellipt2d.txt;   http://www.
                                        ddj.com/ftp/2002/2002_03/

ellipt2d.zip. See correction [Ano02a].

Pleumann:2002:MP

[Ple02b] Jörg Pleumann. The Mail4Me Project. Dr. Dobb’s Journal of Software Tools, 27(6):36, 38, 40, 42, 44, June 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_06/mail4me.
                                        txt.

Pont:2003:CCL

[Pon03] Mick Pont. Calling C library routines from Java. Dr. Dobb’s Journal of Software Tools, 28(7):28, 32, 34– 36, 38, July 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2003/2003_07/
c_java.txt;   http://www.
ddj.com/ftp/2003/2003_07/
c_java.zip.

Porter:2003:MDX

[Por03] Brett Porter. Managing devices with XML-RPC. Dr. Dobb’s Journal of Software Tools, 28(4):66, 68–70, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7827/ddj0304h/.

Pote:2001:IS

[Pot01] John Pote. Interrupt scheduling. Dr. Dobb’s Journal of Software Tools, 26(2):104, 106, February 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_02/
intsched.txt.

Puttick:2002:LDP

[PTB^{+}02] James H. Puttick, Andrew D.

Todd, Joe Block, Bob Mc-Connell, and The Editors. Letters: Determining processor utilization; JPEG prior art; little languages; DDJ archives. Dr. Dobb’s Journal of Software Tools, 27(10):10, October 2002. CODEN DDJOEB. ISSN 1044-789X.

Pugh:2002:ATW

[Pug02] Tom Pugh. Automated testing and Windows CE. Dr. Dobb’s Journal of Software Tools, 27 (12):54, 56, 58, 60, 62, December 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7718/ddj0212g/.

Pumar:2003:DP

[Pum03] Mark A. Pumar. Dynamic programming. Dr. Dobb’s Journal of Software Tools, 28(5):18, 20, 22, 24, 26, May 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7927/
ddj0305b/.

Pendergrass:2002:LCB

[PWS^{+}02] Dan Pendergrass, Jim Williams, Seth Schoen, Cesar A. Gonzalez Perez, and David Charlap. Letters: A call for B++; Visual Basic and VB.NET; security inside and out; the D language; heat pumps. Dr. Dobb’s Journal of Software Tools, 27 (5):10, May 2002. CODEN DDJOEB. ISSN 1044-789X.

Qureshi:2001:AAO

[Qur01] Shehrzad Qureshi. Algorithm alley: Optimization and fixedpoint iteration. Dr. Dobb’s Journal of Software Tools, 26 (12):115–117, December 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_12/
aa1201.txt;   http://www.
ddj.com/ftp/2001/2001_12/
aa1201.zip.

Qureshi:2003:CCM

[Qur03]  Shehrzad Qureshi.  Cross-

correlation and matched filters. Dr. Dobb’s Journal of Software Tools, 28(4):52, 54–56, April 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7827/
ddj0304f/.

Rahavachary:2001:LSS

[Rah01] Saty Rahavachary. Letters:

Setting the sqrt() record straight. Dr. Dobb’s Journal of Software Tools, 26(4):12, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Raisanen:2001:QSV

[Räi01] Vilho Räisänen. Quality of service and voice-over-IP. Dr. Dobb’s Journal of Software Tools, 26(5):50, 54, 56, 58, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Ramaswamy:2001:RSL

[Ram01] Ramkumar Ramaswamy. A request scheduling layer for Apache. Dr. Dobb’s Journal of Software Tools, 26(12): 83–86, December 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_12/
apache.txt.

Rao:2002:JQ

[Rao02] Surlu Rao. Java Q&A: How do message-driver beans work in EJB 2.0? Dr. Dobb’s Journal of Software Tools, 27(2):94– 96, February 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_02/jqa0202.
txt;   http://www.ddj.com/
ftp/2002/2002_02/jqa0202.
zip.

Ratcliff:2002:CPL

[Rat02] John Ratcliff. The Cg programming language. Dr. Dobb’s Journal of Software Tools, 27(10):40, 42, 44, October 2002. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_10/
cgbind.txt;   http://www.
ddj.com/ftp/2002/2002_10/
cgbind.zip.

Roberts:2000:LPM

[RCE^{+}00] Tim Roberts, Ben Combee, Eddie Edwards, Keith McBride, Matt Albrecht, Paul Kinzelman, Ken Anderson, and Arch D. Robison. Letters: Patent madness; C++ identifiers; digital filtering and oversampling; pay phones versus cell phones; NASA and the Space Shuttle; getting the lead out…; how harmful is recursion. Dr. Dobb’s Journal of Software Tools, 25(7):10, 12, July 2000. CODEN DDJOEB. ISSN 1044-789X.

Reed:2000:PMC

[Ree00] Mark Reed. The promise of molecular computing. Dr. Dobb’s Journal of Software Tools, 25(12 (supplement)):13– 15, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Reilly:2000:JQH

[Rei00a] David Reilly. Java Q&A:

How do I user servlets for state and session management? Dr. Dobb’s Journal of Software Tools, 25(5): 111–115, May 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_05/
jqa500.txt;   http://www.
ddj.com/ftp/2000/2000_05/
jqa500.zip.

Reilly:2000:PBD

[Rei00b] Doug Reilly. Programmer’s bookshelf: Data is what it is all about. Dr. Dobb’s Journal of Software Tools, 25(9):145– 146, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Reilly:2001:PBM

[Rei01a] Doug Reilly. Programmer’s bookshelf: Microsoft SQL server cubed. Dr. Dobb’s Journal of Software Tools, 26(6): 159–160, June 2001. CODEN DDJOEB. ISSN 1044-789X.

Reilly:2001:TNF

[Rei01b]  Douglas Reilly.  Thread-

ing and the .Net framework. Dr. Dobb’s Journal of Software Tools, 26(8):30, 32–33, 36, 38, August 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_08/
thrednet.txt.

Reilly:2002:ANS

[Rei02] Douglas Reilly. ASP.NET server components. Dr. Dobb’s Journal of Software Tools, 27 (6):62, 64, 66–68, June 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2002/2002_06/
                                        aspnet.txt;   http://www.
                                        ddj.com/ftp/2002/2002_06/
                                        aspnet.zip.

Rempt:2001:PPT

[Rem01a] Boudewijn Rempt. Python’s PyQt toolkit. Dr. Dobb’s Journal of Software Tools, 26(1):88, 90, 92, 94, January 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_01/
                                        pyqt.txt.

Rempt:2001:SJP

[Rem01b] Boudewijn Rempt. Scripting with Java and Python: Building a Python console window in a Java application. Dr. Dobb’s Journal of Software Tools, 26(10):56, 60–61, October 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_10/
                                        pyconsol.txt;   http://www.
                                        ddj.com/ftp/2001/2001_10/
                                        pyconsol.zip.

Rousselle:2000:PSJ

[RG00] Philip Rousselle and Daniel Greff. Publish, subscribe, and the JMS API. Dr. Dobb’s Journal of Software Tools, 25 (7):44, 46, 48, 51, July 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_07/
pubsub.zip.

Richter:2000:REN

[Ric00] H. Thomas Richter. Reliability and embedded networks. Dr. Dobb’s Journal of Software Tools, 25(8):86, 88, 90, 92, 94, 96, August 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_08/embednw.
zip.

Riemersma:2002:ISB

[Rie02] Thiadmer Riemersma. Image scaling with Bresenham. Dr. Dobb’s Journal of Software Tools, 27(5):21– 22, 24, 26, May 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_05/
image.txt;   http://www.
ddj.com/ftp/2002/2002_05/
image.zip.

Ritke:2000:IT

[Rit00] Ronn Ritke. An Internet toolbox. Dr. Dobb’s Journal of Software Tools, 25(6):84, 86– 87, 89, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Robison:2000:CR

[Rob00] Arch D. Robison. Considering recursion. Dr. Dobb’s Journal of Software Tools, 25(3): 52, 54–55, 57–58, March 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2000/2000_03/
                                        recurs.txt.

Robinson:2001:DSL

[Rob01a] Rich Robinson. DigitaScript: A scripting language for digital cameras. Dr. Dobb’s Journal of Software Tools, 26(1):126, 128– 130, 132, January 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_01/
                                        dscript.txt.

Robison:2001:MB

[Rob01b] Arch D. Robison. Moving up to 64 bits. Dr. Dobb’s Journal of Software Tools, 26(2):66, 68, 70, 72–73, February 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_02/
                                        64bit.txt.

Robison:2003:MCN

[Rob03] Arch D. Robison. Memory consistency and .NET. Dr. Dobb’s Journal of Software Tools, 28 (4):46, 48–50, April 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7827/
                                        ddj0304e/.

Roelofs:2000:JCC

[Roe00] Brian Roelofs. Java custom class loaders. Dr. Dobb’s Journal of Software Tools, 25(6):74, 76, 78, 80, 82, June 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_06/

javaload.txt;   http://www.
ddj.com/ftp/2000/2000_06/
javaload.zip.

Rojas:2000:SKZ

[Roj00]  Raul Rojas.  Simulating

Konrad Zuse’s computers. Dr. Dobb’s Journal of Software Tools, 25(9):64, 66–69, September 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_09/z3.zip.

Rolfe:2000:AAA

[Rol00a] Timothy Rolfe. Algorithm alley: AVL trees. Dr. Dobb’s Journal of Software Tools, 25 (12):149–152, December 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_12/
aa0012.txt.

Rolfe:2000:AAR

[Rol00b] Timothy Rolfe. Algorithm alley: Randomized shuffling. Dr. Dobb’s Journal of Software Tools, 25(1):113– 114, January 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_01/
aa120.txt;   http://www.
ddj.com/ftp/2000/2000_01/
aa120.zip.

Rolfe:2002:AAG

[Rol02] Timothy Rolfe. Algorithm alley: Graph traversals. Dr. Dobb’s Journal of Software Tools, 27(3):97–98, 100– 101, March 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_03/
                                        aa0302.txt;   http://www.
                                        ddj.com/ftp/2002/2002_03/
                                        aa0302.zip.

Rolfe:2003:BBP

[Rol03]  Timothy Rolfe.  Bargain-

basement parallelism. Dr. Dobb’s Journal of Software Tools, 28(2):46, 48, 50, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7790/ddj0302f/.

Roth:2000:PBT

[Rot00] Peter N. Roth. Programmer’s bookshelf: Two-year olds. Dr. Dobb’s Journal of Software Tools, 25(10):159–160, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Roth:2001:PBC

[Rot01a] Peter Roth. Programmer’s bookshelf: A C++ duo. Dr. Dobb’s Journal of Software Tools, 26(8):124–125, August 2001. CODEN DDJOEB. ISSN 1044-789X.

Roth:2001:PBA

[Rot01b] Peter N. Roth. Programmer’s bookshelf: Accelerating C++ programming. Dr. Dobb’s Journal of Software Tools, 26 (3):144, March 2001. CODEN DDJOEB. ISSN 1044-789X.

Roth:2002:PBM

[Rot02a] Peter N. Roth. Programmer’s bookshelf: Mathematics and C++. Dr. Dobb’s Journal of Software Tools, 27(7):100, July 2002. CODEN DDJOEB. ISSN 1044-789X.

Roth:2002:JSA

[Rot02b]  Volker Roth.  Java secu-

rity architecture and extensions. Dr. Dobb’s Journal of Software Tools, 27(4):34, 36–38, April 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_04/jca.txt;
http://www.ddj.com/ftp/
2002/2002_04/jca.zip.

Rousselle:2002:IJP

[Rou02] Philip Rousselle. Implementing the JMS publish/subscribe API. Dr. Dobb’s Journal of Software Tools, 27(4):28, 30– 32, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Rayboud:2001:LSL

[RSW^{+}01] Andrew Rayboud, Larry Smith, Steve Walton, Jared Jackson, Joseph Link, Ralph Moore, Ngaire Atmore, and The Editors. Letters: Shiftleft testing; curley [sic] brackets; trampolines for embedded systems; we’re New Zealanders dammit; correction: “C++ Set-Theoretic Operations On Virtual Containers”. Dr. Dobb’s Journal of Software Tools, 26(12):10, 12, December 2001. CODEN DDJOEB. ISSN 1044-789X. See [BFG01].

Rubin:2000:KVL

[Rub00] Aviel D. Rubin. Kerberos versus the Leighton-Micali protocol. Dr. Dobb’s Journal of Software Tools, 25(11):21–22, 24, 26, November 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Ryland:2002:LLL

[RWD02] Chris Ryland, Dave Waller, and Paul M. Dubuc. Letters: The lightweight language workshop; strike one. … Dr. Dobb’s Journal of Software Tools, 27(4–4):10, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Ross:2000:DFP

[RZ00] Darryl Ross and Con Zymaris. DB Forms: PHP, MySQL, and PHPLIB. Dr. Dobb’s Journal of Software Tools, 25(8):98, 100–104, August 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_08/dbforms.
txt;   http://www.ddj.com/
ftp/2000/2000_08/dbforms.
zip.

Soref:2000:ESD

[SA00] Jeremy Soref and Sheridan Adjei. Electronic software distribution. Dr. Dobb’s Journal of Software Tools, 25(9):100, 102– 106, September 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_09/license.
txt.

Sams:2003:SDS

[Sam03] Mark Sams. The SquareList data structure. Dr. Dobb’s Journal of Software Tools, 28 (5):37–40, May 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7927/
ddj0305c/.

Sassenrath:2000:RSL

[Sas00] Carl Sassenrath. The REBOL scripting language. Dr. Dobb’s Journal of Software Tools, 25 (7):64, 66–69, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
rebol.txt.

Savchenko:2000:TFS

[Sav00] Sergei Savchenko. Tools for flexible scripting. Dr. Dobb’s Journal of Software Tools, 25 (1):52, 54–55, January 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_01/
flexscrp.txt;   http://www.
ddj.com/ftp/2000/2000_01/
flexscrp.zip.

Savchenko:2002:SM

[Sav02] Sergei Savchenko. Shadow mapping. Dr. Dobb’s Journal of Software Tools, 27(7): 32, 34, 36–37, July 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_07/
sdsl.zip.

Shasha:2002:STY

[SB02] Dennis E. Shasha and Philippe Bonnet. Smooth talking your databases. Dr. Dobb’s Journal of Software Tools, 27(7):46. 48, 50–52, 54, July 2002. CODEN DDJOEB. ISSN 1044-789X.

Schneier:2000:SRF

[Sch00] Bruce Schneier. Security research and the future. Dr. Dobb’s Journal of Software Tools, 25(12 (supplement)):33– 35, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Scott:2000:IUI

[Sco00] Neil Scott. Innovations in user interfaces. Dr. Dobb’s Journal of Software Tools, 25(12 (supplement)):23–27, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Scott:2002:AAO

[Sco02] David A. Scott. Algorithm alley: Optimal EOF handling with arithmetic compression. Dr. Dobb’s Journal of Software Tools, 27(1): 99, 102, January 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_01/
                                        aa0102.zip.

Seaman:2002:JQH

[Sea02] Mark Seaman. Java Q&A: How can i generate Java code for tables in my database? Dr. Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2002/2002_08/
jqa0802.txt;   http://www.
ddj.com/ftp/2002/2002_08/
jqa0802.zip.

SantaCruz:2001:JIC

[SEC01] Diego Santa Cruz, Touradj Ebrahimi, and Charilaos Christopoulos. The JPEG 2000 image coding standard. Dr. Dobb’s Journal of Software Tools, 26(4):46, 48, 50–52, 54, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://

jj2000.epfl.ch/;   http:/
/www.ddj.com/; http://www.
ece.ubc.ca/~mdadams/jasper/
;   http://www.imagepower.
com/products/ubcform.htm;
http://www.jpeg.org/;   http:
//www.jpeg.org/public/wavedemo.
zip.

Stallings:2001:PBB

[SG01] William Stallings and Lou Grinzo. Programmer’s bookshelf: A Bluetooth trio and C#. Dr. Dobb’s Journal of Software Tools, 26(9):130, 132, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOj

[Sha00a] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Beasts. Dr. Dobb’s Journal of Software Tools, 25(10):155–157, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOd

[Sha00b] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Blood. Dr. Dobb’s Journal of Software Tools, 25(4):119–120, 122, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOl

[Sha00c] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Causality. Dr. Dobb’s Journal of Software Tools, 25(12):155–156, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOe

[Sha00d] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Centimillionaire. Dr. Dobb’s Journal of Software Tools, 25(5):123–124, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOi

[Sha00e] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Chimps. Dr. Dobb’s Journal of Software Tools, 25(9):142–143, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOh

[Sha00f] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Escape. Dr. Dobb’s Journal of Software Tools, 25(8):141–142, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOf

[Sha00g]  Dennis E. Shasha.  Dr.

Ecco’s omniheurist corner: Microvirus. Dr. Dobb’s Journal of Software Tools, 25(6):139, 141, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOk

[Sha00h] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Mint. Dr. Dobb’s Journal of Software Tools, 25(11):156–157, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOc

[Sha00i] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Simple. Dr. Dobb’s Journal of Software Tools, 25(3):128, March 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOa

[Sha00j] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Stars and starlets. Dr. Dobb’s Journal of Software Tools, 25(1):116, January 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOb

[Sha00k] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Sticks. Dr. Dobb’s Journal of Software Tools, 25(2):131–132, February 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2000:DEOg

[Sha00l] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Wordsnakes. Dr. Dobb’s Journal of Software Tools, 25(7):143–144, July 2000. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOc

[Sha01a] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Ambulance. Dr. Dobb’s Journal of Software Tools, 26(3):141– 142, March 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOg

[Sha01b] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Beams. Dr. Dobb’s Journal of Software Tools, 26(7):135–136, July 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOe

[Sha01c] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Beats. Dr. Dobb’s Journal of Software Tools, 26(5):148–149, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOj

[Sha01d] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Child’s ply. Dr. Dobb’s Journal of Software Tools, 26(10):126, 128, October 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOd

[Sha01e] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Foxy. Dr. Dobb’s Journal of Software Tools, 26(4):148–149, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOf

[Sha01f] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Panamax. Dr. Dobb’s Journal of Software Tools, 26(6):156–157, June 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOh

[Sha01g] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Perimeters. Dr. Dobb’s Journal of Software Tools, 26(8):121, August 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOk

[Sha01h] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Sprawl. Dr. Dobb’s Journal of Software Tools, 26(11):125–126, November 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOb

[Sha01i] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Tundra. Dr. Dobb’s Journal of Software Tools, 26(2):156–157, February 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOl

[Sha01j] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Ultimate tic-tac-toe. Dr. Dobb’s Journal of Software Tools, 26(12):119– 120, December 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOi

[Sha01k] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Who rules? Dr. Dobb’s Journal of Software Tools, 26(9):126– 129, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2001:DEOa

[Sha01l] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Wildfires. Dr. Dobb’s Journal of Software Tools, 26(1):193–194, January 2001. CODEN DDJOEB. ISSN 1044-789X.

Shammas:2002:AAE

[Sha02a] Namir Clement Shammas. Algorithm alley: Enhancing Newton’s method. Dr. Dobb’s Journal of Software Tools, 27 (6):94–97, June 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2002/2002_06/
                                        aa0602.zip.

Shasha:2002:DEOg

[Sha02b] Dennis E. Shasha. Dr. Ecco’s omniheurist corner. Dr. Dobb’s Journal of Software Tools, 27 (7):98–99, July 2002. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2002:DEOh

[Sha02c] Dennis E. Shasha. Dr. Ecco’s omniheurist corner. Dr. Dobb’s Journal of Software Tools, 27 (8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2002:DEOi

[Sha02d] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Adversarial bifurcations. Dr. Dobb’s Journal of Software Tools, 27 (9):97–99, September 2002. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2002:DEOd

[Sha02e] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Color war. Dr. Dobb’s Journal of Software Tools, 27(4):106–107, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2002:DEOa

[Sha02f] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Desert sprinklers. Dr. Dobb’s Journal of Software Tools, 27(1):103– 106, January 2002. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2002:DEOb

[Sha02g] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: Duelists. Dr. Dobb’s Journal of Software Tools, 27(2):101, 103, February 2002. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2002:DEO

[Sha02h] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: The DNA on Noah’s Arc. Dr. Dobb’s Journal of Software Tools, 27 (10):89–90, October 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Shasha:2002:DEOf

[Sha02i] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: The protocol of small numbers2. Dr. Dobb’s Journal of Software Tools, 27(6):98, June 2002. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2002:DEOe

[Sha02j] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: The stone tombs of Zimbabwe. Dr. Dobb’s Journal of Software Tools, 27(5):98–99, May 2002. CODEN DDJOEB. ISSN 1044-789X.

Shasha:2002:DEOc

[Sha02k] Dennis E. Shasha. Dr. Ecco’s omniheurist corner: With eye of newt. Dr. Dobb’s Journal of Software Tools, 27(3):103– 104, March 2002. CODEN DDJOEB. ISSN 1044-789X.

Sharaafath:2003:CCW

[Sha03] Yassar Sharaafath. C++WSP: A C++ web services platform. Dr. Dobb’s Journal of Software Tools, 28(4):18, 20, 22, 24, 26, April 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7827/
ddj0304c/.

Sheong:2001:BDF

[She01] Chang Sau Sheong. Building dynamic fail-over Java servers. Dr. Dobb’s Journal of Software Tools, 26(11):46, 48, 50, 52– 54, 57, November 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_11/
xander.txt;   http://www.
ddj.com/ftp/2001/2001_11/
xander.zip.

Sherer:2003:RTS

[She03] Matt Sherer. Real-time signal analysis and Real-Time Linux: Part 1. Dr. Dobb’s Journal of Software Tools, 28 (7):62–65, July 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2003/2003_07/
rtlp1.txt.

Silverman:2003:SWN

[Sil03] Micah Silverman. Securing wireless networks. Dr. Dobb’s Journal of Software Tools, 28 (6):36, 38–40, 42, 44, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
8213/ddj0306d/.

Simcox:2000:WYW

[Sim00a] Susan Simcox. What are you worth? Dr. Dobb’s Journal of Software Tools, 25(4):??, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Simpson:2000:CVC

[Sim00b] Michael Simpson. CVS version control and branch management. Dr. Dobb’s Journal of Software Tools, 25(10):108, 110–114, October 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Sintes:2000:XSC

[Sin00] Tony Sintes. XML and software configuration. Dr. Dobb’s Journal of Software Tools, 25 (7):56, 58–62, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
xmlcfg.txt;   http://www.
ddj.com/ftp/2000/2000_07/
xmlcfg.zip.

Sipe:2000:VLW

[Sip00] Steve Sipe. Virtual lists for Win32. Dr. Dobb’s Journal of Software Tools, 25(12):56, 58, 61, December 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_12/
vlist.zip.

Siek:2001:CCC

[SL01a] Jeremy Siek and Andrew Lumsdaine. C++ concept checking. Dr. Dobb’s Journal of Software Tools, 26(6): 64, 66–68, 70, June 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_06/
cppcheck.txt.

Smith:2001:AOS

[SL01b] Larry Smith and Cameron Laird. Android: Open-source scripting for testing and automation. Dr. Dobb’s Journal of Software Tools, 26(7): 99–102, July 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_07/android.
txt.

Siek:2000:GGC

[SLL00] Jeremy G. Siek, Lie-Quan Lee, and Andrew Lumsdaine. The Generic Graph Component Library. Dr. Dobb’s Journal of Software Tools, 25(9):29–30, 32–34, 36, 38, September 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_09/
ggcl.txt;   http://www.ddj.
com/ftp/2000/2000_09/ggcl21.
zip.

Stanley-Marbell:2000:IAD

[SM00] Phillip Stanley-Marbell. Inferno application development with Limbo. Dr. Dobb’s Journal of Software Tools, 25 (12):113–114, 116, 118, 120, December 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_12/inferno.
txt.

Smith:2001:SLT

[Smi01a] Larry Smith. Shift-left testing.

Dr. Dobb’s Journal of Software Tools, 26(9):56, 62, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Smith:2001:JQH

[Smi01b] Terry Smith. Java Q&A: How can I simplify Java data access objects. Dr. Dobb’s Journal of Software Tools, 26(5):135–136, 138–139, May 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_05/jqa0501.
txt;   http://www.ddj.com/
ftp/2001/2001_05/jqa0501.
zip.

Smith:2002:AAP

[Smi02] Peter Smith. Algorithm alley:

Prime numbers. Dr. Dobb’s Journal of Software Tools, 27 (7):93–95, July 2002. CODEN DDJOEB. ISSN 1044-789X.

Smith:2003:ADI

[Smi03a]  Kevin W. Smith.  Auto-

mated defect identification. Dr. Dobb’s Journal of Software Tools, 28(2):16, 18, 20, 22, February 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7790/
                                        ddj0302b/.

Smith:2003:DL

[Smi03b] Kevin W. Smith. Distributed loadbuilds. Dr. Dobb’s Journal of Software Tools, 28(7): 58, 60, July 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2003/2003_07/
                                        disbuild.txt.

Smyth:2000:ICY

[Smy00] Gavin Smyth. Infrared control of your PC. Dr. Dobb’s Journal of Software Tools, 25(5):66, 68, 70, 72, May 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_05/
                                        infrared.txt;   http://www.
                                        ddj.com/ftp/2000/2000_05/
                                        infrared.zip.

Snow:2001:IAL

[Sno01]  Ray Snow.  An informa-

tion assembly line in Perl. Dr. Dobb’s Journal of Software Tools, 26(6):72, 76, 78, 80, 82, June 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_06/
infoline.txt.

Snyder:2003:RCG

[Sny03] Martin Snyder. Robotic control and 3D GUIs. Dr. Dobb’s Journal of Software Tools, 28 (1):16–17, 20, 22, 24, January 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7758/ddj0301b/.

Song:2001:Ia

[Son01a]  Della Song.  Of interest.

Dr. Dobb’s Journal of Software Tools, 26(9):134, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Song:2001:Ib

[Son01b] Della Song. Of interest. Dr.

Dobb’s Journal of Software Tools, 26(10):134–135, October 2001. CODEN DDJOEB. ISSN 1044-789X.

Song:2001:Ic

[Son01c] Della Song. Of interest. Dr.

Dobb’s Journal of Software Tools, 26(11):134–135, November 2001. CODEN DDJOEB. ISSN 1044-789X.

Song:2001:Id

[Son01d] Della Song. Of interest. Dr.

Dobb’s Journal of Software Tools, 26(12):126–127, December 2001. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:Ia

[Son02a] Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 27(1):110–111, January 2002. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:Ib

[Son02b] Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 27(2):110–111, February 2002. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:Ic

[Son02c] Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 27(3):109–110, March 2002. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:Id

[Son02d] Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 27(4):110–111, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:Ie

[Son02e] Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 27(5):102–103, May 2002. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:If

[Son02f] Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 27(6):102–103, June 2002. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:Ig

[Son02g] Della Song. Of interest. Dr.

Dobb’s Journal of Software Tools, 27(7):102, July 2002. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:Ih

[Son02h] Della Song. Of interest. Dr.

Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:Ii

[Son02i]  Della Song.  Of interest.

Dr. Dobb’s Journal of Software Tools, 27(9):102, September 2002. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:Ij

[Son02j]  Della Song.  Of interest.

Dr. Dobb’s Journal of Software Tools, 27(10):81, October 2002. CODEN DDJOEB. ISSN 1044-789X.

Song:2002:Ik

[Son02k] Della Song. Of interest. Dr.

Dobb’s Journal of Software Tools, 27(11):81, November 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7644/ddj0211o/.

Song:2003:Ia

[Son03a] Della Song. Of interest. Dr.

Dobb’s Journal of Software Tools, 28(1):74–75, January 2003. CODEN DDJOEB.

ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7758/ddj0301o/.

Song:2003:Ib

[Son03b] Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 28(2):79, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7790/
ddj0302p/.

Song:2003:Ic

[Son03c] Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 28(3):71, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7826/
ddj0303o/.

Song:2003:Id

[Son03d] Della Song. Of interest. Dr. Dobb’s Journal of Software Tools, 28(4):84, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7827/
ddj0304p/.

Schotland:2000:EHC

[SP00] Tom Schotland and Peter Petersen. Exception handling in C without C++. Dr. Dobb’s Journal of Software Tools, 25 (11):102, 104, 106, 108, 110, 112, November 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_11/
except.txt;   http://www.

ddj.com/ftp/2000/2000_11/
except.zip.

Smith:2002:MN

[SPM02] Jason Smith, Nigel Perry, and Erik Meijer. Mondrian for .Net. Dr. Dobb’s Journal of Software Tools, 27(2):28, 30– 32, 34, February 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_02/
mondrian.txt;   http://www.
ddj.com/ftp/2002/2002_02/
mondrian.zip.

Souflis:2002:LCN

[SPRF02] Dimitrios Souflis, Glenn Porter, Bruce Reed, and Carol Flaherty. Letters: The CLR & .NET; real simulator sickness; TSR = oops!; MacroSafe clarification. Dr. Dobb’s Journal of Software Tools, 27(12): 10, December 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7718/
ddj0212l/.

Spurgat:2000:WWC

[Spu00] Jeff Spurgat. Writing Windows CE display drivers. Dr. Dobb’s Journal of Software Tools, 25(9):76, 78–80, 82, 84, September 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_09/wincedd.
txt;   http://www.ddj.com/
ftp/2000/2000_09/wincedd.
zip.

Samwel:2000:LDS

[SSC^{+}00] Bart Samwel, Jiri Soukup, Glenn Crist, Evan Easton, Ron Ruble, David A. Rogers, Al Stevens, Bruce MacDonald, and Scott Venckus. Letters: Data structures as objects; real (Netscape) time; riding the XML bandwagon; porting to CE; nothing new about Open Source; Y2K worries?; version control. Dr. Dobb’s Journal of Software Tools, 25 (2):12, 14, February 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Stallings:2000:GE

[Sta00a] William Stallings. Gigabit Ethernet. Dr. Dobb’s Journal of Software Tools, 25(5): 34, 36–37, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Stallings:2000:PB

[Sta00b] William Stallings. Programmer’s bookshelf: Frame relay wrap-up. Dr. Dobb’s Journal of Software Tools, 25(7):147, July 2000. CODEN DDJOEB. ISSN 1044-789X.

Stallings:2000:SSC

[Sta00c] William Stallings. The SET Standard and E-commerce. Dr. Dobb’s Journal of Software Tools, 25(11):30, 32, 34, 36, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Stallings:2001:AAO

[Sta01a] William Stallings. Algorithm alley: Orthogonal sequences.

Dr. Dobb’s Journal of Software Tools, 26(8):118–120, August 2001. CODEN DDJOEB. ISSN 1044-789X.

Stallman:2001:LSD

[Sta01b] Richard Stallman. Letters:

Setting the Debian record straight. Dr. Dobb’s Journal of Software Tools, 26(4):10, April 2001. CODEN DDJOEB. ISSN 1044-789X. Comments on the GNU HURD kernel, the GNU Project, and the confusion between the Open Source Movement and the Free Software Foundation, and between GNU and Linux.

Stevens:2000:CPB

[Ste00a]  Al Stevens.  C program-

ming: Building Quincy 2000. Dr. Dobb’s Journal of Software Tools, 25(11):133–134, 136, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2000:CPG

[Ste00b] Al Stevens. C programming:

Going undercover. Dr. Dobb’s Journal of Software Tools, 25 (7):113–117, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
cprog.txt.

Stevens:2000:CPW

[Ste00c] Al Stevens. C programming:

Into the world of MP3. Dr. Dobb’s Journal of Software Tools, 25(9):115–116, 118, 120, September 2000. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2000:CPI

[Ste00d]  Al Stevens.  C program-

ming: Iterating the days until Y2K. Dr. Dobb’s Journal of Software Tools, 25(3):107– 108, 110, March 2000. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2000:CPQ

[Ste00e] Al Stevens. C programming: Quincy 2000: Customizing the print dialogs. Dr. Dobb’s Journal of Software Tools, 25(12): 129–130, 132, 134, 136, 138– 139, December 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_12/
                                        cprog012.txt.

Stevens:2000:CPR

[Ste00f] Al Stevens. C programming: Retreading Quincy. Dr. Dobb’s Journal of Software Tools, 25 (6):106, 108, 110, 112, 114, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2000:CPSd

[Ste00g] Al Stevens. C programming: Shareware, adware, spyware. Dr. Dobb’s Journal of Software Tools, 25(10):123–124, 126, 128, 130, October 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:2000:CPSc

[Ste00h] Al Stevens. C programming: Slamming, compilers, E-books. Dr. Dobb’s Journal of Software Tools, 25(5):105–107, 109, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2000:CPSb

[Ste00i] Al Stevens. C programming:

Speaking of rambling. Dr. Dobb’s Journal of Software Tools, 25(4):97–99, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2000:CPSa

[Ste00j]  Al Stevens.  C program-

ming: Spinning purple larks, swirling leaves, rotating ellipses. Dr. Dobb’s Journal of Software Tools, 25(1): 95–98, January 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_01/
cprog120.zip.

Stevens:2000:CPP

[Ste00k]  Al Stevens.  C program-

ming: The S programming language. Dr. Dobb’s Journal of Software Tools, 25(2): 110–113, February 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_02/
cprog220.txt.  See com-

ment on another S language [KSC^{+}00].

Stevens:2000:CPT

[Ste00l] Al Stevens. C programming:

Twelve years and still complaining. Dr. Dobb’s Journal of Software Tools, 25(8):119– 121, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPMa

[Ste01a] Al Stevens. C programming: A moving target. Dr. Dobb’s Journal of Software Tools, 26 (4):129–130, 132–133, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPC

[Ste01b]  Al Stevens.  C program-

ming: Changing soapboxes midstream. Dr. Dobb’s Journal of Software Tools, 26(2): 139–140, 142–143, February 2001. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPD

[Ste01c] Al Stevens. C programming: Delivery woes and music minus whatever. Dr. Dobb’s Journal of Software Tools, 26(11): 103–105, November 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_11/
                                        cpro1101.txt;   http://www.
                                        ddj.com/ftp/2001/2001_11/
                                        cpro1101.zip.

Stevens:2001:CPA

[Ste01d] Al Stevens. C programming: I almost get a Linux editor and compiler. Dr. Dobb’s Journal of Software Tools, 26(7): 113–116, July 2001. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPG

[Ste01e] Al Stevens. C programming: It’s good work when you can find it. Dr. Dobb’s Journal of Software Tools, 26(5):121–124, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPMb

[Ste01f]  Al Stevens.  C program-

ming: More music minus whatever, failures and successes. Dr. Dobb’s Journal of Software Tools, 26(12):102– 105, December 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_12/
cpro1201.txt;   http://www.
ddj.com/ftp/2001/2001_12/
cpro1201.zip.

Stevens:2001:CPP

[Ste01g] Al Stevens. C programming:

Play Mysty for me, I’m on the QT with GCC. Dr. Dobb’s Journal of Software Tools, 26 (9):107–109, 111, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPR

[Ste01h]  Al Stevens.  C program-

ming: Road rage = editors + drivers. Dr. Dobb’s Journal of Software Tools, 26(6):131–134, June 2001. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPS

[Ste01i] Al Stevens. C programming:

Software development, Linux, and the White House. Dr. Dobb’s Journal of Software Tools, 26(3):124–127, March 2001. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPB

[Ste01j] Al Stevens. C programming: The best of the Penguin’s quest. Dr. Dobb’s Journal of Software Tools, 26(8):103, 105– 106, 109, August 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPYb

[Ste01k] Al Stevens. C programming: YAPP: Yet Another Programming Platform. Dr. Dobb’s Journal of Software Tools, 26 (10):105–107, 109, October 2001. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2001:CPYa

[Ste01l] Al Stevens. C programming: You’ve come a long way, baby. Dr. Dobb’s Journal of Software Tools, 26(1):173, 175– 176, January 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_01/
                                        cpro0101.txt.

Stevens:2002:CPab

[Ste02a] Al Stevens. C programming. Dr. Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2002:CPB

[Ste02b] Al Stevens. C programming: Back to the stick bass, … uh bull fiddle, … er contra base. whatever. Dr. Dobb’s Journal of Software Tools, 27(6):

82, 84–85, June 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_06/
cpro0602.txt.

Stevens:2002:CPCb

[Ste02c]  Al Stevens.  C program-

ming: C# and the outhouse paradigm. Dr. Dobb’s Journal of Software Tools, 27(7):79–81, July 2002. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2002:CPH

[Ste02d] Al Stevens. C programming:

Hungarians and LANs. Dr. Dobb’s Journal of Software Tools, 27(5):73–75, May 2002. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2002:CPc

[Ste02e] Al Stevens. C programming:

If this is June, it must be Zurich. Dr. Dobb’s Journal of Software Tools, 27(9):77– 80, September 2002. CODEN DDJOEB. ISSN 1044-789X.

Stevens:2002:CPI

[Ste02f] Al Stevens. C programming:

InstallShield Developer and another ride on the waves. Dr. Dobb’s Journal of Software Tools, 27(4):89–91, April 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2002/2002_04/
cpro0402.zip.

Stevens:2002:CPJ

[Ste02g]  Al Stevens.  C program-

ming: Jack? who’s Jack?

Dr. Dobb’s Journal of Software Tools, 27(10):68, 70, 72, 74, October 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_10/
cpro1002.zip.

Stevens:2002:CPPb

[Ste02h] Al Stevens. C programming: Paying the pipers. Dr. Dobb’s Journal of Software Tools, 27 (12):64, 68, 70, December 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7718/ddj0212h/.

Stevens:2002:CPPa

[Ste02i]  Al Stevens.  C program-

ming: Plug it in, plug it in. Dr. Dobb’s Journal of Software Tools, 27(1):82– 85, January 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_01/
cpro0102.zip.

Stevens:2002:CPR

[Ste02j] Al Stevens. C programming: Riding the waves. Dr. Dobb’s Journal of Software Tools, 27 (2):87–89, February 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_02/
cpro0202.zip.

Stevens:2002:CPW

[Ste02k] Al Stevens. C programming: Wave repairs and a first look at C#. Dr. Dobb’s Journal of Software Tools, 27(3):77– 78, 80, March 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_03/
cpro0302.txt;   http://www.
ddj.com/ftp/2002/2002_03/
cpro0302.zip.

Stevens:2003:CPU

[Ste03a] Al Stevens. C programming:

Underdocumented C++. Dr. Dobb’s Journal of Software Tools, 28(4):72, 74, 76, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7827/ddj0304i/.

Stevens:2003:EJT

[Ste03b] Al Stevens. Epicenter Johnny and typename exposed. Dr. Dobb’s Journal of Software Tools, 28(2):66, 68–69, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7790/ddj0302i/.

Stob:2001:NAV

[Sto01] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 26(12):18, December 2001. CODEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVa

[Sto02a] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 27(1):16, January 2002. CODEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVb

[Sto02b] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 27(2):16, February 2002. CODEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVc

[Sto02c] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 27(3):12, March 2002. CODEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVd

[Sto02d] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 27(4):14, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVe

[Sto02e] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 27(5):12, May 2002. CODEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVf

[Sto02f] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 27(6):12, June 2002. CODEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVg

[Sto02g] Verity Stob. The new adventures of Verity Stob. Dr.

Dobb’s Journal of Software Tools, 27(7):12, July 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVh

[Sto02h] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVi

[Sto02i] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 27(9):12, September 2002. CODEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVj

[Sto02j] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 27(10):12, October 2002. CODEN DDJOEB. ISSN 1044-789X.

Stob:2002:NAVk

[Sto02k] Verity Stob. The new adventures of Verity Stob: Idle thoughts of an idle process. Dr. Dobb’s Journal of Software Tools, 27(11):12, November 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7644/ddj0211m/.

Stob:2002:NAVl

[Sto02l] Verity Stob. The new adventures of Verity Stob: With and without Fin[n]ish. Dr. Dobb’s Journal of Software Tools, 27 (12):12, December 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7718/
ddj0212m/.

Stob:2003:NAV

[Sto03a] Verity Stob. The new adventures of Verity Stob. Dr. Dobb’s Journal of Software Tools, 28(7):12, July 2003. CO-DEN DDJOEB. ISSN 1044-789X.

Stob:2003:DR

[Sto03b] Verity Stob. The new adventures of Verity Stob: Damnation without relief. Dr. Dobb’s Journal of Software Tools, 28(6):12, June 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=8213/
ddj0306m/.

Stob:2003:NAVd

[Sto03c] Verity Stob. The new adventures of Verity Stob: Print job. Dr. Dobb’s Journal of Software Tools, 28(4):14, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7827/ddj0304n/.

Stob:2003:NAVb

[Sto03d] Verity Stob. The new adventures of Verity Stob: Revenge of the anti-patterns. Dr. Dobb’s Journal of Software Tools, 28(2):12, February 2003. CODEN DDJOEB.

ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7790/ddj0302n/.

Stob:2003:NAVe

[Sto03e] Verity Stob. The new adventures of Verity Stob: Soundtrack. Dr. Dobb’s Journal of Software Tools, 28(5):12, May 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7927/ddj0305n/.

Stob:2003:NAVf

[Sto03f] Verity Stob. The new adventures of Verity Stob: Stoblog. Dr. Dobb’s Journal of Software Tools, 28(1):12, January 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7758/ddj0301m/.

Stob:2003:NAVc

[Sto03g] Verity Stob. The new adventures of Verity Stob: We don’t guarantee that the latest Windows API feature will make you sexy. Dr. Dobb’s Journal of Software Tools, 28 (3):12, March 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7826/
ddj0303m/.

Strunk:2001:JQJ

[Str01] Elisabeth Strunk. Java Q&A:

Java & NT authentication. Dr. Dobb’s Journal of Software Tools, 26(2):145–146, 148, February 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2001/2001_02/jqa0201.
                                        txt.

Sutter:2000:MN

[Sut00] Herb Sutter. Migrating to namespaces. Dr. Dobb’s Journal of Software Tools, 25(10): 48, 50, 52, October 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_10/
                                        namespac.txt.

Swaine:2000:PPAb

[Swa00a] Michael Swaine. Programming paradigms: All those zombies. Dr. Dobb’s Journal of Software Tools, 25(10):117–120, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPAa

[Swa00b] Michael Swaine. Programming paradigms: Atom by atom. Dr. Dobb’s Journal of Software Tools, 25(3):101–102, 104–105, March 2000. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPB

[Swa00c] Michael Swaine. Programming paradigms: Breadth-first search. Dr. Dobb’s Journal of Software Tools, 25(6):100, 102, 105, June 2000. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPCa

[Swa00d] Michael Swaine. Programming paradigms: Change is good, but folding money is better. Dr. Dobb’s Journal of Software Tools, 25(1):89, 90–92, January 2000. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPCb

[Swa00e] Michael Swaine. Programming paradigms: Childhood’s end. Dr. Dobb’s Journal of Software Tools, 25(11):127–130, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPD

[Swa00f] Michael Swaine. Programming paradigms: Declarative programming revisited. Dr. Dobb’s Journal of Software Tools, 25(8):113–117, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPM

[Swa00g] Michael Swaine. Programming paradigms: Misinformation. Dr. Dobb’s Journal of Software Tools, 25(4):91– 93, 95, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPO

[Swa00h] Michael Swaine. Programming paradigms: Oeuvre interest. Dr. Dobb’s Journal of Software Tools, 25(12):123–126, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPe

[Swa00i] Michael Swaine. Programming paradigms: Oh behave! Dr. Dobb’s Journal of Software Tools, 25(5):99–100, 102, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPPb

[Swa00j] Michael Swaine. Programming paradigms: Paradigms lost and regained. Dr. Dobb’s Journal of Software Tools, 25(9): 109–112, September 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPg

[Swa00k] Michael Swaine. Programming paradigms: The cluetrain antipasto. Dr. Dobb’s Journal of Software Tools, 25(7):107–110, July 2000. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2000:PPPa

[Swa00l] Michael Swaine. Programming paradigms: The palm, the nose, and other computing platforms. Dr. Dobb’s Journal of Software Tools, 25(2):105, 107–108, February 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:2000:RMS

[Swa00m] Michael Swaine. Rebol and Email services. Dr. Dobb’s Journal of Software Tools, 25(1):28, 30, 32–34, January 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_01/
                                        rebol.txt.

Swaine:2000:SFA

[Swa00n]  Michael Swaine.  Swaine’s

flames: Acrophilia. Dr. Dobb’s Journal of Software Tools, 25 (11):168, November 2000. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/privacy.htm.

Swaine:2000:SFB

[Swa00o]  Michael Swaine.  Swaine’s

flames: Being a blip. Dr. Dobb’s Journal of Software Tools, 25(1):128, January 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ddj/privacy.
htm.

Swaine:2000:SFC

[Swa00p]  Michael Swaine.  Swaine’s

flames: Carly’s way. Dr. Dobb’s Journal of Software Tools, 25(12):168, December 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ddj/privacy.
htm.

Swaine:2000:SFDb

[Swa00q]  Michael Swaine.  Swaine’s

flames: Disgusting Canadians. Dr. Dobb’s Journal of Software Tools, 25(10):168, October 2000. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ddj/privacy.
htm.

Swaine:2000:SFE

[Swa00r]  Michael Swaine.  Swaine’s

flames: Error-correcting journalism. Dr. Dobb’s Journal of Software Tools, 25 (6):152, June 2000. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/privacy.htm.

Swaine:2000:SFI

[Swa00s]  Michael Swaine.  Swaine’s

flames: Invention or hack? Dr. Dobb’s Journal of Software Tools, 25(3):144, March 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ddj/privacy.htm.

Swaine:2000:SFNa

[Swa00t]  Michael Swaine.  Swaine’s

flames: Nag, nag, nag. Dr. Dobb’s Journal of Software Tools, 25(5):136, May 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ddj/privacy.htm.

Swaine:2000:SFNb

[Swa00u]  Michael Swaine.  Swaine’s

flames: No formal wear, no nuns, no chilly willy. Dr. Dobb’s Journal of Software Tools, 25(8):152, August 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ddj/privacy.htm.

Swaine:2000:SFS

[Swa00v]  Michael Swaine.  Swaine’s

flames: Scaffolding. Dr. Dobb’s Journal of Software Tools, 25 (2):144, February 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ddj/privacy.htm.

Swaine:2000:SFDa

[Swa00w]  Michael Swaine.  Swaine’s

flames: The death and resurrection of Rudy Doone. Dr. Dobb’s Journal of Software Tools, 25(4):136, April 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ddj/privacy.htm.

Swaine:2000:SFR

[Swa00x]  Michael Swaine.  Swaine’s

flames: The rant at Foo Bar. Dr. Dobb’s Journal of Software Tools, 25(7):152, July 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ddj/privacy.htm.

Swaine:2000:SFT

[Swa00y]  Michael Swaine.  Swaine’s

flames: The tenors. Dr. Dobb’s Journal of Software Tools, 25 (9):152, September 2000. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ddj/privacy.htm.

Swaine:2001:PPA

[Swa01a] Michael Swaine. Programming paradigms: Alternatives to Java. Dr. Dobb’s Journal of Software Tools, 26(2):135– 137, February 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPB

[Swa01b] Michael Swaine. Programming paradigms: Back to the future. Dr. Dobb’s Journal of Software Tools, 26(8):97, 99–101, August 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPEa

[Swa01c] Michael Swaine. Programming paradigms: Every install is five-sevenths stall. Dr. Dobb’s Journal of Software Tools, 26 (6):125, 127–129, June 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPM

[Swa01d] Michael Swaine. Programming paradigms: Mostly math. Dr. Dobb’s Journal of Software Tools, 26(3):121–123, March 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPO

[Swa01e] Michael Swaine. Programming paradigms: Other worlds. Dr. Dobb’s Journal of Software Tools, 26(10):99–102, October 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPRb

[Swa01f] Michael Swaine. Programming paradigms: Realizable fantasies. Dr. Dobb’s Journal of Software Tools, 26(9):103– 105, September 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPRc

[Swa01g] Michael Swaine. Programming paradigms: Realizable fantasies and conspiracy theories. Dr. Dobb’s Journal of Software Tools, 26(11):99– 101, November 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPRa

[Swa01h] Michael Swaine. Programming paradigms: Roll over, Kemeny and Kurtz. Dr. Dobb’s Journal of Software Tools, 26(5):

115, 117, 119, May 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPSa

[Swa01i] Michael Swaine. Programming paradigms: Secrets and lies. Dr. Dobb’s Journal of Software Tools, 26(4):125– 127, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPSb

[Swa01j] Michael Swaine. Programming paradigms: Size matters. Dr. Dobb’s Journal of Software Tools, 26(12):95–96, 98, December 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPEb

[Swa01k] Michael Swaine. Programming paradigms: The enduring popularity of Basic, the true meaning of the Semantic Web, and other puzzles. Dr. Dobb’s Journal of Software Tools, 26(7):105–106, 109–110, July 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:PPW

[Swa01l] Michael Swaine. Programming paradigms: Worthless gifts? Dr. Dobb’s Journal of Software Tools, 26(1):167–168, 170, January 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:SFB

[Swa01m]  Michael Swaine.  Swaine’s

flames: Breakfast with Burnout Boy. Dr. Dobb’s Journal of Software Tools, 26(7):144, July 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:SFF

[Swa01n]  Michael Swaine.  Swaine’s

flames: Flaw and order. Dr. Dobb’s Journal of Software Tools, 26(8):128, August 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ddj/privacy.htm;
                                         http://www.ddj.com/ddj/
                                        sdmg/.

Swaine:2001:SFGa

[Swa01o]  Michael Swaine.  Swaine’s

flames: Getting skinned. Dr. Dobb’s Journal of Software Tools, 26(4):160, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ddj/privacy.htm;
                                         http://www.ddj.com/ddj/
                                        sdmg/.

Swaine:2001:SFM

[Swa01p]  Michael Swaine.  Swaine’s

flames: Mau-mauing the ambulance chasers. Dr. Dobb’s Journal of Software Tools, 26 (9):136, September 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        cmp.com/delivery/privacy.
                                        html.

Swaine:2001:SFPb

[Swa01q]  Michael Swaine.  Swaine’s

flames: Poverty on parade. Dr. Dobb’s Journal of Software Tools, 26(11):136, November 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.cmp.com/
delivery/privacy.html.

Swaine:2001:SFR

[Swa01r]  Michael Swaine.  Swaine’s

flames: Read me last. Dr. Dobb’s Journal of Software Tools, 26(1):208, January 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ddj/privacy.
htm;   http://www.ddj.com/
ddj/sdmg/.

Swaine:2001:SFS

[Swa01s]  Michael Swaine.  Swaine’s

flames: Saving Private Lyin’. Dr. Dobb’s Journal of Software Tools, 26(2):168, February 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ddj/privacy.
htm;   http://www.ddj.com/
ddj/sdmg/.

Swaine:2001:SFGb

[Swa01t]  Michael Swaine.  Swaine’s

flames: The geek girls and the big-endian conspiracy. Dr. Dobb’s Journal of Software Tools, 26(10): 136, October 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.cmp.com/
delivery/privacy.html.

Swaine:2001:SFPa

[Swa01u]  Michael Swaine.  Swaine’s

flames: The PlayStations of Saddam Hussein. Dr. Dobb’s Journal of Software Tools, 26 (3):152, March 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ddj/privacy.htm;   http:/
                                        /www.ddj.com/ddj/sdmg/.

Swaine:2001:SFT

[Swa01v]  Michael Swaine.  Swaine’s

flames: The taxidermist and the flapper. Dr. Dobb’s Journal of Software Tools, 26(12): 128, December 2001. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.cmp.com/
                                        delivery/privacy.html.

Swaine:2001:SFZ

[Swa01w]  Michael Swaine.  Swaine’s

flames: The Zappa meme. Dr. Dobb’s Journal of Software Tools, 26(5):160, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2001:SFW

[Swa01x]  Michael Swaine.  Swaine’s

flames: Whistling in the dark. Dr. Dobb’s Journal of Software Tools, 26(6):168, June 2001. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:IBM

[Swa02a] Michael Swaine. It brings on many changes. Dr. Dobb’s Journal of Software Tools, 27 (12):??, December 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:2002:PPad

[Swa02b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 27(8): ??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:PP

[Swa02c] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 27(10): ??, October 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:PPQ

[Swa02d] Michael Swaine. Programming paradigms: A quantum computer and a scripting studio. Dr. Dobb’s Journal of Software Tools, 27(3):71–72, 74, March 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:PPB

[Swa02e] Michael Swaine. Programming paradigms: Beautiful math. Dr. Dobb’s Journal of Software Tools, 27(6):79–81, June 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:PPC

[Swa02f] Michael Swaine. Programming paradigms: Community. Dr. Dobb’s Journal of Software Tools, 27(11):73–75, November 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7644/ddj0211h/.

Swaine:2002:PPD

[Swa02g] Michael Swaine. Programming paradigms: Digital hubbub and the wayback machine. Dr. Dobb’s Journal of Software Tools, 27(1):76, 78, 80, January 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:PPE

[Swa02h] Michael Swaine. Programming paradigms: Embracing anarchy. Dr. Dobb’s Journal of Software Tools, 27(5):69–71, May 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:PPF

[Swa02i] Michael Swaine. Programming paradigms: Faux fear. Dr. Dobb’s Journal of Software Tools, 27(2):83–85, February 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:PPS

[Swa02j] Michael Swaine. Programming paradigms: Strategy and secrecy. Dr. Dobb’s Journal of Software Tools, 27(4): 85–87, April 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:PPWa

[Swa02k] Michael Swaine. Programming paradigms: Waiting for QuantumBasic 1.0. Dr. Dobb’s Journal of Software Tools, 27 (7):75–77, July 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:PPWb

[Swa02l] Michael Swaine. Programming paradigms: Wolfram’s computational equivalence. Dr. Dobb’s Journal of Software Tools, 27(9):71, 73–75, September 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:SFA

[Swa02m]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 27(6): 104, June 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.cmp.com/
delivery/privacy.html.

Swaine:2002:SFHb

[Swa02n]  Michael Swaine.  Swaine’s

flames. Dr. Dobb’s Journal of Software Tools, 27(8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:SFDa

[Swa02o]  Michael Swaine.  Swaine’s

flames: Dirges in the dark. Dr. Dobb’s Journal of Software Tools, 27(4):112, April 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

cmp.com/delivery/privacy.
html.

Swaine:2002:SFDb

[Swa02p]  Michael Swaine.  Swaine’s

flames: Do you copy? Dr. Dobb’s Journal of Software Tools, 27(11):84, November 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:SFM

[Swa02q]  Michael Swaine.  Swaine’s

flames: E-mails of the Fortune Five Million, #1. Dr. Dobb’s Journal of Software Tools, 27 (1):112, January 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
cmp.com/delivery/privacy.
html.

Swaine:2002:SFE

[Swa02r]  Michael Swaine.  Swaine’s

flames: Eating my words. Dr. Dobb’s Journal of Software Tools, 27(5):104, May 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        cmp.com/delivery/privacy.
                                        html.

Swaine:2002:SFF

[Swa02s]  Michael Swaine.  Swaine’s

flames: Falling in place. Dr. Dobb’s Journal of Software Tools, 27(2):112, February 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.cmp.com/delivery/privacy.
                                        html.

Swaine:2002:SFK

[Swa02t]  Michael Swaine.  Swaine’s

flames: Kicking the habit. Dr. Dobb’s Journal of Software Tools, 27(7):104, July 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:SFL

[Swa02u]  Michael Swaine.  Swaine’s

flames: Looking for zebras. Dr. Dobb’s Journal of Software Tools, 27(12):84, December 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7718/ddj0212p/.

Swaine:2002:SFH

[Swa02v]  Michael Swaine.  Swaine’s

flames: Request for disposal. Dr. Dobb’s Journal of Software Tools, 27(10):92, October 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:SFHc

[Swa02w]  Michael Swaine.  Swaine’s

flames: The big bug summit. Dr. Dobb’s Journal of Software Tools, 27(9):104, September 2002. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2002:SFP

[Swa02x]  Michael Swaine.  Swaine’s

flames: The pickle and the printout. Dr. Dobb’s Journal of Software Tools, 27(3): 112, March 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.cmp.com/
delivery/privacy.html.

Swaine:2003:LME

[Swa03a] Michael Swaine. Last Macworld Expo. Dr. Dobb’s Journal of Software Tools, 28(2): ??, February 2003. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2003:PPAb

[Swa03b] Michael Swaine. Programming paradigms: Adam Osborne. Dr. Dobb’s Journal of Software Tools, 28(7):68–70, July 2003. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2003:PPAa

[Swa03c] Michael Swaine. Programming paradigms: Attacks and accidents. Dr. Dobb’s Journal of Software Tools, 28(1): 61–63, January 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.

                                        ddj.com/documents/s=7758/
                                        ddj0301h/.

Swaine:2003:PPH

[Swa03d] Michael Swaine. Programming paradigms: History lessons. Dr. Dobb’s Journal of Software Tools, 28(3):60, 62–63, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7826/ddj0303h/.

Swaine:2003:PPI

[Swa03e] Michael Swaine. Programming paradigms: Interesting times. Dr. Dobb’s Journal of Software Tools, 28(5):77–79, May 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7927/ddj0305i/.

Swaine:2003:PPJ

[Swa03f] Michael Swaine. Programming paradigms: Just released. Dr. Dobb’s Journal of Software Tools, 28(6):68, 70–71, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        8213/ddj0306h/.

Swaine:2003:SFB

[Swa03g]  Michael Swaine.  Swaine’s

flames: Big Brother and the Holding Company. Dr. Dobb’s Journal of Software Tools, 28 (3):72, March 2003. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2003:SFJ

[Swa03h]  Michael Swaine.  Swaine’s

flames: Jerry and the Beeb.

Dr. Dobb’s Journal of Software Tools, 28(1):76, January 2003. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2003:MMM

[Swa03i]  Michael Swaine.  Swaine’s

flames: Man-month math. Dr. Dobb’s Journal of Software Tools, 28(2):80, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7790/ddj0302q/.

Swaine:2003:SFM

[Swa03j]  Michael Swaine.  Swaine’s

flames: Minor essays poorly connected. Dr. Dobb’s Journal of Software Tools, 28(7):80, July 2003. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2003:SFP

[Swa03k]  Michael Swaine.  Swaine’s

flames: Prose sampling #1. Dr. Dobb’s Journal of Software Tools, 28(6):80, June 2003. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2003:RMA

[Swa03l]  Michael Swaine.  Swaine’s

flames: Renewing my artistic license. Dr. Dobb’s Journal of Software Tools, 28(5):88, May 2003. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2003:SFT

[Swa03m]  Michael Swaine.  Swaine’s

flames: Tales from the security battlefronts. Dr. Dobb’s Journal of Software Tools, 28 (4):88, April 2003. CODEN DDJOEB. ISSN 1044-789X.

Swaine:2006:VR

[Swa06] Michael Swaine. Is virtualization real? Dr. Dobb’s Journal of Software Tools, 31(12): 18–19, 21–22, December 2006. CODEN DDJOEB. ISSN 1044-789X.

Schneier:2001:BIP

[SZ01] Bruce Schneier and Elizabeth Zwicky. Banners and Internet protocols. Dr. Dobb’s Journal of Software Tools, 26(11): 38, 41, November 2001. CO-DEN DDJOEB. ISSN 1044-789X.

Tan:2001:JWP

[Tan01] C. K. Tan. A JDBC wrapper — in Python! Dr. Dobb’s Journal of Software Tools, 26 (8):50, 52, 54, August 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2001/2001_08/
                                        jdbcpy.txt;   http://www.
                                        ddj.com/ftp/2001/2001_08/
                                        jdbcpy.zip.

Taylor:2000:PBS

[Tay00] Jeffrey L. Taylor. Programmer’s bookshelf: Second time around. Dr. Dobb’s Journal of Software Tools, 25(4):127– 128, 130, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Taylor:2001:PHN

[Tay01] Jeffrey Taylor. Precompiled headers and normal C++ semantics. Dr. Dobb’s Journal of Software Tools, 26(10):36, 38, October 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_10/headers.
txt.

Taylor:2002:PBL

[Tay02] Jeffrey L. Taylor. Programmer’s bookshelf: Lightweight processes for changing environments. Dr. Dobb’s Journal of Software Tools, 27(11): 82, November 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7644/
ddj0211j/.

Truong:2000:MCP

[TC00] Derek Truong and Allen Chan.

Measuring C++ program efficiency. Dr. Dobb’s Journal of Software Tools, 25(10):62, 64–67, October 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_10/
cppmetric.zip.

Tucker:2000:LEP

[TCM^{+}00] Andrew Tucker, Edoardo Comar, Scott Meyers, Yves Piguet, Kevin Ruland, Greg Hadaller, Jonathan Erickson, Mike Zhilin, and Todd Stephan. Letters: Editor preferences; Java enums; labor union harassment; smart pointer update; traveling salesman; granting block grant; porting to CE; analyzing algorithms. Dr. Dobb’s Journal of Software Tools, 25(1):

10, 12, January 2000. CODEN DDJOEB. ISSN 1044-789X.

Todd:2001:LSS

[TEM^{+}01] Andrew W. Todd, Jonathan Erickson, Nadine McKenzie, Chris Cleeland, Richard Huang, Ragae Ghaly, and The Editors. Letters: Shared source and shared secrets; JavaScript fix; CORBA interoperability; EJB application servers update; correction [“The Delphi XML SAX2 Component and MSXML 3.0”]. Dr. Dobb’s Journal of Software Tools, 26(10):10, 12, October 2001. CODEN DDJOEB. ISSN 1044-789X. See [Hei01].

Tamassia:2001:JDS

[TGV^{+}01] Roberto Tamassia, Michael T. Goodrich, Luca Vismara, Mark Handy, Galina Shubina, Robert Cohen, Benoît Hudson, Ryan S. Baker, Natasha Gelfand, and Ulrik Brandes. JDSL: The data structures library in Java. Dr. Dobb’s Journal of Software Tools, 26(4):21–22, 24, 26, 28, 30–31, April 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_04/
                                        jdsl.txt.

Thomas:2001:PR

[TH01] Dave Thomas and Andy Hunt. Programming in Ruby. Dr. Dobb’s Journal of Software Tools, 26(1):44, 46, 48– 51, January 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_01/
ruby.txt.

Thomas:2000:EOS

[Tho00] Gary Thomas. eCos: An operating system for embedded systems. Dr. Dobb’s Journal of Software Tools, 25(1):66, 68– 72, 74, January 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_01/
ecos.txt.

Thomas:2002:PTE

[Tho02] Chris Thomas. Programmer’s toolchest: Examining Ruby-Cocoa. Dr. Dobb’s Journal of Software Tools, 27 (5):64–67, May 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_05/
rubycoco.txt;   http://www.
ddj.com/ftp/2002/2002_05/
rubycoco.zip.

Timmerman:2003:EWC

[Tim03] Martin Timmerman. Examining Windows CE .NET. Dr. Dobb’s Journal of Software Tools, 28(2):62, 64, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7790/ddj0302h/.

Editors:2002:LUC

[TLA^{+}02] The Editors, Kim Reidar Lantz, Ze’ev Atlas, Pete Nelson, and Gus J. Grubba. Letters: URL correction [“The NewOS Operating System”]; passing context to threads; compiling Perl/Tk scripts; standing by Al’s principles; understanding photomosaics. Dr. Dobb’s Journal of Software Tools, 27(1):10, 12, January 2002. CODEN DDJOEB. ISSN 1044-789X. See [Gei01].

Toran-Marti:2002:SNI

[TMVTdM02] Felix Toran-Marti, Javier Ventura-Traveset, and Juan Carlos de Mateo. Satellite navigation and the Internet. Dr. Dobb’s Journal of Software Tools, 27(3):17–18, 20, 22, 24, 26, March 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_03/
satelite.txt;   http://www.
ddj.com/ftp/2002/2002_03/
satelite.zip.

Tomson:2002:DCR

[Tom02]  Phil Tomson.  Distributed

computing with Ruby. Dr. Dobb’s Journal of Software Tools, 27(9):16, 18, 22, September 2002. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2002/2002_09/taskmas.
txt;   http://www.ddj.com/
ftp/2002/2002_09/taskmas.
zip.

Torres:2001:DSD

[Tor01] Mario A. Torres. Developing scalable distributed applications: A generic model implemented in Java. Dr. Dobb’s Journal of Software Tools, 26(9):21–22, 24–25, September 2001. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_09/
scalable.txt.

To:2002:ESR

[TP02] Richard To and Cara Pang.

External SQL rewriters. Dr. Dobb’s Journal of Software Tools, 27(5):42, 44–48, May 2002. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/ftp/2002/2002_
05/sql.txt.

Travers:2000:JQW

[Tra00] Michael Travers. Java Q&A:

What is interactive scripting? Dr. Dobb’s Journal of Software Tools, 25(1):103, 105–111, January 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_01/
jqa120.txt.

Tremblett:2002:JUR

[Tre02a] Paul Tremblett. Java and UDDI registries. Dr. Dobb’s Journal of Software Tools, 27 (9):34, 37–40, September 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2002/2002_09/
juddi.txt;   http://www.
ddj.com/ftp/2002/2002_09/
juddi.zip;   http://www.
ddj.com/ftp/2002/2002_09/
uddi.war.

Tremblett:2002:PTJ

[Tre02b] Paul Tremblett. Programmer’s toolchest: Java’s Generic Connection Framework: Wireless networking for mobile devices. Dr. Dobb’s Journal of Software Tools, 27(4):75– 78, 80, April 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_04/javagcf.
                                        txt;   http://www.ddj.com/
                                        ftp/2002/2002_04/javagcf.
                                        zip.

Trueman:2000:VNR

[Tru00] Chris Trueman. Visualizing network resources using Visio. Dr. Dobb’s Journal of Software Tools, 25(2):98, 100– 102, February 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_02/
                                        visioviz.txt;   http://www.
                                        ddj.com/ftp/2000/2000_02/
                                        visioviz.zip.

Trudell:2001:RTC

[Tru01]  Bill Trudell.  Redirection

through C++ function pointers. Dr. Dobb’s Journal of Software Tools, 26(12):56, 58, 60, 62–64, December 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2001/2001_12/
                                        redirect.txt.

Trueman:2002:VIE

[Tru02] Christopher J. Trueman. Verifying installation environments. Dr. Dobb’s Journal of Software Tools, 27(6):46, 48, 50–52, June 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_06/
echeck.txt;   http://www.
ddj.com/ftp/2002/2002_06/
echeck.zip.

Tucker:2000:WCC

[Tuc00] Andrew Tucker. Window CE’s CESH utility. Dr. Dobb’s Journal of Software Tools, 25(5): 74, 76–78, 80, May 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_05/
cecesh.txt;   http://www.
ddj.com/ftp/2000/2000_05/
cecesh.zip.

Turjanmaa:2001:MOS

[Tur01] Ville Mikael Turjanmaa. The Menuet operating system. Dr. Dobb’s Journal of Software Tools, 26(12):32, 36, December 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_12/msetup.
zip.

VanBeneden:2001:EWC

[Van01] Bart Van Beneden. Examining Windows CE 3.0 real-time capabilities. Dr. Dobb’s Journal of Software Tools, 26(12):66, 68–69, 71–72, December 2001. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2001/2001_12/
wince3.txt.

VanBeneden:2002:EQR

[Van02a] Bart Van Beneden. Examining QNX RTOS 6.1. Dr. Dobb’s Journal of Software Tools, 27 (6):54, 56, 58, 60, June 2002. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2002/2002_06/
                                        qnx61.txt.

VanBeneden:2002:EVA

[Van02b] Bart Van Beneden. Examining the VxWorks AE 1.1 RTOS. Dr. Dobb’s Journal of Software Tools, 27(11):66, 68, 70, November 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7644/
                                        ddj0211g/.

Vaingast:2001:ESI

[VC01] Shai Vaingast and Ehud Cohen. DOS for embedded systems: Interrupt latencies. Dr. Dobb’s Journal of Software Tools, 26(1):138, 140–142, 144, 146, January 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X. URL http://www.ddj.
                                        com/ftp/2001/2001_01/int_
                                        late.txt.

Venema:2000:FRT

[Ven00a] Wietse Venema. File recovery techniques. Dr. Dobb’s Journal of Software Tools, 25(12):74, 76–80, December 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Venema:2000:SN

[Ven00b] Wietse Venema. Strangers in the night. Dr. Dobb’s Journal of Software Tools, 25(11):82, 84–88, November 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Vidales:2000:DEI

[VF00] Carlos E. Vidales and Eugene M. Farrelly. The Digita OS: An extensible imaging platform. Dr. Dobb’s Journal of Software Tools, 25(12):82, 84, 86, 89–90, December 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/ftp/2000/2000_12/
digitaos.txt.

Vilar:2000:JQW

[Vil00] Alexandre Sieira Vilar. Java Q&A: What’s object pooling all about? Dr. Dobb’s Journal of Software Tools, 25(8):123–124, 126, 128, 130, August 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_08/
jqa800.txt;   http://www.
ddj.com/ftp/2000/2000_08/
jqa800.zip.

Vlissides:2000:MC

[Vli00] John Vlissides. McConnell complete. Dr. Dobb’s Journal of Software Tools, 25(10):76, 78, 80–81, October 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Vandevoorde:2003:LCT

[VMNH03] David (“Daveed”) Vandevoorde, Doug McGarrett, Ed Nisley, and Jim Hunter. Letters: C++’s typename keyword; more VB versus the world; RF band letter designations; happy with XSLT. Dr. Dobb’s Journal of Software Tools, 28(4):12, April 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7827/
ddj0304m/.

VanTassle:2002:WNB

[VR02] Ray Van Tassle and Marjorie Richardson. Wireless networking and Berkeley DB. Dr. Dobb’s Journal of Software Tools, 27(1): 50–53, January 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_01/
wireless.txt.

Wadler:2000:GGJ

[Wad00a] Philip Wadler. GJ: A Generic Java. Dr. Dobb’s Journal of Software Tools, 25(2):23–26, 28, February 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_02/gj.txt.

Wadler:2000:NLO

[Wad00b] Philip Wadler. New languages, old logic. Dr. Dobb’s Journal of Software Tools, 25(12 (supplement)):37–41, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Walsh:2002:MJA

[Wal02a]  Aaron E. Walsh.  The

MPEG-4 Java API and MPE-Glets. Dr. Dobb’s Journal of Software Tools, 27(4): 40, 42–45, April 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_04/
jmpeg.txt;   http://www.
ddj.com/ftp/2002/2002_04/
jmpeg.zip.

Walsh:2002:USG

[Wal02b] Aaron E. Walsh. Understanding scene graphs. Dr. Dobb’s Journal of Software Tools, 27 (7):17, 20–26, July 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_07/
scene.txt;   http://www.
ddj.com/ftp/2002/2002_07/
scene.zip.

Walsh:2003:JWS

[Wal03a] Aaron E. Walsh. J2Ee 1.4 Web services. Dr. Dobb’s Journal of Software Tools, 28(4):34, 36, 38, 39, April 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7827/
ddj0304d/.

Walsh:2003:JP

[Wal03b] Aaron E. Walsh. JavaServer Pages 2.0. Dr. Dobb’s Journal of Software Tools, 28(7): 48, 50, 52–57, July 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.

                                        ddj.com/ftp/2003/2003_07/
                                        jsp20.zip.

Wanner:2000:SEY

[Wan00] James F. Wanner. Source-Monitor: Exposing your code. Dr. Dobb’s Journal of Software Tools, 25(3):92, 94–98, March 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_03/
                                        sourcemonitor.zip;   http:
                                        //www.ddj.com/ftp/2000/2000_
                                        03/sourcemonitorv0163.exe.

Anonymous:2003:Le

[WDGO03] David Wright, Stan Day, Kurt Guntheroth, and Larry O’Brien. Letters: 13 ways to loath VB; business process outsourcing; copy protection. Dr. Dobb’s Journal of Software Tools, 28(5):10, May 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/documents/s=
                                        7927/ddj0305m/.

Webb:2000:MFC

[Web00]  Chilton Webb.  The Ma-

cOS Find-By-Content Engine. Dr. Dobb’s Journal of Software Tools, 25(12):62–63, December 2000. CODEN DDJOEB. ISSN 1044-789X.

Wendt:2000:MP

[Wen00]  David Wendt.  Multilan-

guage programming. Dr. Dobb’s Journal of Software Tools, 25(11):68, 70–74, 76, November 2000. CODEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_11/
mldll.txt;   http://www.
ddj.com/ftp/2000/2000_11/
mldll.zip.

Woehr:2000:PBA

[WG00] Jack J. Woehr and Lou Grinzo.

Programmer’s bookshelf: All the world’s A database. Dr. Dobb’s Journal of Software Tools, 25(1):119–120, January 2000. CODEN DDJOEB. ISSN 1044-789X.

White:2001:DPP

[Whi01] Ken White. DBMS: Past, present, and future: Where databases have been, where they are going, and what it means to you. Dr. Dobb’s Journal of Software Tools, 26 (8):21–22, 24, 26, 28–29, August 2001. CODEN DDJOEB. ISSN 1044-789X.

Whittaker:2002:TSS

[Whi02] James A. Whittaker. Testing for software security. Dr. Dobb’s Journal of Software Tools, 27(11):24, 26, 28, 32, 34, November 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7644/
ddj0211c/.

White:2003:UTL

[Whi03] Tom White. Using thread-local variables in Java. Dr. Dobb’s Journal of Software Tools, 28 (7):42, 44–46, July 2003. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2003/2003_07/
                                        thread.txt;   http://www.
                                        ddj.com/ftp/2003/2003_07/
                                        thread.zip.

Wiener:2000:AAH

[Wie00] Michael J. Wiener. Algorithm alley: High-speed cryptography with the RSA algorithm. Dr. Dobb’s Journal of Software Tools, 25(2): 123–126, February 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_02/
                                        aa220.txt.

Williams:2000:TII

[Wil00a] Al Williams. The TINI Internet interface. Dr. Dobb’s Journal of Software Tools, 25(10): 82, 84, 86, 88, October 2000. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

                                        ddj.com/ftp/2000/2000_10/
                                        tini.txt.

Wilson:2000:PBA

[Wil00b] Gregory V. Wilson. Programmer’s bookshelf: All this, and frequent flyer miles, too! Dr. Dobb’s Journal of Software Tools, 25(2):135–136, February 2000. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2000:PBC

[Wil00c] Gregory V. Wilson. Programmer’s bookshelf: Classics old and new. Dr. Dobb’s Journal of Software Tools, 25(11):159– 160, November 2000. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2000:PBH

[Wil00d] Gregory V. Wilson. Programmer’s bookshelf: How matters more than what. Dr. Dobb’s Journal of Software Tools, 25 (3):135–136, March 2000. CO-DEN DDJOEB. ISSN 1044-789X.

Wilson:2000:PBP

[Wil00e] Gregory V. Wilson. Programmer’s bookshelf: Python, C++, and other religions. Dr. Dobb’s Journal of Software Tools, 25(8):145–147, August 2000. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2000:PBS

[Wil00f] Gregory V. Wilson. Programmer’s bookshelf: Stirring the gene pool. Dr. Dobb’s Journal of Software Tools, 25(5): 127, 130, May 2000. CODEN DDJOEB. ISSN 1044-789X.

Wiltamuth:2000:CPL

[Wil00g] Scott Wiltamuth. The C# programming language. Dr. Dobb’s Journal of Software Tools, 25(10):21–22, 24, 26, 28, 30–32, October 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_10/
cshrp.txt.

Williams:2001:JWT

[Wil01a]  Al Williams.  Java and

the Waba toolkit. Dr. Dobb’s Journal of Software Tools, 26(2):92, 94, 96, 98, February 2001. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2001/2001_02/
                                        jwaba.txt.

Williams:2001:SSC

[Wil01b] Chris Williams. SCRAM: Software Configuration, Release, and Management. Dr. Dobb’s Journal of Software Tools, 26 (4):114, 116–118, 120, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2001:PBA

[Wil01c] Gregory V. Wilson. Programmer’s bookshelf: Alien worlds. Dr. Dobb’s Journal of Software Tools, 26(12):122, 124, December 2001. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2001:PBT

[Wil01d] Gregory V. Wilson. Programmer’s bookshelf: Time warps. Dr. Dobb’s Journal of Software Tools, 26(2):159–160, February 2001. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2001:PBX

[Wil01e] Gregory V. Wilson. Programmer’s bookshelf: XML and then some. Dr. Dobb’s Journal of Software Tools, 26(5): 151–152, May 2001. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2002:PB

[Wil02a] Gregory V. Wilson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 27 (8):??, August 2002. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2002:PBA

[Wil02b] Gregory V. Wilson. Programmer’s bookshelf: All over the map. Dr. Dobb’s Journal of Software Tools, 27(6):100, June 2002. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2002:PBD

[Wil02c] Gregory V. Wilson. Programmer’s bookshelf: Doubling down on the good and the okay. Dr. Dobb’s Journal of Software Tools, 27(3): 107, March 2002. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2002:PBF

[Wil02d] Gregory V. Wilson. Programmer’s bookshelf: Filling in the gaps. Dr. Dobb’s Journal of Software Tools, 27(5):100, May 2002. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2002:PBS

[Wil02e] Gregory V. Wilson. Programmer’s bookshelf: Sometimes you get what you want. Dr. Dobb’s Journal of Software Tools, 27(2):107–170, February 2002. CODEN DDJOEB. ISSN 1044-789X.

Williams:2003:ESP

[Wil03a] Al Williams. Embedded systems: Programmable logic and hardware. Dr. Dobb’s Journal of Software Tools, 28(5): 70, 72–75, May 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7927/
ddj0305h/.

Williams:2003:ICB

[Wil03b] Al Williams. Internet connectivity in a box. Dr. Dobb’s Journal of Software Tools, 28 (1):52, 54, 56, 58, January 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7758/ddj0301g/.

Wilson:2003:PB

[Wil03c] Gregory V. Wilson. Programmer’s bookshelf. Dr. Dobb’s Journal of Software Tools, 28 (7):75, July 2003. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2003:PBFb

[Wil03d] Gregory V. Wilson. Programmer’s bookshelf: Facts, fallacies, and numeric computing. Dr. Dobb’s Journal of Software Tools, 28(6):76, June 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
8213/ddj0306j/.

Wilson:2003:PBFa

[Wil03e] Gregory V. Wilson. Programmer’s bookshelf: Five skinny books. Dr. Dobb’s Journal of Software Tools, 28 (2):75, February 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7790/
ddj0302k/.

Wilson:2003:PBP

[Wil03f] Gregory V. Wilson. Programmer’s bookshelf: Parkinson’s other law. Dr. Dobb’s Journal of Software Tools, 28(4):83, April 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7827/
ddj0304k/.

Wilson:2003:PBO

[Wil03g] Gregory V. Wilson. Programmer’s bookshelf: The observation of trifles. Dr. Dobb’s Journal of Software Tools, 28 (1):73, January 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7758/
ddj0301j/.

Wilson:2003:PBT

[Wil03h] Gregory V. Wilson. Programmer’s bookshelf: Three from Manning. Dr. Dobb’s Journal of Software Tools, 28(5):84, May 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7927/ddj0305k/.

Wilson:2003:XBP

[Wil03i] Gregory V. Wilson. XMLbased programming systems. Dr. Dobb’s Journal of Software Tools, 28(3):16, 18, 20, 22, 24, March 2003. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/documents/s=7826/
ddj0303b/.

Wincelberg:2000:AAA

[Win00] David Wincelberg. Algorithm alley: Alphanumeric ordering. Dr. Dobb’s Journal of Software Tools, 25(10):147– 150, October 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_10/
                                        aa010.txt.

Wincelberg:2001:JQH

[Win01]  David Wincelberg.  Java

Q&A: How do you use the

                                        javax.crypto package? Dr.

Dobb’s Journal of Software Tools, 26(4):139–140, April 2001. CODEN DDJOEB. ISSN 1044-789X. URL http:/

                                        /www.ddj.com/ftp/2001/2001_
                                        04/jqa0401.txt;   http://
                                        www.ddj.com/ftp/2001/2001_
                                        04/jqa0401.zip.

Withers:2000:AIZ

[Wit00] Wm. Douglas Withers. Augural image zooming. Dr. Dobb’s Journal of Software Tools, 25(8):48, 50–54, 56, 58, August 2000. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2000/2000_08/aizoom.
                                        zip.

Wilson:2000:IBP

[WK00a] James Y. Wilson and Jason A. Krontz. Inside Bluetooth: Part I. Dr. Dobb’s Journal of Software Tools, 25(3):62, 64–66, 68, 70, March 2000. CODEN DDJOEB. ISSN 1044-789X.

Wilson:2000:BSP

[WK00b] James Y. Wilson and Jason A.

Kronz. The Bluetooth spec: Part II. Dr. Dobb’s Journal of Software Tools, 25(4):58, 60, 62–64, April 2000. CODEN DDJOEB. ISSN 1044-789X.

Wurmsdobler:2000:MMR

[WM00] Peter Wurmsdobler and Nicholas McGuire. MiniRTL: A minimal real-time Linux. Dr. Dobb’s Journal of Software Tools, 25(12):48, 50, 52, 54, December 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_12/minirtl.
txt.

Wilkinson:2000:LFS

[WNL^{+}00] Dennis Wilkinson, Stefan Nilsson, Murray Lesser, Kerry L. Bonin, Jonathan Erickson, Jeff Davis, and Ze’ev Atlas. Letters: Fast sorting; Microsoft: Arrogant or dumb?; patents: Myth versus reality. Dr. Dobb’s Journal of Software Tools, 25(10):10, 12, October 2000. CODEN DDJOEB. ISSN 1044-789X.

Woehr:2000:CLI

[Woe00] Jack Woehr. A C++ library for IBM MQSeries. Dr. Dobb’s Journal of Software Tools, 25 (7):52, 54, July 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_07/
cppmqs.txt;   http://www.

                                        ddj.com/ftp/2000/2000_07/
                                        jxmqv1p0.zip.

Woehr:2001:PBR

[Woe01] Jack J. Woehr. Programmer’s bookshelf: Reading embedded. Dr. Dobb’s Journal of Software Tools, 26(11):128, 130, November 2001. CODEN DDJOEB. ISSN 1044-789X.

Woehr:2002:PBP

[Woe02a]  Jack Woehr.  Program-

mer’s bookshelf: Patterns and MPEG-4. Dr. Dobb’s Journal of Software Tools, 27(12): 78, December 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7718/
                                        ddj0212j/.

Woehr:2002:PBE

[Woe02b] Jack J. Woehr. Programmer’s bookshelf: Embedded systems programming. Dr. Dobb’s Journal of Software Tools, 27 (10):86, October 2002. CO-DEN DDJOEB. ISSN 1044-789X.

Woehr:2002:PBI

[Woe02c] Jack J. Woehr. Programmer’s bookshelf: Intrinsic difficulty. Dr. Dobb’s Journal of Software Tools, 27(9):101, September 2002. CODEN DDJOEB. ISSN 1044-789X.

Woehr:2003:PBL

[Woe03] Jack J. Woehr. Programmer’s bookshelf: Linux, XP, and everything in between. Dr. Dobb’s Journal of Software Tools, 28(3):69, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7826/
ddj0303j/.

Wraxall:2001:JQH

[Wra01] Jason Wraxall. Java Q&A:

How can I extend the Swing JComboBox? Dr. Dobb’s Journal of Software Tools, 26(1):181–182, 184, 186, January 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_01/jqa0101.
txt;   http://www.ddj.com/
ftp/2001/2001_01/jqa0101.
zip.

Wright:2003:JES

[Wri03] William Wright. J2Me and embedded systems. Dr. Dobb’s Journal of Software Tools, 28 (3):54, 56, 58, March 2003. CODEN DDJOEB. ISSN 1044-789X. URL http://www.

ddj.com/documents/s=7826/
ddj0303g/.

Wincelberg:2002:LIE

[WWL^{+}02] David Wincelberg, Sy Wong, Dan Leach, Paul Keister, and Robert Masta. Letters: Inside eBook security; numerical weather forecasting; FrontPage EULA; priority queues; audio watermarking versus compression. Dr. Dobb’s Journal of Software Tools, 27(3): 10, March 2002. CODEN DDJOEB. ISSN 1044-789X.

Yam:2000:EEC

[Yam00] Michael Yam. EC: A Euro calculator for the Palm platform. Dr. Dobb’s Journal of Software Tools, 25(1):56, 58, 60, 62, 64, January 2000. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/ftp/2000/2000_01/
                                        palmeuro.zip.

Yuan:2002:JQH

[Yua02]  Michael J. Yuan.  Java

Q&A: How do I map SQL database tables to Java objects? Dr. Dobb’s Journal of Software Tools, 27(3):89–90, 92–94, March 2002. CODEN DDJOEB. ISSN 1044-789X.

                                        URL http://www.ddj.com/
                                        ftp/2002/2002_03/jqa0302.
                                        txt;   http://www.ddj.com/
                                        ftp/2002/2002_03/jqa0302.
                                        zip.

Yudkowsky:2002:VBA

[Yud02] Moshe Yudkowsky. Voice biometrics and application security. Dr. Dobb’s Journal of Software Tools, 27(11):16, 18, 20, 22, November 2002. CO-DEN DDJOEB. ISSN 1044-

                                        789X.  URL http://www.
                                        ddj.com/documents/s=7644/
                                        ddj0211b/.

Yuen:2001:TPS

[Yue01] Andy Yuen. A tiny Perl Server Pages engine. Dr. Dobb’s Journal of Software Tools, 26 (8):71, 73, 77–78, 80, 82, 84, August 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_08/psp.zip;
http://www.ddj.com/ftp/
2001/2001_08/psplist.zip.

Yun:2000:MNS

[Yun00]  James K. Yun.  Measur-

ing network software performance. Dr. Dobb’s Journal of Software Tools, 25(3):80, 82–91, March 2000. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2000/2000_03/netperf.
txt;   http://www.ddj.com/
ftp/2000/2000_03/netperf.
zip.

Yuen:2002:TPS

[YY02] Andy Yuen and Hung-Wai Yuen. Tiny Perl Server Pages and MySQL. Dr. Dobb’s Journal of Software Tools, 27(6):26, 28, 30, 32, 34, June 2002. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2002/2002_06/
pspsql.zip;   http://www.
ddj.com/ftp/2002/2002_06/
pspsql_listings.zip.

Zanger:2001:SCP

[Zan01] Ana Zanger. Surveying computer programmers. Dr. Dobb’s Journal of Software Tools, 26(4):??, April 2001. CODEN DDJOEB. ISSN 1044-789X.

Zhao:2001:SAL

[ZBKO01] Feng Zhao, Christopher Bailey-Kellogg, and Ivan Ordonez.

The Spatial Aggregation Language. Dr. Dobb’s Journal of Software Tools, 26(4):68, 70, 72–74, 76–77, April 2001. CO-DEN DDJOEB. ISSN 1044-

789X. URL http://www.ddj.
com/ftp/2001/2001_04/sal.
txt.

Zeitlin:2001:WCP

[Zei01] Vadim Zeitlin. The wxWindows cross-platform framework. Dr. Dobb’s Journal of Software Tools, 26(5):106, 109– 110, 112, May 2001. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2001/2001_05/
wxwin.txt;   http://www.
ddj.com/ftp/2001/2001_05/
wxwin.zip.

Zeller:2001:VDD

[Zel01] Andreas Zeller. Visual debugging with DDD. Dr. Dobb’s Journal of Software Tools, 26(3):21–22, 24, 26, 28, March 2001. CODEN DDJOEB. ISSN 1044-789X.

URL http://www.ddj.com/
ftp/2001/2001_03/vdebug.
txt.

Zhang:2000:WMH

[Zha00] Peter Zhang. Webrelay: A multithreaded HTTP relay server. Dr. Dobb’s Journal of Software Tools, 25(2):86, 88, 90–94, 96, February 2000. CO-DEN DDJOEB. ISSN 1044-

789X.  URL http://www.
ddj.com/ftp/2000/2000_02/
webrelay.txt;   http://www.

ddj.com/ftp/2000/2000_02/
webrelay.zip.

Zyzyck:2003:RGP

[Zyz03] Jon Zyzyck. A report generator for PC-Lint. Dr. Dobb’s Journal of Software Tools, 28 (2):52, 54, 58, 60, February 2003. CODEN DDJOEB. ISSN 1044-789X. URL http:/

/www.ddj.com/documents/s=
7790/ddj0302g/.